Mon Jun 23 17:54:51 2008	ZZz	ZZ	zqa AM	zxxx AM	f30cryo	cryoimaging	asas			
Mon Jun 30 14:19:46 2008	morphew	04/15	9:00 AM	4:45 PM	tiltrotate	rtimage	tryps	4		
Mon Jun 30 14:21:55 2008	schwartz	04/16	9:15 AM	5:30 PM	f30cryo	cryotomo	QDs, Giardia	4		
Mon Jun 30 14:22:49 2008	Schwartz	4/17	10:15 AM	10:45 AM	standard	serviceman	new alignments		cindi040708 new alignments
	
Mon Jun 30 14:23:48 2008	schwartz	04/17	10:45 AM	4:30 PM	f30cryo	cryoimaging	Chloroplasts, Giardia			
Mon Jun 30 14:24:30 2008	cedric	04/17	4:30 PM	10:00 PM	f30cryo	cryoimaging	chloro			
Mon Jun 30 14:25:14 2008	morphew	04/18	8:45 AM	2:30 PM	tiltrotate	rttomo	tryps	4		
Mon Jun 30 14:25:52 2008	schwartz	04/18	2:30 PM	4:15 PM	standard	rtimage	eels		couldn't figure out eels	
Mon Jun 30 14:27:03 2008	cedric	04/22	8:15 AM	10:00 PM	914	cryotomo	section chloro	4		motor makes nnoise in one direction 
(right on the screen) at -50
Mon Jun 30 14:27:46 2008	morphew	04/23	1:30 PM	4:30 PM	tiltrotate	rttomo	thylakoid	2		
Mon Jun 30 14:28:37 2008	morphew	04/24	8:45 AM	5:00 PM	tiltrotate	rttomo	thylakoid	7		stage movement at x=500 no good

Mon Jun 30 14:29:31 2008	schwartz	04/25	8:45 AM	4:45 PM	70	rttomo	eels	1		stage locked up 1X, reset goniometer
Mon Jun 30 14:30:07 2008	morphew	04/28	9:00 AM	5:30 PM	70	rttomo	u20s/melanoma	4		
Mon Jun 30 14:30:44 2008	morphew	04/29	9:00 AM	4:00 PM	70	rttomo	melanoma/3D-matrix	2		
Mon Jun 30 14:31:14 2008	morphew	04/30	8:45 AM	12:00 PM	tiltrotate	rttomo	thylakoid	4		
Mon Jun 30 14:31:46 2008	schwartz	04/30	1:00 PM	4:30 PM	standard	rtimage	eels			
Mon Jun 30 14:32:25 2008	schwartz	05/01	9:00 AM	4:30 PM	standard	rtimage	eels		replaced std. o-ring	
Mon Jun 30 14:32:58 2008	morphew	05/02	9:00 AM	2:00 PM	tiltrotate	rttomo	thylakoid	5	
	
Mon Jun 30 14:33:35 2008	Schwartz/Andy	05/02	2:00 PM	3:30 PM	standard	rtimage	eels
eels			
Mon Jun 30 14:34:23 2008	EOT	05/04	12:00 PM	6:30 PM	tiltrotate	rttomo	u2os	3		
Mon Jun 30 14:34:56 2008	EOT	05/06	8:30 AM	4:30 PM	tiltrotate	rttomo	u2os	2		
Mon Jun 30 14:35:44 2008	morphew	05/07	9:00 AM	3:30 PM	tiltrotate	rtimage	AFP transfected COS	4	cindi changed o-ring on t/r rod	
Mon Jun 30 14:36:30 2008	morphew	05/08	10:00 AM	1:00 PM	tiltrotate	rttomo	AFP transfected COS	2		
Mon Jun 30 14:37:32 2008	cedric	05/08	2:00 PM	5:30 PM	f30cryo	cryoimaging	afp sections			stage doesn't move with trackball: 
10 times to move correctly...pain 
in the ass

Mon Jun 30 14:38:08 2008	morphew	05/13	9:00 AM	4:30 PM	tiltrotate	rttomo	melanoma	2		
Mon Jun 30 14:38:37 2008	morphew	05/14	9:00 AM	2:00 PM	tiltrotate	rttomo	melanoma/spc42	2		
Mon Jun 30 14:39:12 2008	EOT	05/15	8:30 AM	2:00 PM	tiltrotate	rttomo	cilia	4		
Mon Jun 30 14:39:45 2008	schwartz	05/15	2:00 PM	2:30 PM	standard	serviceman	reboot			reboot computer to see if beam gets 
brighter
Mon Jun 30 14:40:20 2008	morphew	05/16	9:00 AM	1:00 PM	tiltrotate	rttomo	melanoma	4		rod sluggish on rotation
Mon Jun 30 14:41:09 2008	janet/odde	05/16	1:00 PM	5:00 PM	tiltrotate	rttomo	ase1delete yeast spindles	4		rod sluggish on rotation
Mon Jun 30 14:41:48 2008	morphew	05/22	2:00 PM	4:00 PM	tiltrotate	rttomo	olin's nuclei	1		
Mon Jun 30 14:42:32 2008	Janet/Odde	05/29	8:10 AM	6:00 PM	tiltrotate	rttomo	ase1delete yeast spindles	10		T/R motor not working at all
Mon Jun 30 14:43:01 2008	cedric	05/30	8:00 AM	5:00 PM	914	cryoimaging	sections			
Mon Jun 30 14:43:40 2008	power outage	05/30	10:20 PM	10:40 PM	standard	serviceman	campus wide power outage			
Mon Jun 30 14:44:20 2008	service	06/02	8:00 AM	1:00 PM	standard	serviceman	cold start			cold start, upper camera warmed, 
refilled sf6
Mon Jun 30 14:44:51 2008	morphew	06/02	1:00 PM	6:15 PM	70	rttomo	olins	4		
Mon Jun 30 14:45:26 2008	morphew	06/03	8:45 AM	4:50 PM	70	rttomo	olins	4		
Mon Jun 30 14:46:02 2008	schwartz	06/04	10:00 AM	1:15 PM	f30cryo	cryoimaging	testing vitreous grids		not vitreous :(	
Mon Jun 30 14:46:29 2008	morphew	06/04	1:00 PM	5:00 PM	70	rttomo	olins	4		
Mon Jun 30 14:46:51 2008	morphew	06/05	9:00 AM	4:00 PM	70	rttomo	olins	6		
Mon Jun 30 14:47:31 2008	Janet/Odde	06/06	8:30 AM	3:00 PM	70	rttomo	ase1delete yeast spindles	6		
Mon Jun 30 14:48:06 2008	DNM	06/06	3:30 PM	5:00 PM	standard	development	knife edge on GIF
knife edge on GIF			
Mon Jun 30 14:48:47 2008	cedric	06/10	12:30 PM	4:30 PM	914	cryoimaging	sections			from previous entry: video camera 
stuck in
Mon Jun 30 14:49:25 2008	Janet/Winey	06/11	11:30 AM	4:45 PM	70	rttomo	WT tet bbs	6		
Mon Jun 30 14:49:57 2008	EOT	06/12	9:00 AM	3:30 PM	70	rttomo	dynamin 1,2 KO	3		
Mon Jun 30 14:50:31 2008	EOT	06/13	9:00 AM	2:30 PM	70	rttomo	dynamin 1,2 KO	6		
Mon Jun 30 14:51:00 2008	morphew	06/16	9:00 AM	5:00 PM	70	rttomo	melanoma	2		
Mon Jun 30 14:51:55 2008	morphew	06/19	9:00 AM	4:00 PM	f20cryo	cryoimaging	BPV			vacuum at 32 at start-up (also on 
6/16)
Mon Jun 30 14:52:34 2008	DNM	06/23	8:30 AM	12:00 PM	standard	development	testing		testing rectangular distortion 
field and skipping montage pieces	
Mon Jun 30 14:53:06 2008	morphew	06/24	8:30 AM	1:00 PM	f20cryo	cryotomo	BPV	1		vacuum high (29) at start-up
Mon Jun 30 14:53:43 2008	morphew	06/27	12:00 PM	3:45 PM	70	rttomo	BPV	4		vacuum high at startup
Mon Jun 30 14:58:11 2008	Morphew	06/29	9:00 AM	2:30 PM	f20cryo	cryotomo	BPV	5		
Tue Jul  1 16:41:47 2008	Cindi	7/1	9:15 AM	4:45 PM	standard	rtimage	EELS practice			Rigged air regulator in TV camera 
controller.
Wed Jul  2 16:26:59 2008	Cindi	7/2	2:00 PM	4:25 PM	standard	rtimage	eels		AUGH	
Thu Jul  3 12:35:01 2008	morphew	07/02	9;00 AM	12:30 PM	70	rttomo	PAPILLOMA VIRUS	0	i DUMPED THE VAC UPON ROD 
INSERTION. vAC FAILED TO RECOVER ON 
ITS OWN.	
Thu Jul  3 17:08:13 2008	cindi	7/3	2:00 PM	5:00 PM	standard	rtimage	eels, lipids			
Tue Jul  8 08:57:36 2008	Cindi/Alexandra	07/07	9:00 AM	12:00 PM	standard	rtimage	scope down			turbo wasn't working correctly
Tue Jul  8 08:58:15 2008	Service/Chuck Moe	7/7	12:00 PM	2:00 PM	standard	serviceman	worked on turbo pump			diagnosed and got it working again
Tue Jul  8 08:59:09 2008	service/Chuck Moe	7/8	7:00 AM	9:00 AM	standard	serviceman	drag pump			replaced drag pump, still doesn't 
work. controller card is bad.
Fri Jul 11 08:40:36 2008	Service/Chuck Moe	7/9	10:30 AM	12:30 PM	standard	serviceman	install new turbo pump assembly			received old version of turbo pump 
(68). Installed and it was bad. It 
was the last one available and all 
of the rest are on backorder.
Fri Jul 11 08:41:06 2008	Scope Down	7/10	8:00 AM	6:00 PM	standard	serviceman	scope down, bad turbo			
Fri Jul 11 08:41:33 2008	Scope Down	7/11	8:00 AM	6:00 PM	standard	serviceman	scope down, bad turbo			
Mon Jul 14 14:58:04 2008	Service/Chuck Moe	7/14	10:30 AM	1:00 PM	standard	serviceman	replace turbo pump			turbo pump arrived. It was not 
reconditioned, so it was not 
working. Escalated to Head of 
Logistics.
Tue Jul 15 09:36:06 2008	DNM	7/15	9:15 AM	9:30 AM	standard	development	Checking out shuttering with both 
idle states closed	0		
Tue Jul 22 17:58:30 2008	Julia Cope	07/22	10:30 AM	5:45 PM	914	cryoimaging	Kar3Vik1 mt's	0	Not vitrous vitrobot samples :(	No water flow or current to 
objective lens - Chuck fixed it but 
beam was really wacky.  He did some 
fancy FEG realigments and it was ok 
after that.  Crashed vacuum twice, 
but came back relatively quickly.  
Stage makes a screeching noise when 
moving long distances for lmm. 
Wed Jul 23 16:34:16 2008	Cindi	07/23	9:40 AM	4:30 PM	f30cryo	cryoimaging	cryo sections PTKGFP			GIF beam was offset from regular 
beam
Fri Jul 25 12:06:55 2008	morphew	7/25	11:00 AM	12:00 PM	standard	rtimage	silicon oxide film check	0		
Mon Jul 28 12:16:53 2008	bouchet-marquis	07/25	09:00 AM	17:30 PM	914	cryoimaging	sections	0		crashed two times the vaccum at the 
insertion of the rod.
crashed first rod at RT
then crashed when rod was cold.
Mon Jul 28 16:49:41 2008	morphew	07/28	9;00 AM	4:50 PM	f20cryo	cryotomo	whole mount c127/bpv infect	5		
Tue Jul 29 14:23:33 2008	Cindi	07/29	10:15 AM	2:20 PM	f30cryo	cryoimaging	sxns 			stage will not move in Y, called 
service.
Wed Jul 30 09:22:59 2008	Service R Spilman	12/30/2008	8:45 AM	9:20 AM	standard	serviceman	Compustage call N11529		Y donator Noisey in negative 
direction.  Ordered new donator to 
arrive Thursday 7/31/2008 5322-695-
15858  SEO 48851	Y donator Noisey in negative 
direction.  Ordered new donator to 
arrive Thursday 7/31/2008 5322-695-
15858  SEO 48851
Wed Jul 30 18:36:34 2008	morphew	07/29	9 AM	6:30 PM	70	rttomo	trypanosomes	8 2x1		
Thu Jul 31 10:35:40 2008	Service rspilman	7/31/2008	9:00 AM	10:30 AM	standard	serviceman	Compustage 	0	Replaced donator for Y 
translation.  Adusted Y movements.	Replaced donator for Y 
translation.  Adusted Y movements.
Fri Aug  1 10:50:18 2008	Cindi	07/31	2:30 PM	3:00 PM	standard	rtimage	giardia neg stain			
Fri Aug  1 10:50:36 2008	Cindi	07/31	6:00 PM	8:00 PM	f30cryo	cryoimaging	sxns			
Fri Aug  1 16:29:23 2008	morphew	08/01	9 AM	4:30 PM	70	rttomo	trypanosomes	6		
Mon Aug  4 16:47:00 2008	morphew	8/4	9 AM	4:45 PM	f30cryo	cryotomo	c127 bpv infect	3		
Mon Aug 11 17:05:38 2008	Eileen O'Toole	8/12	8:30 AM	5:00 PM	70	rttomo	Odde metaphase yeast	7		The compustage had a siezure when I 
was adjusting eucentricity this 
morning. Called Chuck and he walked 
me through reseting. 
Tue Aug 12 16:51:48 2008	Eileen	8/12	8:30 AM	5:00 PM	70	rttomo	Odde metaphase yeast	8		
Wed Aug 13 16:42:42 2008	morphew	08/13	9 AM	4:35 PM	f20cryo	cryotomo	bpv isolated	1		Stage started going hay-wire at 
touch of translation (not tilted). 
Disabled stage, removed holder, 
tried to re-enable, wouldn't go. 
Tried 4 times. Abandoned scope, 
called and left message for Chuck, 
waited, started cryo-cycle, tried 
once again, this time it did re-
enable. Left scope without holder 
in stage (since cryocycle was 
going). Called Chuck and left 
another message.
Thu Aug 14 16:43:54 2008	morphew	08/14	9 AM	4:45 PM	70	rttomo	bpv isolated cryo, trypanosomes 
plastic	5		
Mon Aug 18 16:13:27 2008	morphew	08/18	8:45 AM	4:15 PM	70	rttomo	malaria--Tim Spurk, Univ of 
Melbourne	4		
Tue Aug 19 17:07:08 2008	morphew	8/19/08	2 PM	5:15 PM	70	rttomo	melanoma	2		
Tue Aug 19 17:08:00 2008	Jose Maria	8/19	9 AM	2 PM	70	rttomo	plant nuclei	3		
Thu Aug 21 16:07:50 2008	Cindi	8/21	10:00 AM	4:00 PM	f30cryo	cryotomo	Eg5 Microtubules	4		Stage tried to go to -1288 in Y 
during LMM. Took holder out, reset 
goniometer, wouldn't reset. David 
manually moved the stage about 1mm, 
then it reset. C2 aperture was at 
150um instead of 100um
Fri Aug 22 17:16:01 2008	Jose Maria	08/22	10 AM	5:30 PM	standard	rttomo	Tomography of plant nuclear 
structures	8	Hi Cindi,

Besides the problem mentioned 
below, I just wanted to thank you 
for all of your help during these 
days.

I will leave next week, and I hope 
to find you around on tuesday to 
say you goodbye.	At low mag, the beam was doing 
weird things, like losing their 
circular shape or projecting like 
shadows. However, at x20000, 
everything was perfect. It did not 
bother me at all.
Mon Aug 25 12:24:59 2008	Chuck/Service	08/25	8:00 AM	9:30 AM	f30cryo	serviceman	check stage			worked on F30 stage problem and tightened the 
y-drive. tested movement with cryo holder. 
going to each y extreme and tilting with no y-
axis error.
Mon Aug 25 16:07:47 2008	morphew	8/25	11:00 AM	4:00 PM	70	rttomo	melanoma cells correlative	3		
Tue Aug 26 12:16:45 2008	cindi	08/26	9:30 AM	10:30 AM	f30cryo	cryotomo	eg5 MTs			stage got stuck. Chuck had to come
Tue Aug 26 12:21:39 2008	Service/Chuck	08/26	11:00 AM	12:30 PM	f30cryo	cryotomo	eg5 MTs			stage got stuck in Y axis . Rebuilt 
Y axis motor assy and tested the 
movement
Thu Aug 28 15:51:01 2008	morphew	08/28	9 AM	3:50 PM	70	rttomo	freeze sub tests	2		
Fri Aug 29 14:28:02 2008	morphew	8/29	9 AM	2:30 PM	f30cryo	cryotomo	BPV/DTT treat	2		
Tue Sep  2 16:04:15 2008	Eileen	9/2	9 AM	4 PM	tiltrotate	rttomo	chlamy mitotic	2	first session with returned tilt 
rotate rod. Worked well!	
Wed Sep  3 15:52:11 2008	cindi	9/3	9:30 AM	4:00 PM	f30cryo	cryotomo	eg5 MTs	2	Also looked at plastic sections in 
cryo-holder.	
Thu Sep  4 16:26:20 2008	morphew	9/4	9 AM	4:24 PM	tiltrotate	rttomo	tryps	4		
Fri Sep  5 15:50:10 2008	morphew	9/5	2 PM	3 PM	standard	rtimage	silicon antibody binding			
Fri Sep  5 15:51:25 2008	cindi	9/5	3:15 PM	3:50 PM	standard	development	calibrations for ctf noise			
Mon Sep  8 15:23:38 2008	Eileen	9/8	8 AM	3:30 PM	70	rttomo	U2OS mitotics	4		Gave 'stage end of range' two times 
when taking the rod out of the 
scope although it had been centered 
in x/y. No problem, though, just a 
weird message.
Tue Sep  9 16:39:29 2008	cindi	9/9	9:30 AM	4:45 PM	914	cryoimaging	S-layers, BPV infected cells		Tried diffraction to see when cells 
de-vitrify. Also used F30 cryo 
holder.	
Fri Sep 12 08:35:43 2008	cedric	9/11	10 AM	12 PM	914	cryotomo	vitreous sections	0	2 hard vac crashes. quit for the day	
Fri Sep 12 08:36:53 2008	eileen	9/11	1 PM	3:15 PM	tiltrotate	rttomo	U2Os overlap zone	1	1x2 montage 	
Fri Sep 12 15:12:00 2008	eileen	9/12	8:15 AM	3:15 PM	tiltrotate	rttomo	U2OS midzone	5	1x2 montages	
Tue Sep 16 17:39:05 2008	eileen	9/16	9 AM	6 PM	tiltrotate	rttomo	U2OS supermontage, section 1	4	Four,3x3 montages for 2x2 super	
Thu Sep 18 12:54:45 2008	cindi	09/18	9:30 AM	1:00 PM	f30cryo	cryoimaging	PTK sections from cryo-LM		Nothing to see here.	
Mon Sep 22 11:42:04 2008	cindi	9/22	10:00 AM	11:30 AM	914	cryoimaging	giardia		grid bad	crashed vacuum inserting rod

Mon Sep 22 16:07:32 2008	cindi	9/22	11:30 AM	4:15 PM	f30cryo	cryotomo	giardia	4	sample looked good	
Wed Sep 24 13:39:54 2008	cindi	9/24	12:00 PM	1:45 PM	914	cryoimaging	slayers		first grid had cauliflower ice and 
the rest were too thick	lost ability to select holder when 
inserting rod
Wed Sep 24 17:19:08 2008	Matt West	9/23	7:30 AM	7:30 PM	standard	rttomo	Yeast ER	14	This session was for Gia Voeltz lab.
Scope worked great!	
Thu Sep 25 14:54:00 2008	morphew	9/25/08	8:45 AM	3:00 PM	f30cryo	cryoimaging	melanoma whole mount	0		
Fri Sep 26 15:57:15 2008	cindi	9/26	9:30 AM	3:30 PM	f30cryo	cryotomo	s-layers	2		wouldn't let me pick the holder, 
but still worked
Mon Sep 29 16:50:08 2008	morphew	9-29	9 AM	4:45 PM	tiltrotate	rttomo	chloroplast	8		
Tue Sep 30 15:27:12 2008	morphew	9-30	9 AM	3:30 PM	f30cryo	cryotomo	bpv	4		
Thu Oct  2 16:53:35 2008	morphew	10-2	10 AM	5 PM	tiltrotate	rttomo	melanoma	4		
Fri Oct  3 15:48:29 2008	morphew	10-3	1 PM	4 PM	tiltrotate	rttomo	bpv			
Mon Oct  6 18:01:25 2008	morphew	10/6	8:30 AM	6 PM	f30cryo	cryotomo	bpv	6		
Tue Oct  7 13:21:24 2008	Eileen	10/7/08	8:50 AM	1:20 PM	tiltrotate	rttomo	isolated thylakoids	4		
Thu Oct  9 12:30:24 2008	eileen	10/9/08	8:45 AM	12:30 PM	70	rttomo	isolated thylakoids	4	extraction voltage was below 4250 
so beam was dimmer than usual. I 
didn't notice until after the 
alignments.	
Fri Oct 10 16:54:37 2008	morphew	10/10/08	9 AM	5 PM	tiltrotate	rttomo	bpv	2-2x2		
Sat Oct 11 01:34:05 2008	Johanna	10/10	8 PM	01:30 AM	fishione	rttomo	Trypanosomes	4		
Sun Oct 12 04:03:51 2008	Johanna	10/11	5:00 PM	04:00 AM	fishione	rttomo	trypanosomes	10	Montages	
Mon Oct 13 04:10:32 2008	Johanna	12/10	3:00 PM	4:00 AM	fishione	rttomo	trypanosomes	9	montages	
Mon Oct 13 17:00:06 2008	morphew	10/13	3 AM	5 PM	tiltrotate	rttomo	thylakoid	2		
Tue Oct 14 16:42:47 2008	morphew	10-14	8:45 AM	4:45 PM	f30cryo	cryotomo	bpv-infected whole mount	7		
Wed Oct 15 23:47:19 2008	Johanna	10/15	7:00 PM	00:00 PM	fishione	rttomo	trypanosomes	7	montages	
Thu Oct 16 08:22:03 2008	cindi	10/15	10:30 PM	5:00 PM	f30cryo	cryoimaging	NRK cryo-sections		The correlation from LM to EM 
worked quite well. 	
Thu Oct 16 21:10:50 2008	Johanna	10/16	3:00 PM	9:00 PM	fishione	rttomo	trypanosome	8		
Fri Oct 17 16:33:15 2008	Janet Meehl	10/17	12:15 PM	4:45 PM	tiltrotate	rttomo	Tetrahymena exchange junction	3		
Mon Oct 20 11:03:21 2008	Johanna	10/18	3:00 PM	11:00 PM	fishione	rttomo	trypanosome	6		
Mon Oct 20 20:47:42 2008	cindi	10/20	10:30 AM	8:45 PM	f30cryo	cryotomo	ice and cryo-sections	3	5th grid's the charm :)	
Tue Oct 21 18:16:33 2008	eileen	10/21/08	9 AM	6:30 PM	tiltrotate	development	auto-acquire with SerialEM	2 3x3	David and I worked on setting up 
supermontage auto-acquiring 
functions with Macros, Navigator 
and Serialem. Very cool! 2 3x3 
montages from a supermontage map	
Wed Oct 22 01:45:57 2008	JOhanna	10/21	06:30 PM	02:00 AM	fishione	rttomo	leishmania	6		
Wed Oct 22 14:44:01 2008	cindi	10/22	9:15 AM	2:45 PM	f30cryo	cryotomo	giardia	2	test of correlative lm/em	
Thu Oct 23 13:27:33 2008	Janet Meehl/Eric Cole	10/23	8:00 AM	1:30 PM	tiltrotate	rttomo	nuclear exchange junction - Tet	6	WOW!	
Fri Oct 24 18:16:56 2008	Johanna	10/24	9:00 AM	6:15 PM	fishione	rttomo	Leishmania	14		
Tue Oct 28 16:40:04 2008	cindi	10/28	10:15 AM	4:45 PM	f30cryo	cryotomo	giardia + NC1Kinesin	3	4th grid's the charm.	Column Vacuum was at 20 when I 
logged in. O-ring was a bit dirty.
Wed Oct 29 15:46:20 2008	cindi	10/29	9:30 AM	3:45 PM	f30cryo	cryoimaging	cryo-sxns and BPV infected cells		sxns lost and cells not vitreous :(	
Wed Nov  5 14:18:57 2008	Janet Meehl/Eric Cole	11/5	10:50 AM	2:15 PM	tiltrotate	rttomo	Tetrahymena nuclear exchange 
junction	0		
Wed Nov 12 15:29:39 2008	eileen	11/12/08	8:30 AM	3:45 PM	tiltrotate	rttomo	Andrea R. synapses	8		
Thu Nov 13 14:39:11 2008	eileen	11/13	8:30 AM	2:30 PM	tiltrotate	rttomo	Andrea's 3minute synapse	4 1x2		
Sat Nov 15 22:23:58 2008	Johanna	11/15	1:30 PM	11:00 PM	fishione	rttomo	tryps	4	montages	
Mon Nov 17 00:39:35 2008	Johanna	11/17	2:00 PM	00:30 AM	fishione	rttomo	tryps	7	montages	
Tue Nov 18 14:58:11 2008	Ron	11/17	1:30 PM	5:00 PM	standard	serviceman	Replace objective aperture lever			replace objective aperture lever
Tue Nov 18 15:00:19 2008	Ron	11/18	11:30 PM	3:00 PM	standard	serviceman	preventative maintenance			replace hoses, check chiller, 
change pump oil, replace water 
filter, checked image, electronic 
vacuum test
Tue Nov 18 17:10:31 2008	eileen	11/18	3:30 PM	5:15 PM	70	rtimage	chlamy mitotics	0	The +/- rod looked slightly bent.	
Wed Nov 19 16:47:43 2008	morphew	11/19	9 AM	5 PM	tiltrotate	rttomo	tryps	6		
Thu Nov 20 15:46:12 2008	eileen	11/20	8:25 AM	3:45 PM	tiltrotate	rttomo	yeast and chlamy	2 1x2		
Fri Nov 21 15:38:10 2008	eileen	11/21	8:30 AM	3:45 PM	tiltrotate	rttomo	chlamy, yeast	3		
Tue Nov 25 18:30:53 2008	DNM	11/25	2:15 PM	6:30 PM	standard	development	cooker and autocenter beam			
Wed Nov 26 13:26:41 2008	eileen	11/26	8:30 AM	1:30 PM	tiltrotate	rttomo	yeast spindles	6		
Wed Nov 26 17:47:14 2008	DNM	11/26	1:30 PM	5:50 PM	standard	development	autocenter, cook, special film 
command			The HT went off and would not 
restart when I started using the 
adaexp server in "show" mode.  I 
had to reboot.  Then I started 
using it in SerialEM and the HT did 
the same thing.  Another reboot, 
and things went fine for 2 hours, 
so it really was the adaexp 
server.  I will get a proper 
version for the scope from Max 
Otten.
Tue Dec  2 16:28:58 2008	eileen	12/2	8:30 AM	4:30 PM	tiltrotate	rttomo	chlamy	5 1x2		
Wed Dec  3 15:09:08 2008	service	12/1	8:00 AM	5:00 PM	standard	serviceman	F30 computer off			needs new power supply. put in 
temporary power supply
Wed Dec  3 15:09:48 2008	cindi	12/3	10:00 AM	3:00 PM	tiltrotate	rtimage	giardia		looked like crap	
Thu Dec  4 15:26:23 2008	eileen	12/4	11:00 AM	3:30 PM	tiltrotate	rttomo	chlamy	2		
Fri Dec  5 15:00:36 2008	david	12/5	10:30 AM	11:00 AM	standard	serviceman	replace power supply			replace power supply on scope 
computer
Mon Dec  8 15:49:29 2008	morphew	12-8	1 PM	4 PM	fishione	rttomo	yeast	2		
Wed Dec 10 17:10:59 2008	eileen	12/20	10:30 AM	5:00 PM	tiltrotate	rttomo	Pietro De Camilli neurons	4		
Thu Dec 11 12:58:25 2008	eileen	12/11	8:45 AM	1:00 PM	tiltrotate	rttomo	Chlamy mitotics	4 1X2		
Fri Dec 12 16:41:09 2008	Chuck Robertson	12/12	8:30am AM	4:30 PM	fishione	rttomo	Methanosarcina barkerei	4	Ok	
Tue Dec 16 15:45:24 2008	eileen	12/16	9:30 AM	3:45 PM	tiltrotate	rttomo	chlamy mitotics	4 1x2	rod was not put back in the 
microscope from the day before. It 
took a while for the vac to come 
down.	
Wed Dec 17 15:34:46 2008	morphew	12-16	9 AM	4 PM	tiltrotate	rttomo	bpv, baas neurons	3		
Thu Dec 18 15:13:28 2008	eileen	12/17/08	8:45 AM	3:15 PM	tiltrotate	rttomo	DeCamilli serials	3 2x2		
Fri Dec 19 16:47:42 2008	Janet Meehl/Winey	12/19	9:10 AM	4:45 PM	tiltrotate	rttomo	WT tetrahymena duplicatin basal 
bodies	11		
Mon Dec 22 13:24:02 2008	Janet Meehl/Winey	12/22	10:15 AM	1:30 PM	tiltrotate	rttomo	WTtet duplicatin basal bodies	4		
Mon Dec 22 15:49:13 2008	DNM	12/22	2:30 PM	3:45 PM	standard	serviceman	trying to get the GIF going again		Got beam on cameras but it won't 
tune	
Wed Dec 31 02:44:47 2008	Chuck Robertson	12/30	8:30 AM	3:00 AM	fishione	rttomo	Methanosarcina barkeri	20	Previous user (Christine) appears 
not to have made log book entry... 
did not have context for problem 
noted below.

F30 was very stable today, but the 
beam is marginal for my samples. 
Today the beam was 8.26 nA st spot 
size = 2.	On arrival found the extraction 
voltage at 3800... no indication in 
the log of problem or not by 
previous user.  Also, turbo pump 
was off on arrival (I thought it 
was supposed to run all of the time 
now after it cryocycled?)

Brought the EV back to 4250... 
everything appeared not to fail.


Mon Jan  5 14:49:32 2009	eileen	1/5	8:40 AM	2:50 PM	tiltrotate	rttomo	Davis yeast	6		
Mon Jan  5 17:26:39 2009	DNM	1/5	3:00 PM	5:30 PM	standard	serviceman	diagnosing GIF problems with VNC			Scott thinks the replacement board 
is bad or there is something bad 
downstream.
Tue Jan  6 18:38:21 2009	cindi	01/06	8:15 AM	11:45 AM	914	cryotomo	giardia		no samples on grids	
Tue Jan  6 18:38:43 2009	cindi	01/06	2:00 PM	6:40 PM	70	rttomo	giardia		samples look OK	
Tue Jan  6 18:38:52 2009	cindi	01/06	2:00 PM	6:40 PM	70	rttomo	giardia	3	samples look OK	
Thu Jan  8 17:18:24 2009	morphew	1/9	9 AM	5 PM	f30cryo	cryotomo	vessicles	3		
Sat Jan 10 16:36:53 2009	cindi/moises	01/10	10:00 AM	4:30 PM	914	cryotomo	giardia cytoskeletons		no good samples :(	
Mon Jan 12 16:19:50 2009	morphew	1-12	9 AM	4 PM	f30cryo	cryotomo	protein complex-stowell	1		
Tue Jan 13 15:59:36 2009	cindi/moises	01/13	1:00 PM	4:00 PM	914	cryotomo	giardia		crappy grids
	
Tue Jan 13 19:21:23 2009	eileen	1/13	4:00 PM	7:30 PM	tiltrotate	rttomo	Trisha Davis yeast	6		
Wed Jan 14 16:52:56 2009	Julia	1/14	12:45 PM	4:50 PM	f20cryo	cryoimaging	Kar3Vik1 mts, apyrase			
Thu Jan 15 21:40:48 2009	cindi/moises	01/15	8:00 AM	9:45 PM	914	cryotomo	Giardia attached cytoskeletons	10	finally!!	
Fri Jan 16 17:21:45 2009	morphew	1/16	9 AM	5:30 PM	tiltrotate	rttomo	melanoma	4		
Mon Jan 19 16:53:00 2009	cindi	1/19	2:30 PM	4:45 PM	tiltrotate	development	distortion correction images			
Tue Jan 20 17:59:18 2009	eileen	1/20	8:30 AM	6:00 PM	tiltrotate	rttomo	Trisha Davis yeast	12		
Wed Jan 21 17:11:27 2009	morphew	1/21	8:00 AM	5:00 PM	tiltrotate	rttomo	aditi neurons	6, 2x2		
Thu Jan 22 18:02:50 2009	Julia	1/21	9:00 AM	6:00 PM	f20cryo	cryoimaging	Kar3Vik1 mt Apyrase	0	All good	
Fri Jan 23 18:01:25 2009	Janet Meehl/Winey	1/23	10:15 AM	6:00 PM	tiltrotate	rttomo	WT tetrahymena basal bodies	10		
Mon Jan 26 15:41:31 2009	eileen	1/26/09	8:40 AM	3:45 PM	tiltrotate	rttomo	chlamy mitotics	9		
Tue Jan 27 15:18:25 2009	morphew	1-27	8:45 AM	3:15 PM	tiltrotate	rttomo	aditi, neurons	4		
Tue Jan 27 18:11:43 2009	eileen	1/27	3:15 PM	6:30 PM	tiltrotate	rttomo	Chlamy poles	4		
Wed Jan 28 17:53:29 2009	Janet Meehl/Winey	1/28	1:30 PM	6:00 PM	tiltrotate	rttomo	WT tetrahymena basal bodies	4		
Thu Jan 29 19:35:34 2009	Matthew West	01/28	9:00 AM	7:45 PM	tiltrotate	rttomo	WT ER	16	Voeltz lab	
Fri Jan 30 15:29:11 2009	eileen	1/30	11 AM	3:30 PM	tiltrotate	rttomo	chlamy mitotics	4		
Mon Feb  2 14:36:56 2009	eileen	2/2	8:45 AM	2:30 PM	tiltrotate	rttomo	chlamy interphase	6		
Tue Feb  3 19:19:13 2009	cedric	02/03	09:00 AM	20:00 PM	914	cryotomo	section	4		
Wed Feb  4 16:57:43 2009	morphew	2-4	9 AM	5 PM	f30cryo	cryotomo	c127 w/bpv	6		
Thu Feb  5 16:09:43 2009	eileen	2/5	9 AM	4:15 PM	tiltrotate	rttomo	RNAi gamma	5		
Mon Feb  9 15:35:47 2009	eileen	2/9	8:30 AM	3:30 PM	70	rttomo	RNAi gamma	5		
Tue Feb 10 16:05:46 2009	morphew	2/10	8:45 AM	4:15 PM	tiltrotate	rttomo	aditi neurons	4-1x3		
Wed Feb 11 19:15:57 2009	cedric	02/11	09:00 AM	8:00 PM	914	cryotomo	sections	4		
Thu Feb 12 17:32:18 2009	Julia	2/12	9:00 AM	5:30 PM	f30cryo	cryoimaging	kar3vik1 mts apyrase		All good	
Mon Feb 16 16:05:40 2009	morphew	2/16	9 AM	4 PM	f30cryo	cryotomo	c127 virus	3		
Tue Feb 17 16:00:55 2009	cindi	12/17	10:30 AM	4:00 PM	914	cryoimaging	giardia		no good samples	
Wed Feb 18 16:55:26 2009	Chris Arthur	02/18	10:00 AM	5:00 PM	f30cryo	cryotomo	Mouse synaptosomes	0	Crashed vacuum 3 times and figured 
it was not meant to be today. Will 
try again in the morning.	
Thu Feb 19 15:17:51 2009	spilman	2/19	10:00 AM	3:00 PM	standard	serviceman	test			feg hardware falure after HT arc.
Conditioned HT to 305KV

checked rapid Dave open 1194 port 
checked operation with TSG
Thu Feb 19 19:41:02 2009	Chris Arthur	02/19	9:00 AM	8:00 PM	914	cryotomo	Mouse synaptosomes	3	FEG arcing problem but Ron said it 
was okay	
Fri Feb 20 18:07:18 2009	cedric	02/20	13:00 PM	18:00 PM	914	cryoimaging	sections	0		beam blinks when playing with 
intensity
beam vibrates when concentrated
Sun Feb 22 16:29:17 2009	cindi	2/22	9:00 AM	4:30 PM	914	cryotomo	giardia cytoskeletons	8		
Sun Feb 22 17:39:30 2009	cindi	02/22	4:30 PM	5:30 PM	70	rtimage	giardia check-out			
Mon Feb 23 16:05:20 2009	morphew	2-23	8:30 AM	4 PM	f30cryo	cryotomo	3t3's plus motors	5		
Tue Feb 24 18:36:22 2009	Chris Arthur	02/24	9:00 AM	6:45 PM	914	cryotomo	Mouse synaptosomes	4		
Thu Feb 26 17:02:45 2009	Julia	2/26	9:00 AM	5:00 PM	f20cryo	cryoimaging	Eg5A3 mts		scope good, samples bad	
Fri Feb 27 15:50:05 2009	morphew	2-27	9 AM	4 PM	f30cryo	cryotomo	membrane channels	5		
Sun Mar  1 16:33:15 2009	cindi	3/1	9:00 AM	1:30 PM	914	cryoimaging	bad giardia grids			
Sun Mar  1 16:33:35 2009	cindi	3/1	1:30 PM	4:30 PM	tiltrotate	rttomo	giardia	2		
Tue Mar  3 09:36:08 2009	Chris Ackerson	03/02	8:55 AM	4:00 PM	standard	rtimage	EELS attempts			bad columns in CCD?- David M was 
working on this when I left
Tue Mar  3 17:41:03 2009	morphew	3-3	8:15 AM	5:45 PM	tiltrotate	rttomo	aditi neurons	6 2x1		
Wed Mar  4 16:57:36 2009	morphew	3-4	8:15 AM	5 PM	f30cryo	cryotomo	3t3 w/eg5	7		
Thu Mar  5 11:08:07 2009	DNM	3/3	4:00 PM	5:00 PM	standard	serviceman	adjusting defect table in DM for GIF	0		Added some regions to GIF defect 
table.  Aslo determined that the 
bottom/right coordinate in the 
defect table has to be 1 past the 
desired pixel
Thu Mar  5 16:40:20 2009	cindi	3/5	10:30 AM	4:45 PM	914	cryotomo	giardia attached	4		for some reason, my gif was -130 
defocus from the upper camera. just 
changed focus with screen up to 0 
and then everything was fine.
Fri Mar  6 17:10:43 2009	cedric	03/06	10:00 AM	17:30 PM	f20cryo	cryotomo	sections	3		
Mon Mar  9 15:26:20 2009	eileen	3/9	8:30 AM	3:30 PM	70	rttomo	Gamma RNAi	6		
Tue Mar 10 17:15:47 2009	morphew	3-10	8:45 AM	5:15 PM	tiltrotate	rttomo	eric melamoma			
Wed Mar 11 19:45:00 2009	cindi	3/11	9:00 AM	7:45 PM	914	cryotomo	giardia cytoskeletons	6	woo hoo!!!!!!	
Thu Mar 12 18:30:16 2009	cedric-sarah	03/12	9:30 AM	6:30 PM	f20cryo	cryoimaging	tubes-tau1	0		
Fri Mar 13 18:19:35 2009	Matthew West	03/13	7:00 AM	6:30 PM	tiltrotate	rttomo	ER Mutant / Thick Section	18	Voeltz lab	
Mon Mar 16 14:50:16 2009	morphew	3-16	9 AM	3:45 PM	f30cryo	cryotomo	g127 plus BPV	0		gif doesn't seem to always  
continue tuning after needing to 
reset the beam intensity
Tue Mar 17 17:24:47 2009	eileen	3/17	8:30 AM	5:30 PM	70	rttomo	Gamma RNAi	10		
Wed Mar 18 11:26:56 2009	cindi	3/18	9:00 AM	11:30 AM	914	cryotomo	giardia cytoskeletons		cryo sucks.	
Mon Mar 23 14:53:35 2009	eileen	3/23	8:30 AM	3:00 PM	70	rttomo	Odde yeast prometa and maps	6		
Tue Mar 24 20:18:00 2009	Chris Arthur	03/24	9:00 AM	8:00 PM	914	cryotomo	Synaptosomes	9	Thanks Cindi	
Wed Mar 25 17:19:36 2009	Chris Arthur	03/25	9:00 AM	5:30 PM	914	cryotomo	Synaptosomes	6		
Wed Apr  1 17:59:57 2009	Johanna/Kristin	4/1	9:30 AM	6 PM	fishione	rttomo	Trypanosome flagella connector	2		
Thu Apr  2 17:17:06 2009	Matthew West	04/02	11:00 AM	5:15 PM	tiltrotate	rttomo	Plastic Yeast	10	Voeltz lab	crashed the vacuum putting the 
standard rod in at the last step 
before cryo-cycle
Fri Apr  3 14:59:28 2009	cedric	04/03	09:00 AM	15:00 PM	f20cryo	cryoimaging	section-plunge frozen cells	0		
Sun Apr  5 16:36:19 2009	cindi	04/05	9:00 AM	4:30 PM	standard	rtimage	giardia thin sxns			
Mon Apr  6 13:47:27 2009	morphew	4-06	9 AM	2 PM	f30cryo	cryotomo	c127-bpv	3		
Mon Apr  6 21:48:18 2009	Johanna/Kristin	04/06	2:00 PM	10:00 PM	fishione	rttomo	Trypanosomes	3		
Tue Apr  7 17:55:22 2009	eileen	4/7/09	9:15 AM	6:00 PM	70	rttomo	WT centrosomes	8		
Wed Apr  8 16:12:46 2009	morphew	4-8	8:45 AM	4:45 PM	f30cryo	cryotomo	bpv-c127	1		
Thu Apr  9 12:58:26 2009	eileen	4/9/09	9:30 AM	1:00 PM	70	rttomo	WT centrosome series	6		
Thu Apr  9 21:12:06 2009	Johanna/Kristin	4/9	1:00 PM	9:10 PM	fishione	rttomo	trypanosome	4		
Fri Apr 10 17:26:07 2009	cedric	04/10	09:00 AM	18:00 PM	914	cryoimaging	sections	0		
Mon Apr 13 18:52:08 2009	Matthew West	04/13	8:30 AM	7:00 PM	tiltrotate	rttomo	ER triple mutant	16	Voeltz lab tomography	
Tue Apr 14 17:40:54 2009	eileen	4/14	1:15 PM	5:45 PM	70	rttomo	endophilin DKO (De Camilli)	4		
Wed Apr 15 19:55:17 2009	Johanna/Kristin	04/15	1:00 PM	8:15 PM	fishione	rttomo	chlamydomonas	4		
Wed Apr 15 19:56:07 2009	Johanna/Kristin	04/15	1:00 PM	8:15 PM	fishione	rttomo	chlamydomonas	4		Stage lost communication with 
multifunction knobs and trackball.
Thu Apr 16 14:01:51 2009	morphew	04/16	8:45 AM	2 PM	tiltrotate	rttomo	bpv_c127	2-2x1		
Fri Apr 17 11:14:59 2009	eileen	4/17	8:15 AM	11:15 AM	70	rttomo	endophilin TKO (De Camilli)	4		
Mon Apr 20 20:19:59 2009	Kristin/Johanna	04/20	12:30 PM	8:30 PM	fishione	rttomo	trypanosomes	6		
Tue Apr 21 17:39:53 2009	eileen	4/21	8:45 AM	5:45 PM	70	rttomo	gamma RNAi late mitotic; endophilin 
wt control (De Camilli)	10		
Wed Apr 22 15:59:28 2009	morphew	4-23	9 AM	4 PM	f30cryo	cryotomo	bpv	0		
Wed Apr 22 20:52:34 2009	Johanna	04/22	4:00 PM	9:00 PM	fishione	rttomo	trypanosomes	4		
Thu Apr 23 15:50:56 2009	cedric	04/23	9:00 AM	17:00 PM	f30cryo	cryoimaging	metallothionein	0		
Tue Apr 28 15:32:57 2009	morphew	4-28	10:30 AM	3:30 PM	tiltrotate	rttomo	bpv	4		
Thu Apr 30 16:34:04 2009	cedric	04/30	9:00 AM	17:00 PM	f30cryo	cryoimaging	tau-mth2-gold	0		
Fri May  1 19:35:31 2009	Johanna	5/1	10:00 AM	7:30 PM	fishione	rttomo	trypanosomes	7		
Sat May  2 18:25:54 2009	Johanna	5/2	1:30 PM	6:30 PM	fishione	rttomo	chlamydomonas	6		
Sun May  3 19:20:21 2009	Johanna	5/3	12:30 PM	7:15 PM	fishione	rttomo	chlamydomonas/trypsanosoma	4		I had to insert the rod several 
times, cleaning it before and 
greasing the o-ring, since the 
vaccum stayed up. Still, the vaccum 
crashed when rod inserted the 
third/fourth time. After this, no 
more problems though. 
Mon May  4 21:07:24 2009	Johanna	5/4	1:00 PM	9:00 PM	fishione	rttomo	trypanosome	10		
Tue May  5 20:29:44 2009	johanna	5/5	2:00 PM	8:30 PM	fishione	rttomo	trypanosomes	6		
Wed May  6 18:30:31 2009	Julia	5/6	1:30 PM	6:27 PM	f30cryo	cryoimaging	Eg5A3 mts		All good	
Thu May  7 13:25:33 2009	morphew	05-07	7:45 AM	1:22 PM	f30cryo	cryotomo	mechanosensitive channels	4		
Thu May  7 15:27:00 2009	Cindi	05/07	2:15 PM	3:30 PM	standard	rtimage	giardia neg stain		samples sucked	
Fri May  8 16:28:21 2009	Ron Spilman	5/08	1:00  PM	5:00 PM	standard	serviceman	Gold 2.35A resolution	0	PM,  Made electronic and vacuum 
checks, backup hard drive, checked 
resolution, drain compressor,check 
hoses and fittings.	
Mon May 11 16:09:28 2009	Johanna	05/11	10:00 AM	4:00 PM	fishione	rttomo	tryps/chlamy	0		The vacuum was on 20 when I arrived 
and did not go down even after an 
hour with liquid nitrogen. 
I noticed that the compustage had 
not been reset. Coincidence or not? 
Tue May 12 17:01:44 2009	eileen	5/12	8:30 AM	5:00 PM	70	rttomo	DeCamilli Endophilin DKO and 
Dynamin TKO neurons	9		
Wed May 13 15:43:53 2009	Kristin/Cindi	5/13	1:00 PM	3:45 PM	tiltrotate	rttomo	septin delete pombe cells	1		T/R holder does NOT rotate. The 
motor makes no noise. Getting an 
RMA to send to Gatan for repair.
Thu May 14 17:36:22 2009	Janet Meehl/Odde	5/14/09	9:10 AM	5:45 PM	70	rttomo	yeast cik1del spindles	9		
Fri May 15 16:37:07 2009	morphew	5-15	9 AM	4:30 PM	f30cryo	cryotomo	mechanosensitive channel	4		
Mon May 18 18:25:48 2009	Kristin	5/18	12:30 PM	6:20 PM	fishione	rttomo	trypanosomes	4		I selected 'keep intensity below 
current value' but when I started 
the tilt series, the beam didn't 
spread so the image became 
overexposed... Had to perform tilt 
series without that 'keep 
intensity' option selected.
Tue May 19 19:51:32 2009	Matthew West	5/21	8:30 AM	8:00 PM	standard	rttomo	Yeast MVBs / ER	14	6 TS were for Odorizzi (5 hrs) and 
8 TS were for Voeltz (6.5 hrs)	
Wed May 20 17:11:18 2009	Janet Meehl/Odde	5/20/09	9:20 AM	5:10 PM	70	rttomo	yeast cik1delete spindles	12		
Fri May 22 14:06:34 2009	cedric	05/22	11:00 AM	2:00 PM	f30cryo	cryoimaging	unknown			
Wed May 27 17:35:21 2009	Kristin	5/27	11:30 AM	5:30 PM	fishione	rttomo	yeast	2		
Mon Jun  1 11:13:49 2009	eileen	5/28	12:30 PM	4:30 PM	70	rttomo	Chlamy 400nm wt	1		
Mon Jun  1 11:15:19 2009	eileen	5/29	8:30 AM	5:30 PM	70	rttomo	Chlamy 400nm wt	6		E: drive died during session. David 
tested the drive and is returning 
it.
Mon Jun  1 15:53:36 2009	eileen	6/1	1:15 PM	4:00 PM	70	rtimage	mapped out tbg1(RNAi) poles			
Wed Jun  3 09:28:33 2009	eileen	6/2/09	9:15 AM	5:15 PM	70	rttomo	tbg1(RNAi)	7 1x2; 1 2x2		
Wed Jun  3 12:31:45 2009	eileen	6/3/09	9:15 AM	12:30 PM	70	rttomo	tbg1(RNAi)	3 1x2		
Wed Jun  3 17:10:51 2009	Janet Meehl/Winey	6/3/09	12:30 PM	5:20 PM	70	rttomo	WT tet duplicating basal bodies	7		
Thu Jun  4 10:59:30 2009	eileen	6/4/09	8:30 AM	11:00 AM	70	rttomo	tbg1(RNAi) midzone	1 2x2		
Thu Jun  4 18:21:43 2009	Kristin	6/4	2:00 PM	6:30 PM	fishione	rttomo	pombe	2		
Mon Jun  8 16:12:16 2009	morphew	6/8/09	10 AM	4 PM	70	rttomo	low temp freeze sub	4	beam turned-off once, at random	
Tue Jun  9 14:22:54 2009	cindi/derek	06/09	8:00 AM	2:15 PM	f30cryo	cryotomo	telomeric complexes	2		
Thu Jun 11 08:43:41 2009	DNM	6/10	4:30 PM	6:00 PM	70	training	preparing for workshop demo			
Thu Jun 11 08:44:26 2009	DNM	6/11	6:15 AM	8:45 AM	70	training	demo to workshop in England			
Thu Jun 11 16:30:21 2009	Johanna	06/11	10:00 AM	5:00 PM	fishione	rttomo	S. pombe	2		
Fri Jun 12 16:29:45 2009	eileen	6/12/09	8:45 AM	4:30 PM	fishione	rttomo	tbg1(RNAi)	6 ts + 2 1x2	imaged 1x2 montages with +/- 70 rod 
then used fishione for other 6 ts	
Tue Jun 16 21:21:29 2009	Matthew West	06/16	7:30 AM	9:30 PM	fishione	rttomo	ER Mutant, Voeltz lab	18	Voeltz lab tomography--Scope is 
working great!  	+/- 60 standard rod is slightly bent
Wed Jun 17 16:48:10 2009	morphew	6/17	9 AM	5 PM	70	rttomo	low temp freeze sub	4		
Thu Jun 18 16:04:14 2009	Kristin	6/18	1:50 PM	4:00 PM	fishione	rttomo	trypanosomes	0		
Fri Jun 19 15:49:01 2009	cindi	06/19	9:15 AM	3:45 PM	standard	rtimage	giardia/pombe		samples sucked!	Found C2 aperture at 150um instead 
of 100um. put back to 100um
Thu Jun 25 16:49:03 2009	cindi/david	06/25	9:30 AM	4:30 PM	standard	serviceman	alignments			redid EFTEM alignments as well as 
all scope alignments. Use New062509 
as the alignment file
Fri Jun 26 14:57:07 2009	Cynthia	06/24	10:00 AM	4:30 PM	f20cryo	cryoimaging	ifs		Gif needed tuning - Cindi did new 
alignments on 6/25	
Sat Jun 27 15:44:02 2009	cynthia	06/27	10:00 AM	3:30 PM	f30cryo	cryoimaging	ifs		tried to use gif!	Hi Cindi, I was not success full 
using the gif today. I am not sure 
what the problem was but after 
tuning the gif I couldn't find a 
beam. Tried to redo alignments and 
repeat gif tuning but then DM 
didn't read beam even though I 
could see it... message intensity 
too low! Screen was up.
Tue Jun 30 22:20:04 2009	Matt West / Nesia Zurek	06/30	8:45 AM	10:30 PM	fishione	rttomo	ER of Sc at bud sites:  WT, ER 
mutant	18	Voeltz lab session	
Wed Jul  1 14:19:13 2009	eileen	7/1/09	8:30 AM	2:15 PM	fishione	rttomo	tbg1(RNAi) PNM	4		
Thu Jul  2 16:13:02 2009	morphew	7/2	10 AM	4 PM	fishione	rttomo	low temp freeze sub	4		
Wed Jul  8 16:52:57 2009	DNM	07/08	2:30 PM	5:00 PM	70	serviceman	Fixing low mag alignments			
Thu Jul  9 11:25:08 2009	cindi	07/09	9:15 AM	11:20 AM	standard	serviceman	realignments			
Thu Jul  9 18:59:32 2009	Kristin/Johanna	07/09	11:30 AM	7:30 PM	fishione	rttomo	anterior end trypansoma	4		
Fri Jul 10 15:52:58 2009	Johanna	07/10	10:00 AM	4:00 PM	fishione	rtimage	trypanosome telophase	2	montages	
Mon Jul 13 16:44:33 2009	Johanna	07/13	9:00 AM	4:30 PM	fishione	rtimage	trypanosome telophase	4	montages	
Tue Jul 14 19:51:04 2009	Matthew West	07/14	9:30 AM	8:00 PM	fishione	rttomo	MVBs of Yeast, Bro1 Domain 
Overexpression	16	Odorizzi lab	
Wed Jul 15 16:41:46 2009	cindi/chieze	04/15	9:00 AM	4:45 PM	70	rttomo	pombe, wt and septin delete	4		also deleted and made new cschwart 
account
Thu Jul 16 17:33:43 2009	cedric	07/16	10:00 AM	17:30 PM	tiltrotate	rttomo	polyoma-dextran-plastic sections	3		
Fri Jul 17 16:04:07 2009	morphew	7-17	9 AM	4 PM	70	rttomo	low temp freeze sub	2		
Mon Jul 20 10:54:32 2009	cindi	07/20	10:00 AM	10:45 AM	standard	rtimage	giardia hexadecene			column vacuum at 17 this morning
Mon Jul 20 19:10:47 2009	Johanna	07/20	11:00 AM	7:30 PM	fishione	rttomo	tryps telophase + leish 	5		Can still not transfer files to 
simba. 
Tue Jul 21 16:51:03 2009	morphew	7/23	1 PM	5 PM	70	rttomo	low temp freeze sub	1		
Wed Jul 22 16:35:19 2009	cedric	07/21	11:00 AM	16:30 PM	f30cryo	cryoimaging	pombe	0	this scope sucks anyhow. Darn it ...	
Thu Jul 23 16:35:15 2009	cedric	07/23	10:00 AM	16:30 PM	914	cryoimaging	pombe	0	I defenetly hate this microscope.
beam is always shifting. Objectif 
aperture always get in the way. it 
sucks	the dose calibration should some 
how be skrewed up because now to 
take a decent picture I should use 
more than 20 e/A2.
Fri Jul 24 15:57:40 2009	eileen	7/24	8 AM	4 PM	fishione	rttomo	wt centrosome	6		
Tue Jul 28 16:07:26 2009	philip/cindi	07/28	9:15 AM	4:00 PM	70	rttomo	aptamers on cells	9	training went well	
Wed Jul 29 16:12:13 2009	cindi	7/29	3:00 PM	4:15 PM	standard	rtimage	giardia x-sxns			
Thu Jul 30 14:08:40 2009	Kristin	07/30	9:00 AM	2:00 PM	fishione	rttomo	Trypanosomes	2		
Fri Jul 31 16:09:09 2009	cedric	07/31	09:00 AM	16:00 PM	f30cryo	cryoimaging	pombe	0	defenetly sucks :)))	
Wed Aug  5 18:52:49 2009	Matthew West	08/05	8;15 AM	7:00 PM	fishione	rttomo	ER mutant, Voeltz lab	14	Gia Voeltz lab strain	
Thu Aug  6 13:33:52 2009	morphew	8-6	9 AM	1:30 PM	f30cryo	cryoimaging	mt's in vivo	1		
Thu Aug  6 16:08:43 2009	Kristin	08/06	2:10 PM	4:05 PM	fishione	rtimage	pombe	0		
Mon Aug 10 18:58:11 2009	Nesia	09/10	9:00 AM	7:00 PM	tiltrotate	rttomo	HM20 embedded yeast 	12		
Tue Aug 11 16:08:13 2009	Nesia 	09/11	10 AM	4 PM	tiltrotate	rttomo	HM20 yeast 	8		
Wed Aug 12 15:39:03 2009	morphew	8/12	9 AM	3:30 PM	f30cryo	cryotomo	mechanosensitive channel	4		
Thu Aug 13 15:52:56 2009	Johanna	08/13	10:00 AM	4:00 PM	fishione	rttomo	leishmania	4		still can not transfer data without 
ssh-ing them. 
Fri Aug 14 12:17:18 2009	Johanna	08/14	08:30 AM	12:00 PM	fishione	rttomo	LEISHMANIA	6		
Mon Aug 17 12:13:43 2009	Janet Meehl/Winey	8/17	9:15 AM	12:15 PM	tiltrotate	rttomo	WTtet basal bodies, duplicating	2		
Tue Aug 18 17:33:27 2009	Janet Meehl/Winey	8/18	10:00 AM	5:45 PM	tiltrotate	rttomo	WTtet duplicating basal bodies	9		
Wed Aug 19 12:07:33 2009	Johanna	08/19	11:00 AM	12:00 PM	fishione	rttomo	leishmania	0	serial sections	
Wed Aug 19 18:30:12 2009	Janet Meehl/Odde	819	12:15 PM	6:45 PM	tiltrotate	rttomo	WTtetrahymena duplicating basal 
bodies	8		
Thu Aug 20 16:20:01 2009	Kristin	08/20	2:20 PM	5:00 PM	fishione	rttomo	pombe	2		
Fri Aug 21 18:44:37 2009	cedric	08/21	9:00 AM	19:00 PM	tiltrotate	rttomo	polyoma-38h-mannitol	8		
Tue Aug 25 16:41:16 2009	morphew	8-25	9 AM	5 PM	tiltrotate	rttomo	bpv	4		
Wed Aug 26 17:49:58 2009	Robert Kirmse	08/26	10:00 AM	5:50 PM	f30cryo	cryoimaging	CEMOVIS hDesWT	0		
Thu Aug 27 10:38:06 2009	Cindi	08/27	10:15 AM	10:30 AM	standard	serviceman	checking out GIF alignments			Checked out GIF alignments b/c 
Robert had issues yesterday. He had 
a dark corner all the time with all 
apertures and sample out. My 
account is fine.
Fri Aug 28 15:43:05 2009	cedric	08/28	10:00 AM	16:00 PM	f30cryo	cryoimaging	section-polyoma-38h	0		
Mon Aug 31 17:20:28 2009	Nesia Zurek	8/31	9:00 AM	5:30 PM	fishione	rttomo	HM20 embedded Yeast	12	besides resetting goniometer... it 
went perfectly	reset goniometer
Tue Sep  1 16:31:24 2009	morphew	9-1	9 AM	4:30 PM	f30cryo	cryotomo	mpp	2		
Thu Sep  3 17:30:04 2009	cedric	09/03	10:00 AM	17:30 PM	f30cryo	cryoimaging	yeast-section	0		
Fri Sep  4 15:54:03 2009	eileen	9/4	8:15 AM	4:00 PM	fishione	rttomo	DeCamili; endophilin TKO;Dynamin KO	6ts (4 2x1)	Please leave scope with standard 
apertures	
Tue Sep  8 17:35:39 2009	eileen	9/8	8:15 AM	5:30 PM	fishione	rttomo	DeCamili; dynaminKO_Kstimulation	10	vacuum took a while to come down 
after inserted rod. 	
Wed Sep  9 16:14:02 2009	cindi	09/09	9:00 AM	4:00 PM	tiltrotate	rttomo	giardia thick sxns			vacuum very slow coming down
Thu Sep 10 14:11:15 2009	DNM	9/10	11:30 AM	2:15 PM	70	rtimage	Distortion calubration for 13.5 and 
9.3K		Do not ever do this without the 
rotating rod	
Fri Sep 11 15:35:32 2009	eileen	9/11	8:30 AM	3:30 PM	fishione	rttomo	DeCamili dynamin KO recovery 1 min	8		
Sun Sep 13 15:27:38 2009	cindi	09/13	12:30 PM	3:30 PM	tiltrotate	rttomo	giardia serials	1		
Mon Sep 14 17:10:13 2009	eileen	9/14	8:15 AM	5:00 PM	fishione	rttomo	Platelets with Brian Storrie		vac stayed ~12 even though oring 
had new grease	
Wed Sep 16 14:42:54 2009	Nesia 	9/16/09	9:30 AM	3:00 PM	fishione	rttomo	Yeast in HM20	6		none
Mon Sep 21 17:13:32 2009	eileen	9/21	9 AM	5:10 PM	fishione	rttomo	Brian Storrie Platelets	3		
Tue Sep 22 19:22:25 2009	Matthew West	09/22	9:15 AM	7:25 PM	fishione	rttomo	Yeast ER / Endosomes		Voeltz time: 8 hrs
Odorizzi time:  2 hrs	
Wed Sep 23 16:58:13 2009	morphew	9-23	9:30 AM	9 PM	tiltrotate	rttomo	bpv	2		
Thu Sep 24 15:14:09 2009	morphew	9-26	9 AM	3:10 PM	tiltrotate	rttomo	bpv	4		
Fri Sep 25 15:39:19 2009	Nesia Zurek	9/25	9:30 AM	3:30 PM	fishione	rttomo	HM20 yeast 	6	good	
Mon Sep 28 16:26:27 2009	eileen	9/28	9 AM	4:20 PM	tiltrotate	rttomo	eyespot	4		
Tue Sep 29 17:17:06 2009	cedric	09/29	10:00 AM	18:00 PM	fishione	rttomo	polyoma-dextran	4	smooth	
Wed Sep 30 19:43:51 2009	cedric	09/30	10:00 AM	20:00 PM	fishione	rttomo	polyoma-dextran-36H	3	worked good	
Thu Oct  1 15:47:34 2009	Nesia	10/01	10 AM	4 PM	fishione	rttomo	HM20 embedded yeast	12	Cedric was still logged on when i 
came in, not a big deal though	none
Thu Oct  1 16:40:53 2009	Cindi	10/01	4:00 PM	4:40 PM	tiltrotate	rttomo	giardia 400nm sxns	1	just testing :o)	
Fri Oct  2 10:04:57 2009	Cindi	10/02	9:00 AM	10:00 AM	tiltrotate	rttomo	thick sections giardia	1	good	
Fri Oct  2 17:04:27 2009	morphew	10-2-09	10 AM	5 PM	tiltrotate	rttomo	marcus engstler tryps	4-3x1		
Mon Oct  5 12:50:23 2009	morphew	10-5	9 AM	12;45 PM	tiltrotate	rttomo	ENGSTLER TRYPS	2 3X1		
Mon Oct  5 19:55:11 2009	Matt Russell	Oct 5	1:30 PM	8:00 PM	tiltrotate	rttomo	yeast plastic	3		Vacuum IGP1 was increasing (up to 
~40) during tilt series.  Matt W 
helped me remove rod and clean o-
ring, and fill up cold finger 
dewar.  IGP1 vacuum allowed to 
return to 15.  Replaced rod, 
resumed tilt series; vacuum OK.
Tue Oct  6 12:58:35 2009	eileen	10/6	8:15 AM	12:50 PM	fishione	rttomo	EndTKODynKo DeCamilli	3 1x2		
Tue Oct  6 18:28:00 2009	Matt Russell	Oct 6	1:30 PM	6:30 PM	tiltrotate	rttomo	yeast plastic	2		
Wed Oct  7 09:55:17 2009	eileen	10/7	8:15 AM	9:45 AM	fishione	rttomo	EndophilinTKO/dynaminKO DeCamilli	1 1x2		
Wed Oct  7 17:20:34 2009	cedric	10\07	10:00 AM	17:30 PM	fishione	rttomo	polyoma-dextran	12		
Thu Oct  8 16:13:28 2009	Kristin	10/08	12:30 PM	4:15 PM	fishione	rttomo	pombe	2		
Fri Oct  9 13:27:14 2009	eileen	10/8	8:15 AM	1:20 PM	fishione	rttomo	Endophilin TKO DeCamilli	3	o-ring on fishione was bad- this 
has been causing vac problems. 
Column dumped this am. Ron replaced 
and now no vac problems. 	
Fri Oct  9 18:51:33 2009	Matt Russell	10/09	2:00 PM	6:45 PM	tiltrotate	rttomo	yeast plastic	2		
Mon Oct 12 18:49:40 2009	Janet Meehl/Winey	10/12	12:30 PM	6:50 PM	tiltrotate	rttomo	WTtet assembling basal bodies	6		
Tue Oct 13 17:17:12 2009	eileen	10/13	8:15 AM	5:15 PM	fishione	rttomo	Odde MTs and gamma RNAi	9		
Wed Oct 14 16:48:16 2009	cynthia	10/14	9:30 AM	4:45 PM	f30cryo	cryoimaging	ifs		used the gif	
Thu Oct 15 14:06:12 2009	Kristin	10/15	9:30 AM	2:00 PM	fishione	rttomo	pombe	2		
Fri Oct 16 10:00:53 2009	Janet Meehl Fox/Winey	10/16	8:45 AM	10:00 AM	tiltrotate	rttomo	WT tet basal bodies	0	formvar on grid broke	
Fri Oct 16 15:51:19 2009	eileen	10/15	9:15 AM	3:45 PM	fishione	rttomo	Odde MTs and gamma RNAi	6	Please leave microscope in standard 
condition	
Mon Oct 19 17:04:23 2009	Robert Kirmse	10/19	10 AM	16 PM	f30cryo	cryoimaging	Desmin Tetramers	0	Scope was good
samples were shit	
Wed Oct 21 17:43:48 2009	Kristin	10/21	1:00 PM	6:00 PM	fishione	rttomo	pombe	2		
Thu Oct 22 19:28:57 2009	cynthia	10/21	9:30 AM	7:30 PM	f30cryo	cryotomo	ifs	2	worked well! 	Hi Cindi, only problem here, I saw 
the upper right quarter panel in my 
images.
Fri Oct 23 16:15:12 2009	eileen	10/23	8:15 AM	4:10 PM	fishione	rttomo	Odde LLC-PK1	5 (2x2 and 1x1)		
Mon Oct 26 17:21:17 2009	morphew	10-26	9 AM	5:30 PM	f30cryo	cryotomo	bpv	5		dumped column at rod removal--again!
Tue Oct 27 16:28:41 2009	eileen	10/27	8:10 AM	1:15 PM	tiltrotate	rttomo	Odde LLC-PK1	3 2x2's		
Tue Oct 27 16:42:37 2009	morphew	10-27	1 PM	5 PM	tiltrotate	rttomo	hela midbody	2 2x1		
Wed Oct 28 14:46:15 2009	cynthia	10/28	9:15 AM	2:45 PM	f30cryo	cryoimaging	ifs		except for the occasional upper 
right quarter panel, aok!	
Thu Oct 29 18:08:44 2009	Johanna/Neil	10/29	8:30 AM	6:30 PM	fishione	rttomo	trypanosome collar	10	everything perfect!	
Fri Oct 30 19:02:31 2009	Johanna/Neil	10/30	08:30 AM	7:30 PM	fishione	rttomo	trypanosome collar	10		
Sat Oct 31 17:26:21 2009	Johanna/Neil	10/31	9:00 AM	5:00 PM	fishione	rttomo	Trypanosome collar	8		
Sun Nov  1 19:16:02 2009	Johanna/Neil	11/1	09:00 AM	7:15 PM	fishione	rttomo	Trypanosome collar	9		
Mon Nov  2 12:02:18 2009	Kristin	11/2	8:30 AM	12:00 PM	fishione	rttomo	pombe	2		
Mon Nov  2 21:09:29 2009	Johanna/Neil	11/2	12:30 PM	9:15 PM	fishione	rttomo	trypanosome collar	10		
Tue Nov  3 12:36:21 2009	Johanna/Neil	11/3	8:00 AM	12:30 PM	fishione	rttomo	trypanosome collar	7		
Wed Nov  4 09:57:18 2009	sRon	11/04	8:00 AM	10:00 AM	standard	serviceman	HT off			refilled water chiller. brought 
scope back up. reset goniometer
Wed Nov  4 10:08:04 2009	Kristin	11/3	12:30 PM	4:30 PM	tiltrotate	rttomo	HeLa 	1		
Wed Nov  4 15:57:08 2009	eileen	11/4	10:30 AM	3:50 PM	tiltrotate	rttomo	Odde MTs	2 2x2	microscope vac down when I came 
in.Ron got microscope up and 
running.	
Thu Nov  5 17:57:08 2009	Janet Meehl Fox/Winey	11/5	8:30 AM	6:00 PM	tiltrotate	rttomo	poc1del tet and WT tet dupl. bbs	10		I don't know if it was the specimen 
rod or the stage motor that was 
randomly clicking on my last TS.  I 
had to wiggle the specimen rod to 
get it to move back to center.  It 
seems quiet with the std. rod in.  
My last TS also seized up a couple 
of times which I think was the 
specimen rod/stage issue.
Fri Nov  6 13:57:08 2009	cindi	11/06	9:30 AM	11:45 AM	f30cryo	cryoimaging	giardia		samples too thick	
Fri Nov  6 15:28:41 2009	cindi	11/06	11:45 AM	3:30 PM	tiltrotate	rttomo	giardia dextran beads	2	cells looked good!	had to reload alignments--weird 
rotation in LMM. also, accidentally 
unplugged computer, so it got 
rebooted.
Sun Nov  8 18:35:40 2009	DNM	11/8	2:30 PM	6:30 PM	standard	development	Testing drift compensation with 
image shift			
Mon Nov  9 15:14:48 2009	morphew	11-9	9 AM	3 PM	tiltrotate	rttomo	bpv	2		
Mon Nov  9 17:12:48 2009	Johanna/Neil	9/11	15:15 PM	17:15 PM	fishione	rttomo	trypanosome collar	2		
Mon Nov  9 18:33:33 2009	DNM	11/9	5:30 PM	6:30 PM	standard	development	Drift compensation with image shift			
Tue Nov 10 16:44:27 2009	cindi	11/10	9:15 AM	4:45 PM	tiltrotate	rttomo	giardia	4	some montaging tests	
Wed Nov 11 15:12:12 2009	Janet Meehl Fox/Winey	11/11	9:00 AM	3:20 PM	tiltrotate	rttomo	WTtet duplicating basal bodies	8		
Thu Nov 12 12:33:07 2009	Johanna	11/12	9:15 AM	12:30 PM	fishione	rttomo	tryps	3		
Thu Nov 12 17:26:43 2009	Kristin/John	11/12	12:30 PM	5:30 PM	tiltrotate	rttomo	hela telophase	2	montages	
Thu Nov 12 18:39:30 2009	DNM	11/12	5:30 PM	6:45 PM	standard	development	Recording shifting image during 
exposure on GIF video camera		Yes, there is vibration in the GIF	
Fri Nov 13 15:56:32 2009	DNM	11/13	3:25 PM	4:00 PM	standard	development	more drift image shift measurements			
Fri Nov 13 18:26:39 2009	Johanna	11/13	4:00 PM	6:30 PM	fishione	rttomo	trypanosome	3		
Mon Nov 16 17:34:02 2009	Matthew West	11/16	9:00 AM	5:30 PM	tiltrotate	rttomo	Yeast endosomes, Bro1 oe	10	This work was for Greg Odorizzi's 
lab.	
Tue Nov 17 18:00:26 2009	eileen	11/17	8:30 AM	6:00 PM	fishione	rttomo	C.elegans SPD5 RNAi	10		
Wed Nov 18 17:51:04 2009	Robert Kirmse	11/18	9:30 AM	6:00 PM	f30cryo	cryoimaging	desmin	0	no problems	
Thu Nov 19 15:39:36 2009	morphew	11-19	8:30 AM	3:30 PM	f30cryo	cryotomo	s9 cells	0		
Fri Nov 20 16:19:33 2009	cynthia	11/19	9:30 AM	4:00 PM	f30cryo	cryoimaging	if		Seems to be a bug in Refine ZLP 
according to David	
Mon Nov 23 16:56:21 2009	cynthia	11/23	10:00 AM	5:00 PM	f30cryo	cryoimaging	ifs		very good!	
Tue Nov 24 18:03:39 2009	eileen	11/24	9:00 AM	6:00 PM	fishione	rttomo	SPD5 RNAi and Dyn1KO recovery 	8		
Wed Nov 25 15:55:44 2009	morphew	11-26-09	8:30 AM	4 PM	f30cryo	cryotomo	c127 bpv	4		
Sat Nov 28 14:28:53 2009	Johanna	11/28	9:00 AM	2:30 PM	fishione	rttomo	trypanosome collar	6		
Tue Dec  1 16:34:55 2009	Johanna	12/1	11:00 AM	4:30 PM	fishione	rttomo	trypanosome fusion	6	left myself logged in so that the 
files could transfer. 	
Thu Dec  3 21:09:39 2009	cedric	12/03	10:00 AM	21:30 PM	tiltrotate	rttomo	polyoma	10		
Fri Dec  4 15:49:43 2009	cindi	12/04	9:00 AM	3:45 PM	tiltrotate	rttomo	Giardia sections Low Temp	6		
Mon Dec  7 14:30:31 2009	morphew	12-7	9 AM	12 AM	f30cryo	cryotomo	s9	0		
Mon Dec  7 18:00:23 2009	Johanna	12/7	2:30 PM	6:00 PM	fishione	rttomo	trypanosome fusion	2		
Tue Dec  8 17:28:49 2009	Matthew West	12/08	10:15 AM	5:45 PM	tiltrotate	rttomo	Bro1p over-expressed yeast	8	Lower than normal beam energy for 
spot size 3 so I used spot size 2, 
otherwise good alignments.	
Wed Dec  9 18:30:00 2009	Kirmse Robert	12/09	09:00 AM	7:00 PM	f30cryo	cryoimaging	cryo negativ staining test	0	All good during session,
crahed vacuum, while reinserting
RT-holder	
Thu Dec 10 13:08:02 2009	Johanna	12/10	9:30 AM	1:00 PM	fishione	rttomo	tryps fusion	1		
Thu Dec 10 16:22:51 2009	Kristin	12/10	1:00 PM	5:00 PM	tiltrotate	rttomo	HeLa	2		
Fri Dec 11 15:24:29 2009	cedric	12/11	9:30 AM	16:00 PM	tiltrotate	rttomo	polyoma virus	1		
Mon Dec 14 18:21:43 2009	Janet Meehl Fox/Winey	12/14	9:00 AM	6:30 PM	tiltrotate	rttomo	Tet poc1del bbs	12		
Tue Dec 15 11:31:54 2009	eileen	12/15	8:50 AM	11:30 AM	fishione	rttomo	spd5 RNAi	2		
Tue Dec 15 17:35:58 2009	cindi	12/15	12:00 PM	5:30 PM	tiltrotate	rttomo	giardia	6		
Wed Dec 16 17:14:40 2009	Kirmse Robert	12/16	9:30 AM	5:30 PM	f30cryo	cryoimaging	Desmin	0	all good	
Thu Dec 17 16:43:17 2009	Kristin	12/17	9:00 AM	5:00 PM	tiltrotate	rttomo	HeLa 	6		
Fri Dec 18 16:35:16 2009	eileen	12/18	9:15 AM	4:30 PM	fishione	rttomo	spd5 RNAi	2		
Mon Dec 21 18:08:18 2009	Matthew West	12/21	8:30 AM	6:30 PM	tiltrotate	rttomo	WT & Bro1 over expression	14	Odorizzi lab	
Tue Dec 22 16:26:58 2009	cindi	12/22	10:00 AM	4:30 PM	tiltrotate	rttomo	giardia	5		
Sun Dec 27 15:24:55 2009	Cindi	12/27	10:00 AM	3:00 PM	tiltrotate	rttomo	giardia	5		The e: drive is sounding very 
strange. Cleaned out the drive and 
Dave is going to send it back for 
repair.
Mon Dec 28 16:58:42 2009	cindi	12/28	9:00 AM	5:00 PM	tiltrotate	rttomo	Giardia	6		camera vacuum high. chuck to come 
fix this wednesday.
Tue Dec 29 14:46:51 2009	eileen	12/29	9:30 AM	2:40 PM	fishione	rttomo	Chlamy bb duplication	6		
Tue Dec 29 16:15:47 2009	cindi	12/29	3:00 PM	4:00 PM	tiltrotate	rttomo	giardia	1		
Wed Dec 30 12:56:07 2009	Chuck Moe	12/30	9:00 AM	11:00 AM	standard	serviceman	fix camera vacuum			cleaned seals/orings. Penning guage 
bad. Ordered penning guage, to be 
replaced on Monday.
Wed Dec 30 15:56:06 2009	Cindi	12/30	2:00 PM	4:00 PM	tiltrotate	rttomo	Giardia	2		Had to wait for cameras to cool 
down. Had to shut down ultracam and 
restart. Chuck also lubed the 
binocs.
Thu Dec 31 10:51:19 2009	Cindi	12/31	8:00 AM	10:45 AM	tiltrotate	rttomo	Giardia	3		
Thu Dec 31 16:38:54 2009	Kristin	12/31	11:00 AM	4:45 PM	tiltrotate	rttomo	HeLa	4		
Tue Jan  5 17:23:38 2010	eileen	1/5	8:40 AM	5:20 PM	fishione	rttomo	chlamy bb, spd5 rnai, llcpk1 mts	4 + 3 2x2	put the scope in standby mode after 
session for power outage	
Wed Jan  6 14:23:21 2010	Ron Spilman	1/6	8:40 AM	5:00 PM	standard	serviceman	Preventive maintenance			Replace buffer hose, prevac hose, 
vac oil.  Replace red soop lamp.  
Drain water from Jun-Air.  Made 
electronic and vacuum checks as per 
PM check list.
Thu Jan  7 15:03:25 2010	Kristin	1/7	9:00 AM	3:15 PM	tiltrotate	rtimage	trypanosomes	0		
Mon Jan 11 18:59:19 2010	Matthew West	01/11	10:30 AM	7:00 PM	tiltrotate	rttomo	Snf7 point mutants and WT yeast	12		
Tue Jan 12 13:32:37 2010	eileen	1/12	8:30 AM	1:30 PM	fishione	rttomo	WT gamma-GFP controls	5		
Wed Jan 13 13:21:09 2010	morphew	1-13-10	9 AM	1 PM	tiltrotate	rttomo	low temp freeze sub	4		
Wed Jan 13 16:56:02 2010	Kristin	1/13	1:30 PM	5:30 PM	tiltrotate	rttomo	trypanosomes	2		
Thu Jan 14 17:30:26 2010	Kirmse Robert	01/14	9:00 AM	5:30 PM	f20cryo	cryoimaging	cryo negative stain 	0		
Tue Jan 19 17:27:46 2010	eileen	1/19	8:50 AM	5:30 PM	fishione	rttomo	spd5 RNAi and Odde mts	6 ts + 2 2x2 ts		
Wed Jan 20 18:12:08 2010	Kristin	1/20	7:50 AM	6:30 PM	tiltrotate	rttomo	trypanosomes	8		
Fri Jan 22 10:02:12 2010	Ron Spilman	1/22	8:30 AM	10:15 AM	standard	serviceman	Service (Replace PV pump)		
Replace prevac pump	pump was slow starting ok now
Fri Jan 22 17:36:00 2010	cedric	01/22	10:30 AM	17:30 PM	f30cryo	cryoimaging	sections	0		
Mon Jan 25 18:22:22 2010	Johanna	01/25	11:00 AM	6:30 PM	fishione	rttomo	trypanosomes	5	montages	
Tue Jan 26 11:34:44 2010	Kristin	1/26	8:30 AM	11:30 AM	tiltrotate	rtimage	hela	0		
Tue Jan 26 17:56:00 2010	Johanna	01/26	01:00 PM	6:00 PM	fishione	rttomo	trypanosomes	5	2 x1 montages	
Wed Jan 27 16:50:16 2010	morphew	1-27	8:30 AM	5 PM	f30cryo	cryotomo	polyoma	3		I don't think the cryobox is 
working properly.
Thu Jan 28 16:55:55 2010	Kristin	1/28	8:30 AM	5:00 PM	tiltrotate	rttomo	hela	6		
Fri Jan 29 11:44:25 2010	eileen	1/29	8:15 AM	11:40 AM	fishione	rttomo	De Camilli WT Recov. 1 min	4		
Fri Jan 29 15:58:49 2010	morphew	1-29	12 PM	4 PM	fishione	rttomo	polymom virus	3		
Mon Feb  1 15:58:11 2010	morphew	2-1-10	9 AM	4 PM	fishione	rttomo	polyoma	4 2x1		
Tue Feb  2 14:45:54 2010	Kristin	2/2	8:15 AM	2:45 PM	tiltrotate	rttomo	hela	3		crashed vacuum, computer lost 
communication with scope
Wed Feb  3 19:41:22 2010	Johanna	2/4	10:00 AM	7:45 PM	fishione	rttomo	trypanosome collar	2		
Fri Feb  5 15:02:12 2010	ghoshn	02/05	1:45 PM	3:00 PM	standard	training	training			
Fri Feb  5 15:29:13 2010	cschwart	02/05	3:00 PM	3:30 PM	70	rtimage	checking grids, they sucked			
Mon Feb  8 14:37:41 2010	ghoshn	02/08	1:00 PM	2:32 PM	standard	rtimage	training			
Tue Feb  9 15:59:25 2010	Kristin	2/9	11:30 AM	4:30 PM	tiltrotate	rttomo	hela	2		
Wed Feb 10 17:16:43 2010	Kristin	2/10	8:30 AM	5:30 PM	tiltrotate	rttomo	hela	4		
Thu Feb 11 12:51:09 2010	Kristin	2/11	9:00 AM	1:00 PM	tiltrotate	rttomo	hela	2		
Fri Feb 12 15:11:56 2010	cedric	02/12	12:30 PM	15:30 PM	f20cryo	cryoimaging	sections	0		
Mon Feb 15 13:50:30 2010	morphew	2-15	8:15 AM	1:45 PM	f30cryo	cryotomo	mpp	1		
Tue Feb 16 18:09:44 2010	eileen	2/16	8:50 AM	6:00 PM	tiltrotate	rttomo	C.elegans spd5; chlamy flagella	10		
Wed Feb 17 17:03:37 2010	eileen	2/17	8:20 AM	5:00 PM	tiltrotate	rttomo	chlamy flagella	4		
Thu Feb 18 17:12:05 2010	cindi	02/18	9:00 AM	5:00 PM	f30cryo	cryotomo	isolated giardia cytoskeletons	7		
Fri Feb 19 17:09:24 2010	Robert Kirmse	2/19	9:30 AM	5:00 PM	f30cryo	cryoimaging	Desmin	0		
Mon Feb 22 14:22:28 2010	Nesia Zurek	2/22	10 AM	2 PM	fishione	rttomo	yeast HM20	4	no	none
Mon Feb 22 19:40:45 2010	Matthew West	02/22	3:00 PM	7:45 PM	tiltrotate	rttomo	Voeltz lab, ER mutant	6	Voeltz lab	Serial EM was a little strange; 
worked better after restarting 
program
Tue Feb 23 16:44:14 2010	eileen	2/23	8:30 AM	4:40 PM	tiltrotate	rttomo	Odde Ase1 spindles and chlamy 
flagella maps	8		extraction voltage was below 4000
Wed Feb 24 21:39:22 2010	Cindi/Moises	02/24	12:00 PM	2:00 PM	f30cryo	cryoimaging	giardia cytos			
Wed Feb 24 21:39:45 2010	Cindi/Moises	02/24	2:00 PM	9:30 PM	tiltrotate	rttomo	giardia	5		
Thu Feb 25 17:58:53 2010	Cindi/Moises	02/25	10:20 AM	6:00 PM	tiltrotate	rttomo	Giardia	9		
Fri Feb 26 17:50:05 2010	cindi/Moises	02/26	9:00 AM	3:00 PM	f30cryo	cryoimaging	giardia cytos		samples sucked	
Fri Feb 26 17:50:24 2010	cindi/Moises	02/26	3:00 PM	5:45 PM	tiltrotate	rtimage	giardia			
Sat Feb 27 17:37:13 2010	cindi/moises	02/27	10:00 AM	5:30 PM	f30cryo	cryotomo	giardia cytos	2		
Mon Mar  1 16:43:37 2010	Cindi/Moises	03/01	9:00 AM	11:30 AM	tiltrotate	rttomo	giardia	2		
Mon Mar  1 16:43:54 2010	Cindi/Moises	03/01	11:30 AM	4:30 PM	f30cryo	cryoimaging	giardia cytos		samples too thick	
Mon Mar  1 16:48:03 2010	Cindi	03/01	4:45 PM	4:46 PM	standard	serviceman	Warm-up UltraCam			Warmed up Ultra-Cam overnight.
Tue Mar  2 16:34:08 2010	eileen	3/2	8:30 AM	4:30 PM	tiltrotate	rttomo	Odde Ase1 yeast	10		
Wed Mar  3 11:37:00 2010	cindi	03/03	9:00 AM	11:30 AM	standard	serviceman	quadrant correction UltraCam			Quadrant correction UltraCam
Wed Mar  3 21:47:59 2010	cindi/moises	03/03	12:30 PM	9:45 PM	f20cryo	cryotomo	giardia cytos	6		
Thu Mar  4 18:59:14 2010	cindi/moises	03/04	2:00 PM	7:00 PM	f20cryo	cryotomo	giardia WT cytos	4		
Fri Mar  5 12:37:35 2010	morphew	3/5	9 AM	12:30 PM	tiltrotate	rttomo	polyoma virus	2		
Fri Mar  5 17:15:26 2010	cindi/moises	03/05	2:30 PM	4:30 PM	tiltrotate	rttomo	giardia chem fix	2		
Tue Mar  9 17:37:50 2010	eileen	3/9	8:30 AM	5:15 PM	tiltrotate	rttomo	chlamy flagella	4		
Thu Mar 11 15:00:53 2010	DNM	3/11	2:35 PM	3:00 PM	standard	rtimage	looking at aperture sizes, getting 
carbon film pictures			
Fri Mar 12 15:01:02 2010	morphew	3/13	8:30 AM	3 PM	tiltrotate	rttomo	aditi neurons	0		
Mon Mar 15 17:06:13 2010	Cindi	03/15	2:30 PM	5:00 PM	standard	training	making maps		writing paper on how to make 
various maps	
Tue Mar 16 12:32:16 2010	Robert Kirmse	03/16	10:00 AM	1:00 PM	914	cryoimaging	Vitrobot test	0		
Wed Mar 17 14:06:13 2010	DNM	3/17	2:00 PM	2:15 PM	standard	rtimage	New noise files for ultracam			
Thu Mar 18 15:38:32 2010	cindi	03/18	10:00 AM	1:00 PM	standard	serviceman	vacuum bad			vacuum was at 30 at logon. taking 
holder in and out didn't help. 
called service. suggested to 
replace o-ring and leave holder out 
for awhile. vacuum slowly started 
dropping. put holder in/out a few 
times and vacuum stayed at 6 all 
day. chuck to come by in the 
morning and double-check.
Fri Mar 19 09:08:56 2010	Chuck Moe	3/19	08:30 AM	09:30 AM	standard	serviceman	checked stage and column vacuum		checked stage and column vacuum	found no vacuum problems
Mon Mar 22 16:25:17 2010	Workshop	03/22	10:00 AM	4:15 PM	914	training	Workshop demo			
Fri Mar 26 14:28:06 2010	Cindi	03/26	9:00 AM	2:00 PM	f20cryo	cryotomo	Giardia	4		seems like the camera vacuum may be 
spiking when the slit goes in/out.
Mon Mar 29 14:57:30 2010	spilman	3/29	8:00 AM	17:00 PM	standard	serviceman	service		vacuum issues after weekend.  
Appears to be V5.
Replace 100u C2 aperture.	
Tue Mar 30 14:31:49 2010	Spilman	3/30	8:00 AM	2:30 PM	standard	serviceman	service		vacuum issues with single tilt 
holder.	
Tue Mar 30 16:54:41 2010	cindi	03/30	3:00 PM	5:00 PM	tiltrotate	rttomo	Giardia	2		
Wed Mar 31 08:33:34 2010	Matthew West	3/24	9:00 AM	7:00 PM	tiltrotate	rttomo	ER triple delta	12	The camera insertion trick works!	The cryo box was still in 
preventing me from acquiring the 
beam and delaying my work.  Cindi 
pointed it out and I should be able 
to fix future cryo box problems
Wed Mar 31 11:24:36 2010	morphew	03/31/10	8:30 AM	11:30 AM	tiltrotate	rttomo	aditi neurons	2 1x2		
Thu Apr  1 14:33:26 2010	Kristin	4/1	8:00 AM	2:30 PM	tiltrotate	rttomo	hela	3		
Mon Apr  5 14:50:40 2010	Kristin	4/5	8:00 AM	3:00 PM	tiltrotate	rttomo	hela	4		
Tue Apr  6 16:22:51 2010	eileen	4/6	8:45 AM	4:20 PM	tiltrotate	development	supermontage test runs with David	1 3x3 plus test runs		
Thu Apr  8 17:43:35 2010	Johanna/Cindi	04/08	09:00 AM	5:30 PM	f30cryo	cryotomo	Trypanosomes whole cells	2		Vaccum was at 24 in the morning. 
Fri Apr  9 16:40:40 2010	eileen	4/9/10	8:00 AM	4:40 PM	70	rttomo	2x2 supermontage (3x3 panels) auto 
acquisiton with macro- YEA!!!!	4 3x3 montage	vac was at 6 this morning. 	
Mon Apr 12 08:09:56 2010	Spilman	4/12	8:00 AM	8:15 AM	standard	serviceman	none		check vacuum	vacuum at log 6
Mon Apr 12 16:51:31 2010	eileen	4/12	9 AM	4:45 PM	tiltrotate	rttomo	supermontage aquisition with 
Needleman	2 3x3	column vac at 6 after the weekend!	
Tue Apr 13 12:38:34 2010	Nesia	4/13	10 AM	12:30 PM	tiltrotate	rttomo	HM20 Yeast 	1		trouble engaging rod in airlock 
Wed Apr 14 11:18:34 2010	mo0rphew	4-13	8:30 AM	11:30 AM	tiltrotate	rttomo	aditi neurons	2 2x1		
Thu Apr 15 21:09:38 2010	Johanna	15/4	9:30 AM	9:30 PM	fishione	rttomo	Trypanosome collar	10		At one point, an x-ray warning came 
up saying it was not safe. It 
disappeared by itself. 
Mon Apr 19 16:20:52 2010	eileen	4/19	11:00 AM	4:15 PM	fishione	rttomo	DynaminKO 30min recovery (DeCamilli)	6		
Tue Apr 20 16:36:25 2010	Johanna/Cindi	04/20	9:30 AM	16:30 PM	f30cryo	cryotomo	trypanosomes whole mount/lysed	1		
Wed Apr 21 16:15:19 2010	morphew	4-21	8:15 AM	4:15 PM	tiltrotate	rttomo	polyoma, aditi neurons	4		
Thu Apr 22 12:56:43 2010	Nesia	4/22	9:30 AM	1 PM	tiltrotate	rttomo	HM20 yeast	4		
Thu Apr 22 16:38:13 2010	Matt Russell	04/22	1:30 PM	4:30 PM	tiltrotate	rtimage	Practice, checking sample			
Fri Apr 23 15:08:47 2010	Cindi	04/23	9:00 AM	3:00 PM	standard	rtimage	EELS		warmed upper camera	ZLP not centered between imaging 
and spectroscopy mode. Need ALOT of 
dose for EFTEM imaging. Beam moving 
as eV changes. Sent email to Dan 
Ray from Gatan. Warmed upper camera
Mon Apr 26 18:54:25 2010	Johanna	04/26	9:30 AM	7:00 PM	f30cryo	cryotomo	trypanosome detergent extracted 
flagella	8	bad vaccum in the morning. 	
Tue Apr 27 16:37:20 2010	eileen	4/27	8:30 AM	4:30 PM	tiltrotate	rttomo	spindle supermontage and synapses 
(decamili)	1x2 supermontage (3x3s) and 3 other ts	vac at 6 this morning.	
Wed Apr 28 17:26:12 2010	Johanna	04/28	9:00 AM	5:30 PM	f30cryo	cryotomo	whole cell tryps	3		
Thu Apr 29 18:06:17 2010	Kirmse Robert	04/29	10:00 AM	6:00 PM	f30cryo	cryotomo	desmin	1	put the camera in before opening 
the c-valves but in my case the 
vaccum didn't spike and stayed at 6
	
Fri Apr 30 12:48:54 2010	spilman	4/30	8:00 AM	1:00 PM	standard	serviceman	Vibration issues		Resolved issue	
Mon May  3 12:49:42 2010	eileen	5/3	8:30 AM	12:30 PM	fishione	rttomo	WT neurons 30 min recovery 
(DeCamili)	3	column vac at 32 when I came in. 	Hey Cindi-
column vac was at 32 when I came 
in. Took ~ hour to come down. Also, 
no e drive available. Saved to 'my 
document' folder.
Mon May  3 13:04:23 2010	cindi	05/03	1:00 PM	1:01 PM	standard	serviceman	turned on/off external drive			external drive was not recognized. 
turned off, then on and now 
recognized
Wed May  5 16:11:16 2010	Robert Kirmse	05/03	9:15 AM	04:10 PM	f30cryo	cryoimaging	Desmin	0		Vaccum was at 26 in the morning
Fri May  7 20:50:21 2010	cedric	05/07	10:00 AM	21:00 PM	f30cryo	cryotomo	section	4		
Mon May 10 14:51:24 2010	Spilman	5/10	8:00 AM	3:00 PM	standard	serviceman	service	0	Replaced V5	Vacuum deteriates over long period 
of time after cryo cycle when turbo 
is off
Tue May 11 16:07:20 2010	morphew	5-11	1:30 PM	4:15 PM	tiltrotate	rttomo	heart muscle	2		
Wed May 12 15:43:12 2010	morphew	5-13	8:30 AM	3:30 PM	tiltrotate	rttomo	cardiac myocytes	3		
Thu May 13 11:52:47 2010	morphew	5-13	9 AM	12 AM	tiltrotate	rttomo	cardiac myocytes	2 2x1		
Thu May 13 15:29:42 2010	Robert Kirmse	05/13	1:00 PM	4:00 PM	f30cryo	cryoimaging	Desmin	0		
Fri May 14 15:41:51 2010	morphew	5-14	9 AM	4 PM	tiltrotate	rttomo	cardiac myocyte	4		
Mon May 17 15:36:32 2010	morphew	5-17	8:30 AM	3:30 PM	tiltrotate	rttomo	cardiac myocyte	4 2x1		
Wed May 19 16:50:56 2010	Kristin	5/19	1:30 PM	4:45 PM	standard	rtimage	neurons	0		
Thu May 20 12:33:07 2010	Janet Meehl Fox/Winey	5/20/10	9:30 AM	12:30 PM	tiltrotate	rttomo	poc1del, WT  tet basal bodies	2		
Fri May 21 16:40:57 2010	Robert Kirmse	05/21	9:00 AM	5:00 PM	f30cryo	cryoimaging	Desmin	0		
Mon May 24 16:55:12 2010	eileen	5/24	8:30 AM	5:00 PM	tiltrotate	rttomo	Chlamy flagella/bbs	8		
Tue May 25 14:24:46 2010	Kristin	5/25	8:45 AM	2:20 PM	tiltrotate	rttomo	neurons	6		
Wed May 26 16:18:41 2010	cedric	05/26	14:00 PM	16:30 PM	f30cryo	cryoimaging	sections	0		
Thu May 27 13:39:37 2010	Spilman	5/27	9:00 AM	2:00 PM	standard	serviceman	none		Replace V15 & IGP4	
Fri May 28 18:01:56 2010	cedric	05/28	09:00 AM	18:00 PM	f30cryo	cryoimaging	cryosection	0	do not let Ron fix the microscope 
the day before you leave the lab.	lots of them
1. software stuck on cryocycle the 
whole night. IGP1 was at 99 in the 
morning. had to go to supervisor 
mode with david to restart the 
whole vaccum process twice.
2. the upper camera was not cooled 
down either. had to reboot the 
camera controler to cool it down.
3. tracking ball was set on 
joystick and beam moving all other 
the place without being able to 
stop it.
4. bottom left quadrond (I cannot 
remember how to write it) of the 
ultracam dead black. impossible to 
use the gif until fixed.
in summary: a day in hell...
Wed Jun  2 09:11:09 2010	Chuck Robertson	05/19	9:00 AM	12:30 PM	fishione	rttomo	S. solfataricus	1		
Wed Jun  2 09:11:45 2010	Chuck Robertson	05/26	8:30 AM	11:40 AM	fishione	rttomo	S. solfataricus	1		
Wed Jun  2 12:39:20 2010	Chuck Robertson	06/02	9:00 AM	12:40 PM	fishione	rttomo	M barkeri	2	On arrival: col. vac: 32, 
extraction voltage: 3800. (hv off)

Beam current today: 22.6 nA, twice
as high as one week ago	The default rod seems to be a bit 
leaky...
Fri Jun  4 17:59:56 2010	Janet Meehl Fox/Winey	6/4/10	10:45 AM	6:00 PM	tiltrotate	rttomo	Tet poc1delete bbs, WTtet bbs	6		Before taking the gain reference, 
the upper right quadrant of the 
camera image was lighter, both in 
the trial and record mode.  After 
the gain ref was taken, everything 
was fine.
Mon Jun  7 12:18:48 2010	morphew	6-7	8 AM	12 AM	tiltrotate	rttomo	aditi neurons	1		
Mon Jun  7 15:46:30 2010	Joanna	06/07	12:30 PM	3:40 PM	standard	training	training with Kristin	0		
Tue Jun  8 17:38:49 2010	eileen	6/8	8:45 AM	5:30 PM	tiltrotate	rttomo	chlamy flagella	6		
Wed Jun  9 12:05:38 2010	eileen	6/9	9 AM	12 PM	tiltrotate	rttomo	chlamy flagella	2		
Wed Jun  9 17:33:15 2010	Janet Meehl Fox/Winey	6/9/10	12:30 PM	5:45 PM	tiltrotate	rttomo	WTtet new basal bodies	6		
Thu Jun 10 14:47:35 2010	parkkm	6-10	7:30 AM	2:45 PM	tiltrotate	rttomo	hela midbody	4		upper right camera quadrant very 
light in comparison
Fri Jun 11 12:00:09 2010	cedric	06/11	8:15 AM	12:00 PM	tiltrotate	rttomo	pml-minus-virus	1		
Fri Jun 11 16:47:19 2010	eileen	6/11	12:15 PM	4:45 PM	tiltrotate	rttomo	chlamy flagella	4		
Mon Jun 14 16:54:20 2010	Kristin	6/14	9:30 AM	5:30 PM	tiltrotate	rttomo	HeLa	4		
Tue Jun 15 16:21:26 2010	eileen	6/15	8:15 AM	4:15 PM	70	rttomo	Odde Ase1 delete	8		
Thu Jun 17 17:39:45 2010	cedric	06/17	10:00 AM	5:30 PM	tiltrotate	rttomo	polyoma virus pml-- cells	2		
Fri Jun 18 16:58:41 2010	cedric	06/18	10:00 AM	17:00 PM	tiltrotate	rttomo	polyoma-pml-- cells	2		jump or shift of the beam around 38 
degrees. start the tilt series at 
60 and jumps after 21 or 22 images 
2 times in a row. I do not 
understand why the tracking in 
SerialEM  didn't account for it.
Mon Jun 21 17:12:08 2010	cedric	06/21	10:00 AM	17:30 PM	tiltrotate	rttomo	polyoma-pml--cells	2		
Tue Jun 22 17:30:33 2010	cedric	06/22	10:00 AM	17:30 PM	tiltrotate	rttomo	polyoma-pml--cells	2		
Wed Jun 23 13:12:00 2010	Chuck Robertson	06/22	8:10 AM	1:10 PM	fishione	rttomo	M barkeri	2	beam: 6.6nA	
Fri Jun 25 12:31:57 2010	morphew	6-25	8:30 AM	12:15 PM	tiltrotate	rttomo	aditi neurons	4		
Fri Jun 25 17:59:33 2010	Janet Meehl Fox/Winey	6/25/10	12:40 PM	6:00 PM	tiltrotate	rttomo	WTtet new basal bodies	4	Upper right quadrant looked 
different on taking the first 
record image before doing the gain 
ref.  Once the gain ref was taken, 
everything looked fine.

Ultracam work done during session 
by David M. and Cindi S.	
Tue Jun 29 11:49:29 2010	eileen	6/29	8:30 AM	11:45 AM	tiltrotate	rttomo	Odde yeast		E data drive making funny sound- 
Cindi turned off. Upper right 
quadrant funky.	
Wed Jun 30 12:19:52 2010	Chuck Robertson	06/30	8:00 AM	12:15 PM	fishione	rttomo	Mbarkeri	2	beam: 8.54 nA	right trackball has problems
Wed Jun 30 17:56:14 2010	cedric	06/30	13:30 PM	18:00 PM	tiltrotate	rtimage	polyoma-a31	0		
Thu Jul  1 10:24:00 2010	Janet Meehl Fox/Winey	7/1/10	9:15 AM	10:15 AM	standard	rttomo	WTtet	0		Couldn't get rid of the quadrant 
lines after taking a gain ref. The 
left 1/2 looks ok but the right 
upper and lower quadrants are 
noticeable as well as a few 
horizontal lines in the upper right 
quadrant.  I didn't start a tilt 
series.
Tue Jul  6 13:52:21 2010	Johanna/Neil	07/06	9:00 AM	1:45 PM	fishione	rttomo	trypanosome collar	1		Microscope shut down HT three times 
in one hour. Service called.
Thu Jul  8 15:10:36 2010	spilman	7/7	9:00 AM	4:30 PM	standard	serviceman	HT issues & PM	0	HT 50 volt supply replaced, HD 
backup, water leak electronics rack 
(replaced hose & clamps	PM & HT issues
Thu Jul  8 15:12:20 2010	spilman	7/8	8:00 AM	4:00 PM	standard	serviceman	PM	0	PM & HT issues	PM & HT issues
Fri Jul  9 18:07:39 2010	Janet Meehl Fox/Winey	7/9/10	12:15 PM	6:20 PM	tiltrotate	rttomo	WTtet assembling basal bodies	6		Camera was still showing 
quadrants.  David worked some magic 
and got a homogenous gain ref.
Sun Jul 11 16:48:39 2010	DNM	7/11	2:00 PM	4:45 PM	standard	serviceman	quadrant correction, setting up 
data disks	0	I don't like the bumps in the 
quadrant correction curves.  Right 
now binnedx2 works fine on a 
binnedx2 gain reference but not so 
well on an unbinned reference	
Mon Jul 12 17:59:34 2010	DNM	7/12	11:00 AM	11:45 AM	standard	serviceman	Camera problems	0	Upper right quadrant gain sometimes 
varies a few percent	
Mon Jul 12 18:00:26 2010	DNM	7/12	12:01 PM	6:00 PM	standard	development	Contemplation and fixing settling 
in walkup	0		
Tue Jul 13 12:28:47 2010	eileen	7/13	8:50 AM	12:30 PM	tiltrotate	rtimage	Odde yeast			
Tue Jul 13 16:55:34 2010	cedric	07/13	13:00 PM	17:00 PM	tiltrotate	rttomo	pml--	1		upper cam sucks
Wed Jul 14 18:35:56 2010	cedric	07/14	13:00 PM	18:45 PM	tiltrotate	rttomo	A31-accumulation-virus	1		camera sucked again
Thu Jul 15 11:56:39 2010	morphew	7-14	8 AM	11:50 AM	tiltrotate	rttomo	PyV	3		
Thu Jul 15 14:05:36 2010	DNM	7/14	11:55 AM	2:05 PM	tiltrotate	development	testing tilt range assessor	0		
Fri Jul 16 16:22:56 2010	DNM	7/16	9:30 AM	10:30 AM	standard	development	more tilt range testing			
Fri Jul 16 16:23:29 2010	DNM	7/16	10:30 AM	4:25 PM	standard	serviceman	camera video board work			
Mon Jul 19 10:57:46 2010	morphew	7-19	9 AM	11 AM	tiltrotate	rttomo	tour to richard easingwood	1		
Mon Jul 19 16:24:52 2010	Chuck Robertson	07/19	1:00 PM	4:30 PM	fishione	rttomo	Mbarkeri	2	Beam = 7.85 nA	
Tue Jul 20 13:47:42 2010	Cindi/Joey	07/16	9:00 AM	1:45 PM	f30cryo	cryoimaging	Giardia cytos		Too humid to work and grids not so 
good.	
Tue Jul 20 18:37:03 2010	cedric	07/20	2:30 PM	19:00 PM	tiltrotate	rttomo	a31-polyoma	2		
Wed Jul 21 18:32:29 2010	cedric	07/21	7:00 AM	6:40 PM	tiltrotate	rttomo	a31-polyoma-accumulation	3		
Thu Jul 22 15:16:38 2010	morphew	7-22	8 AM	3:30 PM	tiltrotate	rttomo	melanoma	4		
Fri Jul 23 12:51:19 2010	DNM	7/23	9:40 AM	12:55 PM	standard	serviceman	Replacing card in upper camera 
controller, rerunning quadrants			
Fri Jul 23 18:00:41 2010	cedric	07/23	1:30 PM	6:00 PM	tiltrotate	rttomo	a31-virus-accumulation	2		
Mon Jul 26 18:02:39 2010	Cindi/Joey	07/26	9:30 AM	6:00 PM	f20cryo	cryotomo	Giardia Cytos	6	Great day!	Needed to reload alignments. Beam 
was weird.
Tue Jul 27 12:21:06 2010	Cindi	07/27	8:00 AM	12:20 PM	f30cryo	cryoimaging	Microtubules		samples sucked	vacuum crash removing rod.
Wed Jul 28 16:46:56 2010	eileen	7/28	11 AM	4:50 PM	tiltrotate	rttomo	DeCamilli endophilin TKO	4		
Fri Jul 30 15:13:25 2010	eileen	7/30	8:50 AM	3:15 PM	tiltrotate	rttomo	DeCamilli endophilin control	4		
Mon Aug  2 16:44:57 2010	Chuck Robertson	08.02	1:00 PM	4:45 PM	fishione	rttomo	M barkeri	2	beam = 9.52 nA	
Tue Aug  3 14:35:59 2010	eileen	8/3	8:40 AM	2:30 PM	tiltrotate	rttomo	DeCamilli endophilin TKO	6		
Fri Aug  6 16:30:15 2010	cindi/guobin	08/06	8:30 AM	4:30 PM	f30cryo	cryotomo	polio infected cells	1		
Sat Aug  7 16:45:51 2010	Cindi/Guobin	08/07	9:00 AM	5:00 PM	f30cryo	cryotomo	Poliovirus infected Hela cells			
Sun Aug  8 20:22:35 2010	Guobin	8/8	9:00 AM	8:00 PM	f30cryo	cryotomo	Poliovirus infected Hela cells			The column vacuum read 18 with RT 
holder in the column this morning 
when I came in. The vacuum read 
decreased to 6 in 15 minutes after 
RT holder was removed.
Mon Aug  9 15:56:34 2010	Chuck Robertson	08/09	1:00 PM	3:52 PM	fishione	rtimage	Mbarkeri, desiccated	0	Beam = 8.92 nA	Vacuum on the rods was very 
unpredictable today.  Both the 
default rod and fischione were 
very, very slow to drop below 
13...  ended my session early 
because the vacuum would not go 
below 20.  The pattern was, start 
the twist, about 2/3 around the 
vacuum jumps to 35 or greater.  
Insert rod, vacuum either stays at 
35 ("forever", eg > 15 min) or 
drops to ~20 and goes no lower 
(last attempt, let it set for 20 
minutes, no change.)  
Fri Aug 13 13:32:35 2010	Scott Wadewitz	08/10	9:30 AM	4:00 PM	standard	serviceman	Checking out camera/GIF			
Fri Aug 13 13:32:42 2010	Scott Wadewitz	08/11	9:30 AM	4:00 PM	standard	serviceman	Checking out camera/GIF			
Fri Aug 13 13:32:46 2010	Scott Wadewitz	08/12	9:30 AM	4:00 PM	standard	serviceman	Checking out camera/GIF			
Fri Aug 13 13:33:37 2010	Scott Wadewitz	08/13	9:30 AM	4:00 PM	standard	serviceman	Checking out camera/GIF			Camera getting sent back for o-ring 
service. Small vacuum leak found 
1min after air compressor turns 
off. Can't find evidence of other 
vacuum leak within Gatan system.
Sun Aug 15 16:53:19 2010	DNM	8/15	2:00 PM	5:00 PM	70	development	Testing multiple TS	1		
Mon Aug 16 17:58:57 2010	DNM	8/16	11:00 AM	6:00 PM	70	development	Testing multiple TS	3		
Tue Aug 17 12:31:04 2010	spilman	8/16	8:30 AM	9:00 AM	standard	serviceman	test		Z axis inoperative from panel, 
ordered new	
Wed Aug 18 17:00:14 2010	Kirmse	8/18	9:30 AM	5:00 PM	f30cryo	cryoimaging	IFs	0	during LM search on the grid column 
vacuum went up to 26 once closed 
valves waited for it to come down 
afterwards everything was ok,

had some trouble with beam 
alignments, reloaded them from file 
then	
Thu Aug 19 18:41:28 2010	DNM	8/19	10 AM	6:50 PM	70	development	multiple TS, cooking and beam 
center problems with LD			
Fri Aug 20 17:33:45 2010	DNM	8/20	3.00 PM	6:30 PM	70	development	testing multiple TS			Mouse left button is unreliable
Mon Aug 23 15:46:13 2010	DNM	8/23	10:30 AM	4:50 PM	70	development	checking balance shifts in LD and 
navigator			
Tue Aug 24 16:10:01 2010	cindi	08/24	9:00 AM	4:00 PM	f30cryo	cryoimaging	giardia mbp-gfp immuno			column vacuum was at 25 when I 
arrived--very rainy day. coldfinger 
brought vacuum down quickly. stage 
had very jerky movements, 
especially when moving large 
distances in X and near the edge.
Wed Aug 25 17:39:48 2010	Johanna	08/25	9:00 AM	5:40 PM	fishione	rttomo	trypanosome collar	0		After alpha-wobbler, the stage did 
not go back to zero but to 15. 
Thu Aug 26 13:08:12 2010	morphew	8-26	8 AM	1 PM	tiltrotate	rtimage	PyV			
Thu Aug 26 16:30:46 2010	DNM	8/26	10:45 AM	12:00 AM	tiltrotate	development	recalibrate mag offsets, check out 
view area shifting			
Thu Aug 26 16:31:49 2010	DNM	8/26	3:00 PM	4:30 PM	70	development	Check out rotated low dose axis			
Wed Sep  1 15:57:42 2010	morphew	9-1	8 AM	4 PM	70	rttomo	PyV	1		
Thu Sep  2 14:23:17 2010	morphew	9-2	8 AM	2 PM	70	rttomo	PyV	1		col vac at 34 at 8am.; Added LN2, 
pumped to 6 after 2 hours
Wed Sep  8 16:47:50 2010	DNM	9/8	2:30 PM	4:45 PM	standard	development	checking out pixel/intensity 
matching	0		Gain of cameras seems to be wrong 
in properties file.  UScan is about 
18 not 12, and UCam is 8.1 to 
10.4.  The latter could be due to 
the last card change where one 
quadrant had a gain of 80% relative 
to the others... but the UScan 
change is inexplicable.
Thu Sep  9 17:10:09 2010	Johanna	9/9	09:20 AM	5:20 PM	fishione	rttomo	trypanosome collar	2		
Mon Sep 13 15:28:35 2010	Chuck R	9/13	1:00 PM	3:30 PM	fishione	rtimage	E COLI X1411		BEAM = 9.50nA

Vacuum was very nice... very solid.	
Tue Sep 14 14:53:28 2010	Johanna	9/14	9:30 AM	3:00 PM	fishione	rttomo	trypansome collar	2		Had a bad quadrant in the same 
place as it was before. Only 
happened once though.
Tue Sep 14 18:17:47 2010	DNM	9/14	3:45 PM	6:15 PM	70	development	testing HQ stage montaging	0		
Thu Sep 16 09:56:45 2010	morphew	9-16	8 AM	10 AM	tiltrotate	rttomo	PyV	1-2x2		
Thu Sep 16 09:57:28 2010	morphew	9-15	8 AM	4 PM	tiltrotate	rttomo	PyV	3-2x2		
Thu Sep 16 15:48:12 2010	Robert Kirmse	9/16	10:30 AM	4:00 PM	f30cryo	cryoimaging	IFs	0		Stage was always giving "Tracking 
Error" and moving vigorously: but 
doing the LMM nevertheless.

Camera Quadrant was messed up 
(controller restart helped a bit)

Crashed Vaccum once because the 
program wouldn't accept holder 
choice but lock-in timer started 
and red light on the stage went 
out. Inserting the rod resultet in 
vac crash.
Fri Sep 17 13:50:50 2010	DNM/Cindi	09/17	10:00 AM	12:00 AM	standard	serviceman	camera quadrant issues			Swapped boards/cables and figured 
out the new board is bad. Gatan 
shiipping one to arrive on Monday 
to replace. The bad board is the 
bottom board.
Fri Sep 24 17:35:58 2010	DNM	9/24	10:15 AM	5:30 PM	tiltrotate	development	New GIF distortion images, mag 
gradients
	0	Tuned GIF 9400-34000, saved GIB 
settings	
Tue Sep 28 10:25:51 2010	DNM	9/28	9:45 AM	10:30 AM	70	development	Mary's mag gradient problem	0	Quadrant syill bad, saw small and 
large glitches	
Tue Sep 28 17:08:51 2010	cedric	09/28	11:00 AM	5:30 PM	tiltrotate	rtimage	plastic virus stuff	0		
Wed Sep 29 15:57:48 2010	morphew	9-29	8 AM	4 PM	fishione	rttomo	giardia ltfs	6		
Thu Sep 30 14:01:57 2010	morphew	9-30	8 AM	2 PM	fishione	rttomo	PyV	6		
Fri Oct  1 18:34:01 2010	Cindi/David	10/1	9:30 AM	6:30 PM	fishione	training	Grand user trial and training on 
multi-TS	3		
Tue Oct  5 18:01:29 2010	eileen	10/5	8:45 AM	6:00 PM	fishione	rttomo	auto acquire test series	4	lots of stage track errors. 
Navigator software for auto acquire 
is great!	
Wed Oct  6 15:41:32 2010	morphew	10-6	8 AM	4 PM	fishione	rttomo	PyV	6		
Thu Oct  7 15:06:46 2010	Cindi	10/07	11:00 AM	3:00 PM	standard	serviceman	Taking off upper camera			Tried to get images of bad quadrant 
and couldn't get it to go super 
bad. Tried different binnings and 
doses. Anyway, stole some images 
from other sources. Removed camera 
and got it ready for shipping. 
Requested shipping box from Gatan. 
Returning controller and both extra 
video boards.
Thu Oct 14 17:19:44 2010	Spilman	10/11	8:00 AM	6:45 PM	standard	serviceman	Tip change		New Tip	
Thu Oct 14 17:20:59 2010	Spilman	10/12	8:00 AM	7:00 PM	standard	serviceman	Tip change		tip change	
Thu Oct 14 17:22:05 2010	Spilman	10/13	7:30 AM	6:15 PM	standard	serviceman	Tip change		Tip change	
Thu Oct 14 17:23:10 2010	Spilman	10/14	8:00 AM	5:45 PM	standard	serviceman	Tip change		Tip change	ip
Fri Oct 15 12:31:25 2010	cedric	10/15	10:00 AM	12:30 PM	tiltrotate	rtimage	virus sections			scope screwed up. alignments done 
with gun lens bug. shutter left in 
open state. corner of aperture in 
every GIF image.
Fri Oct 15 16:34:12 2010	cindi	10/15	3:30 PM	4:30 PM	standard	serviceman	retune GIF			retuned GIF and aperture went away. 
It was just way off centered. Did 
not cryo-cycle since P15 isn't 
reading right. Just let the cold-
finger warm up gradually.
Fri Oct 22 16:56:51 2010	cedric	10/22	9:30 AM	17:00 PM	f30cryo	cryoimaging	section	0		
Fri Oct 22 17:01:15 2010	DNM	10/22	3:45 PM	5:00 PM	f30cryo	development	fixing montage problem			
Mon Oct 25 11:59:47 2010	Ron Spilman	10/18	8:00 AM	5:00 PM	standard	serviceman	work on compustage			work on compustage
Mon Oct 25 11:59:52 2010	Ron Spilman	10/19	8:00 AM	5:00 PM	standard	serviceman	work on compustage			work on compustage
Mon Oct 25 11:59:56 2010	Ron Spilman	10/20	8:00 AM	5:00 PM	standard	serviceman	work on compustage			work on compustage
Mon Oct 25 12:00:00 2010	Ron Spilman	10/21	8:00 AM	5:00 PM	standard	serviceman	work on compustage			work on compustage
Mon Oct 25 16:56:39 2010	DNM	10/25	12:30 PM	5:00 PM	tiltrotate	development	interpolated dark refs and aligning 
with rotation			
Tue Oct 26 12:12:59 2010	Cindi	10/26	10:30 AM	12:15 PM	standard	serviceman	Install Controller			Controller came bad. It cannot read 
temperature and goes to -199 
quickly, then reads -1. Happens 
when no cable is attached, so cable 
is good. Sending controller back.
Tue Oct 26 16:44:48 2010	roy	10/26	1 PM	5 PM	standard	training	EELS			
Wed Oct 27 16:24:31 2010	roy	10/27	10:30 AM	4:30 PM	standard	development	EFTEM FeO particles			
Wed Oct 27 19:39:07 2010	DNM	10/27	4:30 PM	7:40 PM	standard	serviceman	complete waste of time		Computer gave a blue-screen crash 
for previous user.  DM prefs files 
were unreadable and it was very 
funky, with bad pixel alignments.
Got prefs files from backup disk, 
it is OK now.  But then the HT went 
out.	nothing but problems
Thu Oct 28 09:55:31 2010	cindi	10/28	9:30 AM	9:50 AM	standard	serviceman	HT out			HT won't go stabily beyond 4kV, 
sometimes will spike up to 10kV. 
Ron suggests new tank or cable 
needed. Will come in this afternoon 
to see.
Fri Nov 19 15:39:41 2010	cindi	11/19	9:30 AM	3:30 PM	standard	serviceman	checking fat dewar and column vacuum			column vacuum at 6 using Ron's 
holder at 9:30. Changed to our 
holder and vacuum spiked to 26. 
Vacuum took a long time to settle 
around 10, but fluctuated a lot 
(after adding LN2). Re-inserted 
Ron's holder and it spiked to 46 
and also took a long time to settle 
with LN2 on. Fat dewar looks like 
it might last 8 hours in current 
configuration. Need to make more 
insulation.
Tue Nov 30 17:37:23 2010	DNM	11/30	4:00 PM	5:40 PM	standard	development	a few bugs			
Fri Dec  3 10:36:14 2010	moe	12/3	8:30 AM	10:45 AM	standard	serviceman	Align GIF	0		
Fri Dec  3 15:04:01 2010	DNM	11/3	1:30 PM	3:00 PM	standard	development	Testing automatic termination on 
dim images	4	Right, 4 tilt series of 6 images 
each!	
Sun Dec  5 10:02:35 2010	DNM	11/5	9:20 AM	10:05 AM	standard	serviceman	Checking gain reference and redoing		"Burned" spot now down to 4% dimmer	
Mon Dec  6 17:01:19 2010	eileen	12/6	8:30 AM	5:00 PM	fishione	rttomo	spd5(RNAi) 	6	Scope worked great! Some 
adjustments to exposure time and 
target counts needed.	
Tue Dec  7 15:00:45 2010	Chuck Robertson	12/07	8:15 AM	3:00 PM	fishione	rttomo	E. coli MC4100	2	Beam = 10.95 nA	None!
Wed Dec  8 12:00:08 2010	Cindi	12/8	8:00 AM	12:00 PM	fishione	rttomo	Giardia CPD cytos	2	3X3 montages	
Wed Dec  8 16:17:59 2010	Geiss	Dec 8	1:00 PM	4:00 PM	standard	rtimage	Carbon mtl	0		
Thu Dec  9 12:12:51 2010	Janet Meehl Fox/Winey	12/9/10	8:15 AM	12:15 PM	tiltrotate	rttomo	Tetrahymena poc1delete	4		
Thu Dec  9 15:55:21 2010	Cindi	12/9	1:00 PM	4:00 PM	standard	rtimage	Giardia CPD cytoskeletons			Had to reset goniometer because I 
went to far in the stage. mage 
didn't shift between panels of 
montage one time only. Bug in 
SerialEM that David saw about image 
shift limits in montaging. Vacuum 
spiked pretty high inserting rod 
several times.
Fri Dec 10 16:57:23 2010	DNM	12/10	4:15 PM	5:00 PM	70	serviceman	Fixed shutter connections!			
Mon Dec 13 15:30:43 2010	Cindi	12/13	11:30 AM	3:30 PM	fishione	rttomo	giardia cytos	6		vacuum was spiking VERY high when 
inserting rods (T/R with Mary, 
standard, and Fischione)--up to 64. 
Sometimes, tilting stage to 60, 
waiting a bit, then tilting to 0 
brought the vacuum down rapidly. 
Both IGPs spiked.
Tue Dec 14 14:49:31 2010	morphew	12-14-10	12:30 AM	3 PM	tiltrotate	rttomo	PyV	2		
Wed Dec 15 12:06:15 2010	Janet Meehl Fox/Winey	12/15	8:15 AM	12:00 PM	tiltrotate	rttomo	Tet poc1delete bbs	5	No problems with the column vac.	
Wed Dec 15 16:15:04 2010	Roy Geiss	12/15	12:30 PM	4:30 PM	standard	rtimage	FeO - ZnO		Image in focus on viewing screen 
was not in focus on ULTRACAM	Image in focus on viewing screen 
was not in focus on ULTRACAM
Thu Dec 16 20:05:34 2010	Johanna	12/16	11:00 AM	8:00 PM	fishione	rttomo	trypanosome collar	8		
Mon Dec 20 14:26:44 2010	Cindi	12/20	9:30 AM	2:30 PM	fishione	rttomo	giardia	4		column vacuum at 13 when I got in. 
Checked Fischione oring before 
using and it was covered with hair. 
Checked standard oring and it had a 
few hairs. Stage had difficult time 
moving in X. 
Thu Dec 23 13:49:50 2010	Cindi	12/23	8:00 AM	2:00 PM	tiltrotate	rttomo	Giardia	4		vacuum at 6 when I came in. Vacuum 
spiked high inserting both 
fischione and T/R holders. Had to 
tilt to -60 to get vacuum to drop. 
Several times, vacuum spiked into 
20s with large stage movements.
Tue Dec 28 10:12:28 2010	Chuckr	12/28	8 AM	10 AM	standard	rtimage	none	0	not usable due to vacuum problems	Vacuum on arrival: 34
After 30 min cooling: 34
At -60 degrees: 34
Remove rod drifts to 24
Standard rod has many fiber under 
disecction scope... some as long as 
8 mm

Replace std rod... vac spikes to 38
Goes to 24 on full insert

Vacuum is erratic... jump 5-10 
values between measurements, eg 12 
to 22 in successive col vac values.

Not using scope today because of 
this.
Wed Dec 29 17:12:51 2010	Janet Meehl Fox/Winey	12/29	9:30 AM	5:30 PM	tiltrotate	rttomo	Tet poc1delete bbs	9	Vacuum was good. I had to tilt 
every time to get vacuum to drop 
after rod insertion but it always 
came down. I was careful when 
inserting the rod and didn't have 
any spikes over 35.  No spikes when 
stage moved.  Went over o-rings 
with a fine-toothed tweezer.	
Tue Jan  4 15:19:06 2011	Chuck R	01/04	8:00 AM	3:15 PM	fishione	rttomo	M barkeri	4	Beam = 8.21 nA	Rods still fuzzy when they come out 
of the scope, but not as bad as 
last week.

Col vac spiked to ~28 when 
inserting the rods... had to go 
very, very slowly... when the rods 
were first seated the vacuum was 
often 16-20.  The magic recipe du 
jour was to rotate to +60... this 
always brought the col vac back to 
6.

Other than twitchy rod insert, the 
machine was very stable today.
Wed Jan  5 14:35:50 2011	Chuck Moe	1/5	08:30 AM	14:30 PM	standard	serviceman	stage service		cleaned stage orings and realigned 
stage. Performed EFTEM alignments.	
Thu Jan  6 15:47:30 2011	Cindi	01/06	1:00 PM	3:30 PM	standard	rtimage	Giardia cpd cytos			
Fri Jan  7 16:19:26 2011	eileen	1/7	8:15 AM	4:15 PM	fishione	rttomo	c. elegans spd5	9	scope vac and stage GREAT! 
Completely fixed.	
Mon Jan 10 16:58:07 2011	eileen	1/10	9 AM	5:00 PM	fishione	rttomo	C. elegans spd5	10	O=rings are still  a little fuzzy. 
be sure to look at them under the 
disection scope.	
Tue Jan 11 14:46:36 2011	morphew	1-11	9 AM	3 PM	f20cryo	cryoimaging	PyV A31			
Wed Jan 12 16:32:32 2011	eileen	1/12	8:45 AM	4:30 PM	fishione	rttomo	De Camilli; WT recovery and Dynamin	6		
Thu Jan 13 15:33:15 2011	Cindi	01/13	8:00 AM	1:00 PM	standard	rtimage	Giardia CPD cytos			Stage felt a little squirely. 
vacuum spiked once to 46 when 
inserting rod. 
Fri Jan 14 16:28:18 2011	roy	01/14	1:00 PM	4:30 PM	standard	development	HRTEM Apatite			
Tue Jan 18 14:37:18 2011	morphew	1/18	12:30 AM	2:30 PM	tiltrotate	rttomo	ltfs/giardia	2		
Wed Jan 19 15:04:24 2011	eileen	1/19	8:30 AM	3:00 PM	tiltrotate	rttomo	Tomoko H40 spindles	3 2x3		
Thu Jan 20 15:49:32 2011	Cindi	01/20	8:00 AM	3:45 PM	tiltrotate	rtimage	Giardia Sxns			Vacuum was at 13 when I logged in. 
It wasn't until noon that it got to 
6 with holder in. 
Fri Jan 21 11:48:39 2011	eileen	1/21	9:30 AM	11:45 AM	tiltrotate	rttomo	Tomoko H40	1 2x3		
Fri Jan 21 14:00:09 2011	Cindi	01/20	12:00 PM	2:00 PM	tiltrotate	rtimage	Giardia control samples			had it happen twice with a bad 
lower left quadrant (much, much 
brighter than the others). 
Tue Jan 25 11:11:48 2011	Chuckr	01/25	7:45 AM	11:15 AM	fishione	rtimage	Mbarkeri	0		Mag knob sticks/jumps
Tue Jan 25 17:32:49 2011	Johanna/Neil	01/25	1:00 PM	5:30 PM	fishione	rttomo	trypanosome collar	4		
Wed Jan 26 09:16:30 2011	cindi	01/26	8:00 AM	9:00 AM	standard	serviceman	vacuum recovery			Last night vacuum crashed while 
inserting rod. Could not recover. 
Told Johanna to go home. This 
morning, vacuum still didn't 
recover. Tried cycling vacuum. No 
help. Removed holder. Vacuum 
instantly recovered to 49 and 
dropped to 21 within 30sec. Found a 
couple of hairs on o-ring. Decided 
to replace o-ring. Inserted rod 
(burped to 47) and cooled cryo-box. 
Vacuum down to 6.
Fri Jan 28 07:39:38 2011	Ron Spilman	01/26	1:30 PM	3:30 PM	standard	serviceman	try to fix compustage leak			Ordered new o-rings
Fri Jan 28 07:40:42 2011	Ron Spilman	01/27	8:00 AM	5:30 PM	standard	serviceman	try to fix compustage leak			replaced various o-rings, stage still leaking at 
high tilt and extreme X. Major burp of IGP
Fri Jan 28 17:08:44 2011	Spilman	1/28	8:00 AM	5:30 PM	standard	serviceman	Service	0	Compustage vacuum leak.  Ordered 
hemispherical bearing, bearing cup. 
oring set, & IGP	Scope has not pump below 14 without 
houlder or 23 with holder.  Was 
able to tilt +-79 degrees at 
extreme Y cordinace without 
leaking. 
Mon Jan 31 10:57:23 2011	JohannaCollar	01/27	09:30 AM	11:00 AM	fishione	rttomo	Collar	0		Vaccum broke as I did eucentricity. 
Fri Feb  4 19:53:15 2011	spilman	1/31	9:00 AM	5:00 PM	standard	serviceman	service	0	stage vacuum leak	
Fri Feb  4 19:53:37 2011	spilman	1/31	9:00 AM	5:00 PM	standard	serviceman	service	0	stage vacuum leak	
Fri Feb  4 19:54:20 2011	spilman	1/31	9:00 AM	5:00 PM	standard	serviceman	service	0	stage vacuum leak	
Fri Feb  4 19:55:20 2011	spilman	2/1	9:00 AM	5:00 PM	standard	serviceman	service	0	stage vacuum leak	
Fri Feb  4 19:56:04 2011	spilman	2/2	9:00 AM	5:00 PM	standard	serviceman	service	0	replace ion getter pump	
Fri Feb  4 19:57:26 2011	spilman	2/3	9:00 AM	6:30 PM	standard	serviceman	service	0	stage alignment issue clamping block	
Fri Feb  4 19:58:32 2011	spilman	2/3	6:30 AM	9:00 PM	standard	serviceman	service	0	stage alignment issue clamping block
 vacuum issues returned	
Mon Feb  7 14:17:37 2011	Cindi	02/07	1:45 PM	2:15 PM	standard	rtimage	Giardia CPD			Stage moved OK, but did no tilting
Wed Feb  9 12:23:05 2011	spilman	2/7	7:00 AM	1:00 PM	standard	serviceman	stage issues	0	shift in X when tilting	ordered parts
Wed Feb  9 12:24:48 2011	spilman	2/9	7:00 AM	1:00 PM	standard	serviceman	stage issues	0	Replaced X donator, but still have 
X shift during tilting	ordered parts
Thu Feb 10 18:11:19 2011	Johanna	02/10	10:30 AM	18:10 PM	fishione	rttomo	trypanosome collar	6		-first the scope did not tilt more 
than to 22 degrees. Ron fixed that
-the lower left quadrant of the 
camera showed brighter than the rest
-the vacuum broke as i did the walk 
up. Holder is stuck in compustage
Tue Feb 15 16:05:37 2011	spilman	2/10	3:00 PM	5:00 PM	standard	serviceman	stuck rod & vac leak		Removed rod and checked for vacuum 
leak.  Scheduled to install 
floating rod on the 11th.  	
Tue Feb 15 16:07:59 2011	spilman	2/11	8:00 AM	6:00 PM	standard	serviceman	Floating Rod Replacement		Disassembled and re-assembled 
compustage with new floating rod.  
Allowed to pump down over weekend.	
Tue Feb 15 16:09:09 2011	spilman	2/14	8:00 AM	5:30 PM	standard	serviceman	Floating Rod Replacement		Started stage alignment and vacuum 
checking.  Vacuum in column was 14 
over the weekend.  Trouble-shot gun 
shift issues.	
Tue Feb 15 16:10:04 2011	spilman	2/15	8:00 AM	5:00 PM	standard	serviceman	Floating Rod Replacement		Aligned compustage and checked for 
vacuum leaks and tilt shift 
issues.  Job complete.	
Wed Feb 16 10:18:07 2011	Cindi	02/16	8:00 AM	10:15 AM	tiltrotate	serviceman	Testing compustage			Vacuum at 6. Put in T/R holder, 
would not move in X. Tried gonio 
reset and had unresolvable pole 
touch. Called Ron and reset 
compustage screw on clamp. Reset 
gonio worked. Fischione holder 
works fine with test TS and test 
LMM. Put in F30 cryo holder. Does 
not move in X. Reset gonio. Ron to 
come in morning to fix v-groove.
Wed Feb 16 17:08:19 2011	Johanna	02/16	10:30 AM	05:00 PM	fishione	rttomo	trypanosome collar	8		
Thu Feb 17 13:19:03 2011	morphew	2-17	8 AM	11 PM	f30cryo	cryoimaging	liposomes nanogold			stage would not translate--service 
called
Thu Feb 17 13:58:37 2011	Ron Spilman	02/17	11 AM	12 PM	f30cryo	serviceman	worked on compustage			compustage was sticking with 
holder. Ron came in and removed the 
silver ring on the purple plate. 
Did extensive testing. Should be OK 
now.
Fri Feb 18 15:25:37 2011	spilman	2/17	12:00 PM	3:00 PM	f30cryo	serviceman	test		stage failed on cryo holder removed 
ring from front plate	
Fri Feb 18 15:27:51 2011	spilman	2/17	10:00 AM	4:00 PM	tiltrotate	serviceman	test		stage failed on tily/rotate  
adjusted set screws for horse shoe 
and set screw under X motor	
Mon Feb 21 15:32:18 2011	morphew	2-21	8 AM	3:30 PM	f30cryo	cryotomo	liops	2		
Tue Feb 22 09:54:29 2011	cindi	02/22	8:00 AM	10:00 AM	tiltrotate	rtimage	Giardia CPD			Yesterday, vacuum at 27 when 
arrived. Today, vacuum at 20 when 
arrived. Both days: Stage not 
translating past -500 in Y. Grid 
doesn't seem 'centered'. Made call 
to service.
Wed Feb 23 08:29:11 2011	Chuck Moe	2/23	06:30 AM	09:00 AM	standard	serviceman	service		Adjusted Y stage movement. Found 
bad P1 lens current. Lens checked 
and good. Reseated all lens 
electronics	P1 now working and rotation problem 
fixed. Stage movement good
Wed Feb 23 12:43:46 2011	DNM	11/23	10:15 AM	12:45 PM	standard	serviceman	checking stage cal, looking at 
quadrant problem			
Wed Feb 23 17:16:24 2011	Johanna	2/23	1:00 PM	5:15 PM	fishione	rttomo	tryps flagella	4		Moving sample using the red cross 
does not work at low mag. Davod has 
called service
Thu Feb 24 17:20:16 2011	eileen	2/24	8:30 AM	5:20 PM	fishione	rttomo	C. elegans montage embryos Thomas MR	2 2x2		
Fri Feb 25 17:46:33 2011	Johanna	2/25	10:40 AM	6:00 PM	fishione	rttomo	trypanosome flagellum	3		
Fri Feb 25 20:13:32 2011	Chuck Moe	2/22	12:00 PM	6:00 PM	standard	serviceman	compustage			worked on compustage and Projector lens 1 
issues.
Tue Mar  1 17:43:09 2011	eileen	3/1	8:30 AM	5:30 PM	tiltrotate	rttomo	CDC20 serials	8	Scope working great. Lots of fibers 
on o-rings so be sure to check!	
Wed Mar  2 16:34:42 2011	Johanna	3/2	10:30 AM	5:00 PM	fishione	rttomo	Trypanosome flagella	4		
Thu Mar  3 13:21:13 2011	morphew	3-2	8 AM	1 PM	f30cryo	cryotomo	liposomes	1		extra low contrast, stage not 
translating properly
Thu Mar  3 18:27:18 2011	DNM	3/3	2:00 PM	6:30 PM	standard	development	Chasing montage bugs			
Fri Mar  4 13:06:42 2011	eileen	3/4	8:30 AM	1:00 PM	fishione	rttomo	wt_K-Stim DeCamilli neurons	6		
Fri Mar  4 15:41:06 2011	Cindi	03/04	1:10 PM	3:40 PM	fishione	rttomo	Giardia CPD of MBP-GFP	1	Finally!	
Mon Mar  7 19:11:45 2011	Cindi	03/07	11:00 AM	7:15 PM	fishione	rttomo	Giardia CPD of MBP-GFP	5		Stage seemed to lose some ability 
on the right-click and drag.
Tue Mar  8 15:38:06 2011	morphew	3/8	11 AM	3:30 PM	tiltrotate	rttomo	u2os ltfs	4		
Wed Mar  9 15:10:43 2011	morphew	3-9	8 AM	3 PM	tiltrotate	rttomo	u2os ltfs	5		
Wed Mar  9 19:01:54 2011	Kirmse	03/09	4:00 PM	7:00 PM	f20cryo	cryoimaging	IFs	0		
Thu Mar 10 17:53:07 2011	Johanna	3/10	10:30 AM	6:00 PM	fishione	rttomo	Chlamydomonas	6		
Fri Mar 11 15:06:40 2011	Johanna	3/11	9:00 AM	3:00 PM	f20cryo	cryotomo	trypanosomes	0		F20 cryo holder crashed vacuum 
twice on exit. 
Sat Mar 12 18:27:04 2011	Johanna	3/12	12:30 PM	6:30 PM	fishione	rttomo	trypanosomes	4		
Mon Mar 14 21:55:57 2011	Johanna	3/14	9:30 AM	10:30 PM	fishione	rttomo	trypanosome collar/connector	16		
Tue Mar 15 18:55:20 2011	eileen	3/15	8:30 AM	7:00 PM	tiltrotate	rttomo	U2OS augmin RNAi	4 3x2		
Wed Mar 16 16:53:06 2011	eileen	3/16	8:45 AM	5:10 PM	tiltrotate	rttomo	CDC20	6 1x2		
Thu Mar 17 15:18:11 2011	morphew	3-17	8 AM	3:15 PM	tiltrotate	rttomo	ltfs u2os	6		
Fri Mar 18 16:45:50 2011	eileen	3/18	10 AM	4:45 PM	tiltrotate	rttomo	cik1 delete Odde	8		
Tue Mar 29 17:16:15 2011	eileen	3/29	8:45 AM	5:15 PM	tiltrotate	rttomo	CDC20	10	seemed ok	
Thu Mar 31 12:48:35 2011	Ron	03/21	8:00 AM	5:00 PM	standard	serviceman	Work on Compustage			Worked on Compustage
Thu Mar 31 12:48:39 2011	Ron	03/22	8:00 AM	5:00 PM	standard	serviceman	Work on Compustage			Worked on Compustage
Thu Mar 31 12:48:44 2011	Ron	03/23	8:00 AM	5:00 PM	standard	serviceman	Work on Compustage			Worked on Compustage
Thu Mar 31 12:48:49 2011	Ron	03/24	8:00 AM	5:00 PM	standard	serviceman	Work on Compustage			Worked on Compustage
Thu Mar 31 12:48:54 2011	Ron	03/25	8:00 AM	5:00 PM	standard	serviceman	Work on Compustage			Worked on Compustage
Thu Mar 31 12:49:03 2011	Ron	03/28	8:00 AM	5:00 PM	standard	serviceman	Work on Compustage			Worked on Compustage
Thu Mar 31 12:49:53 2011	Cindi	03/31	6:30 AM	12:00 PM	tiltrotate	rttomo	MBP-GFP Sections	4		Vacuum at 26 when I came in. Found a bug in 
SerialEM for using exposure time during tilt-
series acquisition.
Fri Apr  8 16:01:27 2011	Spilman	4/1	1:00 PM	7:00 PM	standard	serviceman	Stage issues			
Fri Apr  8 16:01:58 2011	Spilman	4/2	8:00 PM	7:00 PM	standard	serviceman	Stage issues			
Fri Apr  8 16:02:35 2011	Spilman	4/5	8:00 PM	7:00 PM	standard	serviceman	Stage issues			
Fri Apr  8 16:02:43 2011	Spilman	4/6	8:00 PM	7:00 PM	standard	serviceman	Stage issues			
Fri Apr  8 16:02:52 2011	Spilman	4/7	8:00 PM	7:00 PM	standard	serviceman	Stage issues			
Fri Apr  8 16:03:19 2011	Spilman	4/8	8:00 PM	7:00 PM	standard	serviceman	Stage issues			
Mon Apr 11 15:34:28 2011	Cindi	04/11	8:00 AM	10:00 AM	tiltrotate	rttomo	MBP-GFP 1to50	0		vacuum at 38 when arrived. Removed 
holder, dropped to 16. Reinstert 
holder to start pre-pump, dropped 
to 6. Vacuum crept up to 38 at one 
point during routine imaging (with 
valves closed!).
Mon Apr 11 15:34:54 2011	Cindi--tour	04/11	10:00 AM	11:00 AM	tiltrotate	training	tour	0		
Mon Apr 11 15:35:48 2011	Ron Spilman	04/11	11:00 AM	4:00 PM	tiltrotate	serviceman	work on compustage leak	0		pulled stage, lubricated orings, 
adjusted stage. Cindi double-
checked and seems to be working
Tue Apr 12 07:59:32 2011	Cindi	04/12	7:35 AM	8:00 AM	standard	rtimage	shrinkage dataset			vacuum at 6
Tue Apr 12 15:20:58 2011	Spilman	11/12	11:40 AM	3:20 PM	standard	serviceman	Vacuum leak stage		Replace oring with 6.63 from 6.62mm	
Tue Apr 12 17:49:12 2011	eileen	4/12	8:30 AM	5:45 PM	tiltrotate	rttomo	chalmy mitotic	2	vac issues in the morning. Ron 
replaced o-rings and is working 
now. May be intensity callibration 
issues? Used at spot size 1 and it 
was ok.	
Wed Apr 13 07:18:00 2011	Cindi	04/13	7:00 AM	7:15 AM	standard	serviceman	Checking beam brightness		Redid supervisor alignments. Got 
beam to 8.10na. Logged in as myself 
and loaded new alignments. Got 
8.59na. Decided not to reboot. Beam 
is just brighter today.	
Wed Apr 13 11:33:58 2011	eileen	4/13	8:30 AM	11:30 AM	tiltrotate	rttomo	chlamy mitotics	2	intensity up over 8 at spot 2 now. 
Now issues with vac or intensity 
today.	
Wed Apr 13 19:16:44 2011	Janet Meehl Fox/Winey	4/11	1:00 PM	7:30 PM	tiltrotate	rttomo	WT and poc1del tetrahymena basal 
bodies	6		Got an error in serial EM about 
beam intensity so I went to spot 
size 1 and it worked fine.
Thu Apr 14 15:27:25 2011	kirmse	04/14	9:30 AM	3:30 PM	914	cryotomo	IFs	0		really messed up, objective 
aperture was way out, couldn't get 
a e- dose reading at first, beam 
centering is off between images.

might be my alignment file, or the 
reloaded service one and everything 
has to be redone for my account
Mon Apr 18 16:09:40 2011	eileen	4/15	1:30 PM	5:15 PM	tiltrotate	rttomo	tub4-25degrees	3 1x2		
Mon Apr 18 16:10:57 2011	eileen	4/18	8:30 AM	4:10 PM	tiltrotate	rttomo	tub4 25 degrees	3 plus 3 1x2		
Tue Apr 19 16:17:38 2011	chuckr	04/19	8:20 AM	4:15 PM	fishione	rttomo	M barkeri	4	beam marginal (7.5 nA)

Vacuum very nice, very solid...	
Tue Apr 19 18:21:52 2011	Janet Meehl Fox/Winey	4/15	11:15 AM	1:15 PM	tiltrotate	rttomo	poc1del tetrahymena	2		
Thu Apr 21 15:55:11 2011	Cindi	04/21	8:00 AM	4:00 PM	tiltrotate	rttomo	Giardia MBP-GFP ICC Cytos	1	1X2. It took a LONG time to find a 
sample.	David and I redid the beam 
intensity calibrations for spot 2,3 
(on right side of crossover). David 
redid the crossover calibrations.
Fri Apr 22 12:57:33 2011	Cindi	04/22	7:15 AM	1:00 PM	tiltrotate	rttomo	Giardia Immuno-labeled Cytos	3		Vacuum was at 15 the last 3 days. 
Fri Apr 22 18:08:18 2011	Janet Meehl Fox/Winey	4/22	1:00 PM	6:00 PM	tiltrotate	rttomo	WTtetrahymena basal bodies	6		
Sat Apr 23 18:35:17 2011	Johanna	4/23	10:00 AM	6:30 PM	fishione	rttomo	trypanosomes old flagellum	2		
Mon Apr 25 13:25:15 2011	Cindi/David	04/25	8:00 AM	1:00 PM	standard	serviceman	Working out serialEM calibs			alignment seems to be way off. Ron 
said to try a reboot. EFTEM is 
definitely way off.
Mon Apr 25 19:25:44 2011	Cindi	04/25	1:45 PM	7:30 PM	tiltrotate	rttomo	Kin14b GFP Cytoskeletons	8	Had some serious drift/focus 
issues. Redid rotation center and 
it seemed to help a little.	
Tue Apr 26 16:20:37 2011	Chuck R	04/26	8:30 AM	4:20 PM	fishione	rttomo	M barkeri	4	used c2 = 150u = 19 nA beam	none
Tue Apr 26 18:02:28 2011	morphew	4-25	4:15 PM	6:30 PM	tiltrotate	rttomo	ptk			
Wed Apr 27 15:45:57 2011	Ron	04/27	8:00 AM	2:00 PM	standard	serviceman	redo alignments			redid all alignments (including 
EFTEM). New service alignment 
called Service042711
Wed Apr 27 15:46:33 2011	Cindi	04/27	3:00 PM	3:30 PM	tiltrotate	rtimage	check sample			checked alignments and eftem. 
seemed fine. 
Thu Apr 28 21:52:06 2011	Matthew	West	8:00 AM	5:00 PM	tiltrotate	rttomo	yeast ER Mitochondria	8		Pole touches close the valves--
easily corrected, thanks for the 
assistance Cindi!
Thu Apr 28 21:52:52 2011	Matthew	West	5:00 PM	10:00 PM	tiltrotate	rttomo	yeast MVBs	5	Odorizzi Time	
Fri Apr 29 17:27:20 2011	eileen	4/29	8:30 AM	5:30 PM	tiltrotate	rttomo	tub4 25 degrees	6	HT/Feg shut off- message ODP 
Heater. Recycled vacuum and all was 
ok.	HT/FEG shut off, camera vac shot 
up. Column valves closed at the 
time and the column vac not 
affected. Message: ODP Heater. 
Called Ron and he suggested 
recycling vacuum. The system 
recovered and all was fine the rest 
of the day. ?? He is going to order 
another ODP heater.
Sun May  1 09:51:32 2011	chuckr	5/1	9:15 AM	9:50 AM	fishione	rttomo	none	0		peoui.exe would not start... hangs 
at "Loading control panels".
Mon May  9 16:45:23 2011	eileen	5/9	8:30 AM	4:45 PM	tiltrotate	rttomo	WT yeast 25 degrees	4 1x2	scope worked great!	
Tue May 10 15:47:45 2011	Chuck R	05/10	8:30 AM	3:50 PM	fishione	rttomo	M barkeri montage	3	beam was puny for my purposes at 
spot 2,c2 = 100 micron.  Otherwise 
the machine was very stable.	
Wed May 11 06:44:16 2011	Spilman	5/3	8:00 AM	5:00 PM	standard	serviceman	test		Scope intermintently shuts down	
Wed May 11 06:45:23 2011	Spilman	5/3	7:30 AM	12:00 PM	standard	serviceman	test		Scope intermintently shuts down	
Wed May 11 06:46:31 2011	Spilman	5/5	7:30 AM	12:00 PM	standard	serviceman	test		Scope intermintently shuts down	Replace faulty relay
Wed May 11 15:24:47 2011	Janet Meehl Fox/Winey	5/11/11	9:00 AM	3:30 PM	tiltrotate	rttomo	WTtet basal bodies	6		Nanoamp reading was very high (over 
200) but the beam looked normal.  
David checked it out and couldn't 
figure out why it was high.  I 
didn't have any issues.
Fri May 13 17:17:29 2011	eileen	5/13	1:30 PM	5:15 PM	tiltrotate	rttomo	CDC20	4		
Fri May 13 17:17:41 2011	eileen	5/13	1:30 PM	5:15 PM	tiltrotate	rttomo	CDC20	4		
Fri May 13 17:19:37 2011	spilman	5/13	9 AM	1 PM	standard	serviceman	service	0	Ron came in yesterday (5/12) and 
today to fix the screen meter. 
Replaced a board and it is now 
working.	
Mon May 16 17:13:22 2011	eileen	5/16	8:30 AM	5:15 PM	tiltrotate	rttomo	CDC20 	7 1x2		
Tue May 17 17:11:34 2011	Chuck R	05/17	8:30 AM	5:10 PM	fishione	rttomo	M barkeri	4	spot size 2 beam = 6.13 nA to dim 
to use... used spot size 1 beam = 
12.6 nA	No alignments for spot size 1 below 
2000x?  Beam appears as once 
concentrated spot, controls have no 
impact on it.
Wed May 18 16:52:28 2011	Janet Meehl Fox/Winey	5/18/11	9:15 AM	5:00 PM	tiltrotate	rttomo	Tetrahymena WT & poc1Del basal 
bodies	7		None!
Thu May 19 13:41:22 2011	morphew	5-19	8:30 AM	1:30 PM	tiltrotate	rttomo	ptk, chromatin	4		
Fri May 20 17:04:51 2011	eileen	5/20	8:30 AM	5:00 PM	tiltrotate	rttomo	P.pacificus	5 2x2		
Sat May 21 18:05:49 2011	chuck r	05/21	8:55 AM	6:05 PM	fishione	rttomo	M barkeri	4	spot size 1 beam = 12.26 nA	none...
Mon May 23 16:36:29 2011	eileen	5/23	8:30 AM	3:30 PM	tiltrotate	rttomo	P.pacificus	2 2x2		Cindi accidentally dumped power 
while working with UPS. Called 
service and restarted scope from 
wall. The ZEM chiller did not come 
back online after the reboot, so 
lenses aren't working. A ticket was 
opened to fix chiller.
Tue May 24 15:05:12 2011	Aztlano Archuleta	05/24	12:30 PM	3:05 PM	standard	serviceman	ZEM Chiller not running	0	Found ZEM Chiller powered off.  The 
main 6 A fuse was blown.  Replaced 
and watched it run for a few 
minutes.  Looks good but would like 
the temp stabilize before using the 
scope.	
Wed May 25 07:44:59 2011	Atzlano	05/24	10:00 AM	2:00 PM	standard	serviceman	work on ZIM chiller			replaced fuse on zim chiller. 
Wed May 25 07:50:46 2011	Cindi	05/25	7:30 AM	7:45 AM	standard	serviceman	checked beam			beam not very bright ~6na at spot 
2, 20K. 
Wed May 25 14:34:40 2011	morphew	5-25	8 AM	2:30 PM	tiltrotate	cryotomo	chromatin	2		
Thu May 26 19:47:12 2011	Matt West	5/26	8:10 AM	6:00 PM	tiltrotate	rttomo	wt yeast ER	12	Voeltz lab; Scope worked superbly	
Thu May 26 19:51:32 2011	Matt West	5/26	6:00 PM	7:45 PM	tiltrotate	rttomo	wt MVBs with pGO642	1	I was planning 4 more TS, but had 
to end my session early	I had problems with the stage map 
not recognizing the location on the 
section.  Worse, the tracking was 
lost at high tilt near the center 
of the grid for no obvious reason 
even after multiple attempts with 
different cells and realignments.  
I'm stumped...ending my Odorizzi 
session early.  
Thu May 26 19:56:37 2011	Matt West	5/26	6:00 PM	7:45 PM	tiltrotate	rttomo	wt with pGO642	1		Follow-up to problem:  The screw 
was not entirely screwed into the 
holder.  Hopefully no damage was 
done to the threads.  Likely 
explains the short TS of my first 
Odorizzi sample.
Sun May 29 08:20:43 2011	Chuck R	5/28	8:55 AM	6:10 PM	fishione	rttomo	M barkeri	4		
Tue May 31 08:12:29 2011	Chuck R	5/29	8:15 AM	5:30 PM	fishione	rttomo	M barkeri	4		
Wed Jun  1 16:28:01 2011	eileen	6/1	8:30 AM	4:30 PM	tiltrotate	rttomo	augmin U2OS	2 2x3		
Thu Jun  2 09:58:28 2011	Maria-Cindi	6/2	8 AM	10 AM	standard	training	training		maria is learning	
Fri Jun  3 13:09:10 2011	Cindi	06/03	8:00 AM	11:00 AM	standard	serviceman	F30 Down			UPS Failed. Took scope to 'line' power to bypass 
UPS. Scope comes up on standby, but computer 
doesn't work right. Looks like bad video card. 
Called service. Aztlano came to look at it.
Fri Jun  3 16:34:01 2011	Aztlano Archuleta	06/03	11:00 AM	4:45 PM	standard	serviceman	Computer would not turn on.			Found Video card in the CPU was 
bad.  Replaced with one from 
another site.  Ordered one and will 
return that one to the other site.  
Cold started the FEG and slowly 
rasied the HT.  All appears to be 
normal so back to CU.
Sat Jun  4 16:35:07 2011	Chuck R	06/04	8:15 AM	4:45 PM	fishione	rttomo	M barkeri	4	Used 150um C2 and spot size 2 = 
13.18 nA	
Mon Jun  6 17:01:38 2011	eileen	6/6	2:00 PM	5:00 PM	tiltrotate	rttomo	tub4 high mag ends	2	screen meter reading is not correct 
(<< below zero na). Cindi called 
for service.	
Tue Jun  7 16:42:57 2011	eileen	6/7	9:00 AM	4:45 PM	tiltrotate	rttomo	tub4/Vogel and augmin U2OS/Tomoko	2 plus 2 2x3		
Wed Jun  8 12:48:35 2011	Aztlano Archuleta	06/08	8:45 AM	12:45 PM	standard	serviceman	Exposure Time on Screen			Found the wire connecting the large 
screen to the exp. read out circuit 
was broke.  Repaired and checked 
exp from large to small screen.  It 
appears to be working fine.  System 
is back to CU.
Thu Jun  9 16:23:33 2011	chuck r	06/09	11:00 AM	4:30 PM	fishione	rttomo	M barkeri	2	c2 = 150
spot = 2
beam = 10.5 down almost 3 nA from 
previous session with same spog and 
same c2	
Fri Jun 10 10:28:17 2011	cindi	06/10	7:45 AM	10:30 AM	standard	serviceman	Scope was down when I got in			called service and Aztlano walked 
me through bringing scope back up. 
Cold start after vacuum and desired 
states were "ready". brought 
extraction volt. to 4250, started 
HT at 20 and ramped every 10 to 
200, then to 300.
Fri Jun 10 13:07:20 2011	Cindi	06/10	10:45 AM	1:00 PM	standard	serviceman	scope shut down			Scope shut down again. Brought up 
like last time, only with warm 
start.
Fri Jun 10 14:54:53 2011	Aztlano Archuleta	06/10	1:30 PM	2:50 PM	standard	serviceman	Unexplained Powering Off			Tool was running fine when I came 
in but it turned off twice today.  
Checked the relay in mains power, 
K1, that was replaced a few weeks 
ago, was fine.  Checked the +/- 
15v, 24v and 5v power supplies.  
All were fine.  The 75v PS was 
reading 76.7v which is good but the 
center supply is very loud and is 
blowing much hotter air then the 
otehr two.  Ordered a replacement 
75vdc PS PN 5322 695 16042.  I left 
the back and front panels off or 
open to get more air flow.  I also 
am leaving the curtain open on one 
side to get more cooling to the 
mains power rack.  It can be used 
this way for now.  Back to CU.
Sat Jun 11 17:42:07 2011	Chuck R	06/11	8:30 AM	6:45 PM	fishione	rttomo	M barkeri	4	C2 = 150 um
spot 2
beam = 12.01nA	
Mon Jun 13 13:56:30 2011	Spilman	06/13	8:30 AM	2:00 PM	standard	serviceman	Power supply			Wrong power supply ordered. Looked 
for Fan, but couldn't find one. put 
back in old power supply and will 
replace fan tomorrow.
Mon Jun 13 16:52:41 2011	eileen	6/13	2:30 PM	5:00 PM	tiltrotate	rtimage	U2OS mitotics- scanning			
Tue Jun 14 16:22:33 2011	spilman	6/14	12:30 PM	4:00 PM	standard	serviceman	test	0	Replace power supply fan	
Wed Jun 15 18:00:14 2011	DNM	6/15	4:00 PM	6:00 PM	standard	development	fixing stage movement on reading 
montage			
Thu Jun 16 13:30:23 2011	morphew	6-16	8 AM	1:30 PM	tiltrotate	rttomo	tryps	2		
Fri Jun 17 16:42:30 2011	Janet Meehl Fox/Winey	6/17	9:00 AM	5:45 PM	tiltrotate	rttomo	Tetrahymena poc1delete basal bodies	10		
Mon Jun 20 15:01:09 2011	Spilman	6/20	8:30 AM	6:00 PM	standard	serviceman	test	0	PM due	Made electronic and vacuum checks 
as per PM checklist.  Replaced 
filter, back up harddrive,  
replaced apertures checked for 
normal operation
Tue Jun 21 09:22:16 2011	Cindi	06/21	08:45 AM	9:00 AM	standard	serviceman	fixed key on back			key on back was loose. Reseated and now peoui 
works fine.
Tue Jun 21 17:25:08 2011	Chuck R	06/21	8:15 AM	5:45 PM	fishione	rttomo	M barkeri	4	Beam current (c2 = 150 um),
spot 2: 11.57nA	Security key loose prevented peoui 
from starting
Wed Jun 22 14:59:37 2011	morphew	6/22	8 AM	3 PM	tiltrotate	rttomo	u2os, ltfs	4		
Fri Jun 24 15:45:08 2011	Spilman	6/24	10:30 AM	4:00 PM	standard	serviceman	test	0	Replaced C2 and Objective apertures.
apertures look much better.	
Mon Jun 27 10:52:16 2011	Matt West	06/28	10:00 AM	4:45 PM	tiltrotate	rttomo	COS cell ER-Mito Contacts	7	Voeltz lab	
Tue Jun 28 16:18:27 2011	Chuck R	06/28	8:15 AM	4:15 PM	fishione	rttomo	M barkeri	4	beam = 15.36 nA spot 2, c2 ap size 
not known... c2 position = 3	None...  well behaved
Wed Jun 29 20:59:22 2011	Matt West	06/29	8:00 AM	9:00 PM	tiltrotate	rttomo	Plastic COS	14		Had a slow start with a poorly 
placed grid in the tilt-rotate (my 
error), but once corrected worked 
well
Fri Jul  1 17:29:28 2011	Janet Meehl/Winey	7/1	9:30 AM	5:45 PM	tiltrotate	rttomo	WTtet basal bodies	8		Couldn't get full tilt range in the 
negative direction, got to about 53-
56 degrees and the rotation was 
hindered.  But those tilt series 
got to the +60 ok.  Also, it seems 
like the tracking ball that centers 
the beam wasn't working toward the 
end of my session.
Sat Jul  2 12:16:16 2011	DNM	7/2	10:00 AM	12:15 PM	70	serviceman	trying to calibrate mag gradients			Beam shift alignment and beam-image 
alignment were both very screwed 
up.  I had to realign these in 
order to take montages.  Saved 
alignments in Mast070211 with this.
Tue Jul  5 11:15:38 2011	DNM	7/5	10:30 AM	11:20 AM	70	serviceman	Calibrating mag gradients			
Tue Jul  5 17:15:30 2011	eileen	7/5	10:30 AM	5:15 PM	tiltrotate	rttomo	U2OS supermontage setup/tests			
Wed Jul  6 17:29:11 2011	eileen	7/6	8:15 AM	5:30 PM	tiltrotate	rttomo	U2OS Anaphase supermontage	4 3x3 (2 serial 1x2 supermontage)		
Thu Jul  7 16:56:59 2011	cedric	07/07	8:30 AM	5:00 PM	standard	rtimage	virus	0		
Fri Jul  8 17:15:34 2011	eileen	7/8	8:30 AM	5:15 PM	tiltrotate	rttomo	U2OS Anaphase supermontage	1 3x3 plus mapping		
Sat Jul  9 14:35:02 2011	chuckr	07/09	8:30 AM	2:30 PM	fishione	rtimage	M barkeri	0	Beam = 13.79 nA	
Mon Jul 11 17:12:00 2011	eileen	7/11	8:45 AM	5:10 PM	tiltrotate	rttomo	U2OS supermontage	4 3x3 (2 serial 1x2 supermontage)		
Tue Jul 12 15:03:29 2011	Chuckr	07/12	8:45 AM	3:15 PM	fishione	rtimage	M barkeri, Ecoli, Grid mapping	0		Column vacuum very slow to come 
below 15...
Fri Jul 15 17:06:26 2011	eileen	7/15	8:30 AM	5:10 PM	tiltrotate	rttomo	basal body dup	3		
Mon Jul 18 17:01:13 2011	eileen	7/18	8:30 AM	5:00 PM	tiltrotate	rttomo	chlamy basal bodies	3		
Tue Jul 19 11:36:05 2011	chuck r	07/19	8:00 AM	11:30 AM	fishione	rtimage	E. coli		Beam = 12.33 nA	
Tue Jul 19 18:06:44 2011	eileen	7/19	12 PM	6 PM	fishione	rttomo	chlamy basal bodies	7 plus maps		
Wed Jul 20 18:13:57 2011	Kirmse	07/20	9:30 AM	6:00 PM	fishione	rtimage	muscle biopsy	1		
Fri Jul 22 15:12:09 2011	R Spilman	7/22	1:30 PM	3:30 PM	standard	serviceman	Zem chiller shut down	0	Checked zem, added water, open 
curtain a little, monitored ZEM 
temp. Left scope a 200KV over 
weekend to alow objective to run 
cooler over weekend. Can't say yet 
if Zem needs to be replace.  See 
how it does next few days.
	
Fri Jul 22 15:18:08 2011	eileen	7/22	8:45 AM	1:00 PM	fishione	rttomo	chlamy mitotics	4		The objective lens went out- 
chiller failed (30C). Called Ron 
and he turned it off, let it cool 
down and it came back. 
Mon Jul 25 16:59:55 2011	eileen	7/25	8:45 AM	5:00 PM	fishione	rttomo	chlamy	6		
Tue Jul 26 16:10:52 2011	chuckr	07/26	8:40 AM	4:15 PM	fishione	rttomo	M barkeri mapping	3		Even with C2 = 150 um, beam current 
is too low: 9.40 nA.  Had to go to 
spot size 1.

Also... had a column vacuum burp... 
col vac went to 99 after std rod 
removed, but turbo pump stayed on.
VAC System self corrected within 
about 5 minutes... no other issues 
with vaccuum in this session.
Wed Jul 27 16:15:37 2011	eileen	7/27	8:30 AM	4:15 PM	tiltrotate	rttomo	LLCPK mts	2 2x2		
Fri Jul 29 14:52:36 2011	morphew	7/29	8 AM	3 PM	tiltrotate	rttomo	ptk hec1	4 2x1		
Mon Aug  1 15:52:44 2011	eileen	8/1	8:45 AM	3:30 PM	tiltrotate	rttomo	Odde LLCPK	2 2x2		
Mon Aug  8 16:52:54 2011	eileen	8/8	8:30 AM	4:50 PM	fishione	rttomo	Augmin RNAi Tomoko	2x3 montage	The objective lens chiller failed 
again. Ron ordered a new one that 
will be installed tomorrow.	
Wed Aug 10 15:50:41 2011	morphew	8-10	8 AM	3:45 PM	f30cryo	cryotomo	neurons	4		gif was a trouble-maker
Thu Aug 11 14:54:59 2011	Kirmse	08/11	11:00 AM	3 PM	fishione	rtimage	Myosin	0		
Fri Aug 12 16:00:58 2011	eileen	8/13	8:30 AM	4:00 PM	fishione	rttomo	Augmin RNAi Tomoko	3 2x3 montage	took 3 gain refs due to quadrant 
issues.	
Wed Aug 17 16:44:06 2011	eileen	8/17	8:45 AM	4:40 PM	fishione	rttomo	Augmin RNAi Tomoko	2 2x3 montages		
Wed Aug 24 13:04:08 2011	Cindi	08/24	7:00 AM	1:00 PM	fishione	rttomo	WT Septin			lower left quadrant really bad and 
fluctuating. Used GIF. Tuned great, 
but the image shifts depending on 
mag. Need to do alignment.

Thu Aug 25 08:29:54 2011	Cindi	08/25	7:00 AM	8:30 AM	fishione	rttomo	Septin	6	forgot to put TS number yesterday	
Thu Aug 25 13:27:36 2011	Cindi	08/25	8:45 AM	1:30 PM	standard	serviceman	quadrant correction/alignments			Did quadrant correction on upper 
camera. Redid image alignments for 
both cameras. Should stay in 
position while changing mag.
Fri Aug 26 17:45:35 2011	Robert	08/29	4:00 PM	6:00 PM	fishione	rttomo	IFs	3		Was just a small test thats why I 
used the Ultracam, did the 15 frame 
averages for the gain ref, but the 
lower left quadrant is really bad, 
lines and all
Wed Aug 31 16:08:52 2011	morphew	8/31	8 AM	4 PM	f30cryo	cryotomo	neuron whole mount	3		
Tue Sep  6 13:11:22 2011	Cindi/Joey	09/05	8:30 AM	1:00 PM	f30cryo	cryoimaging	Giardia cytos		grids super contaminated	had a real hard time getting vacuum 
to drop after adding holder. Grids 
very contaminated. GIF tuning went 
very long. Had to re-tune several 
times. 
Tue Sep  6 15:17:30 2011	Joanna Duncan	09/06	8:30  AM	3:00 PM	f30cryo	cryoimaging	Giarida cytos	0		No problems that you don't already 
know about.
Thu Sep  8 10:34:41 2011	DNM	9/7	3:30 PM	5:00 PM	standard	serviceman	trying old camera video card			
Thu Sep  8 14:17:51 2011	morphew	9/8	8 AM	2 PM	f30cryo	cryoimaging	neurons			
Fri Sep  9 16:12:50 2011	DNM	9/9	8:45 AM	4:15 PM	standard	serviceman	Swapping camera cards, seems to 
have fixed it			
Mon Sep 12 13:21:38 2011	Joanna Duncan	09/12	8:00 AM	1:17 PM	f30cryo	cryoimaging	Giardia cytos with gold	0		Bad alignment between view and 
record. Trouble finding samples as 
they were not where they appeared 
on the LMM.
Tue Sep 13 11:03:02 2011	morphew	9/13	8 AM	11 AM	tiltrotate	cryoimaging	neurons			"sample rod used" won't affirm
Tue Sep 13 17:32:32 2011	eileen	9/13	1 PM	5:30 PM	fishione	rttomo	tub4 S360
tub4-S360A	6		
Wed Sep 14 17:31:30 2011	Aztlano Archuleta	09/14	8:15 AM	5:30 PM	standard	serviceman	Vacuum Leak Related to Z Axis 
Movement			Replaced the hemisherical bearing o-
rings.  This solved the vacuum leak 
related to Z but caused one with 
alpha.  Replaced the larger bearing 
o-ring with the non-45 degree 
seam.  This solved all vacuum leaks 
related to stage movement but it 
now has a leak when the rod is 
inserted.  Will continue to work on 
this tomorrow.  Dr. Mastronarde 
requested that the stage height be 
returned to the height which gives 
us ~88% on the objective lens in 
focus.  This hopefully will solve 
the image, beam and aperture shifts 
he has been seeing.  Obj is 
currently at 93% and historically 
it was closer to 88%.  The schedule 
has been changed so FEI service can 
have the system until Friday 
afternoon.

Aztlano Archuleta 720 236-8159 cell.
Fri Sep 16 12:29:21 2011	Aztlano Archuleta	09/16	8:30 AM	12:30 PM	standard	serviceman	Compustage height adjustment and 
vacuum leak			Replaced the bearing cup and vacuum 
was fine with and with out the 
holder.  Stage moved to limits with 
no vacuum increases.

Set stage height lower in the 
column.  Now obj. lens current is 
90.8% at focus in eucentric 
height.  Did do alignments but not 
EFTEM.  Back to CU for further 
testing.
Fri Sep 16 15:53:57 2011	Aztlano Archuleta	09/16	2:00 AM	4:00 PM	standard	serviceman	Tilt Axis Adjustment			Adjusted tilt axis.  It looks good 
now.  Back to CU.
Fri Sep 16 17:35:42 2011	Kirmse	09/16	4:00 PM	6:00 PM	standard	rtimage	Diffraction experiment	0		
Wed Sep 21 12:55:13 2011	Cindi	09/21	7:00 AM	1:00 PM	fishione	rttomo	Septins WTB	4		
Thu Sep 22 17:12:14 2011	Spilman-Archeleta	9/22	8:00 AM	5:30 PM	standard	serviceman	alignments			Aligned eftem
Fri Sep 23 16:58:47 2011	eileen	9/20	8:30 AM	4:00 PM	fishione	rttomo	Augmin RNAi Tomoko	2 2x3		
Fri Sep 23 16:59:24 2011	eileen	9/20	10 AM	5:00 PM	fishione	rttomo	bld2_5 bbs	7		
Mon Sep 26 16:55:21 2011	eileen	9/26	10 AM	5 PM	fishione	rttomo	bld2-5 	6		
Tue Sep 27 17:05:14 2011	eileen	9/27	8:30 AM	5:10 PM	tiltrotate	rttomo	C.elegans metaphase supermontage	1x2 supermontage of 3x3  montages		
Wed Sep 28 18:50:30 2011	spilman	9/28	8:40 AM	6:55 PM	standard	serviceman	align	0		ETEM Alignements
Fri Sep 30 10:24:57 2011	Joanna	09/29	8:30 AM	1:00 PM	f30cryo	cryoimaging	Giarid gold controls	0		FEM alignment so far off it was not 
usable. :( Digital micrograph 
crashed twice. 
Fri Sep 30 14:17:16 2011	cindi	09/30	8:00 AM	11:15 AM	fishione	rttomo	2738B septins	2	punched a hole in the grid when 
rotation :(	
Sat Oct  1 15:41:21 2011	Johanna/Per	10/01	08:00 AM	03:45 PM	f30cryo	cryoimaging	In vitro microtubules			
Sun Oct  2 22:39:17 2011	Johanna/Per	10/02	11:30 AM	10:45 PM	f30cryo	cryoimaging	In vitro MTs			LMMs are all off, as are where 
record and autofocus is in 
comparison to where you aim. 
Tue Oct  4 12:07:27 2011	Chuck Moe	10/04	08:30 AM	12:00 PM	standard	serviceman	EFTEM alignment			Checked EFTEM alignments and 
verified successful GIF tuning in 
supervisor and customerservice 
workspaces
Wed Oct  5 09:44:18 2011	cindi	10/5	7:00 AM	9:45 AM	standard	serviceman	figuring out why account bad			Alignments looked good in 
supervisor, mary's account. Bad in 
mine. Turns out I was just way out 
of focus in EFTEM mode, causing the 
beam to be super out of alignment. 
I fixed focus and all is fine.
Wed Oct  5 11:23:09 2011	Cindi	10/5	9:45 AM	11:25 AM	tiltrotate	rttomo	2738B	2		
Thu Oct  6 14:03:55 2011	Johanna	9/6	10:00 AM	11:00 AM	f30cryo	cryoimaging	In vitro Mts			
Fri Oct  7 12:56:20 2011	Spilman	10/07	11:00 AM	1:00 PM	standard	serviceman	ZEM chiller	0		ZEM chiller failed, reset & adjusted
Fri Oct  7 14:35:54 2011	DNM	10/7	1:15 PM	2:30 PM	standard	development	switching SerialEM to mercurial			
Mon Oct 10 14:01:03 2011	Charles Moe	10-10	09:00 AM	2:00 PM	standard	serviceman	Water Flow problem			ZEM chiller is not working. Re-
routed hoses for Haskras Chiller to 
cool OBJ lens temporarily untill 
new ZEM chiller arrives.
Tue Oct 11 18:40:34 2011	Eileen	10/11	8:30 AM	6:40 PM	tiltrotate	rttomo	C.elegans supermontage with Steffi	3 2x3	Stage moved by itselft- Ron 
rebooted the comupter and it fixed 
the problem.	
Wed Oct 12 18:21:27 2011	eileen	10/12	8:30 AM	6:20 PM	tiltrotate	rttomo	C.elegans anaphase supermontage/ 
Steffi	6 2x3	file transfer very slow	
Thu Oct 13 14:39:07 2011	Joanna Duncan	10/13	8:30 AM	2:35 PM	f30cryo	cryoimaging	Giarida cytos	0		
Fri Oct 14 17:54:30 2011	DNM	10/14	10:30 AM	6:30 PM	standard	development	Recalibrate IS offsets, beat head 
against realign to item	0		
Sun Oct 16 17:45:24 2011	DNM	10/16	1:15 PM	5:45 PM	70	development	Testing realign to item			
Mon Oct 17 16:50:18 2011	eileen	10/17	8:45 AM	4:50 PM	fishione	rttomo	tub4-360A Vogel/Winey	6		
Tue Oct 18 13:24:35 2011	Joanna	10/18	9:30 AM	1:21 PM	f20cryo	cryoimaging	Not successful 	0		
Tue Oct 18 16:32:23 2011	eileen	10/18	2:00 PM	4:30 PM	fishione	rttomo	tub4-S360A Vogel/Winey	4		
Wed Oct 19 09:57:36 2011	Spilman	10/19	7:45 AM	10:00 AM	standard	serviceman	none			Installed and adjusted ZEM chiller
Wed Oct 19 14:55:44 2011	eileen	10/19	11:00 AM	3:00 PM	fishione	rttomo	tub4S360A Vogel/Winey	4		
Thu Oct 20 13:30:09 2011	Joanna Duncan	10/20	8:00 AM	1:00 PM	f30cryo	cryotomo	Tilt series of WT cytos	2		
Mon Oct 24 16:39:25 2011	eileen	10/24	1 PM	4:40 PM	fishione	rttomo	DeCamilli Dynamin 1/3 KO K+	5		
Tue Oct 25 12:56:38 2011	Kristin	10/25	9:30 AM	1:00 PM	fishione	rttomo	lens capsule	2		
Tue Oct 25 17:27:34 2011	eileen	10/25	1:00 PM	5:30 PM	fishione	rttomo	DeCamilli dynamin1/3ko and wt yeast	4		
Wed Oct 26 12:52:36 2011	Cindi	10/26	6:30 AM	12:50 PM	f30cryo	cryotomo	Giardia WT Cytoskeletons	2	Had some drift	Power went out hardcore at 7AM, but 
scope did fine with UPSs.
Thu Oct 27 13:10:47 2011	Cindi	10/27	6:30 AM	1:06 PM	f30cryo	cryotomo	F30 WT giardia cytos	0		drift really bad at high tilt. 
Turns out the aperture is affecting 
drift with some sort of charging. 
No charging near grid bar, no 
charging with standard holder at 
high tilt. Unclear what the issue 
is. FEI needs to come and redo tilt-
axis offset and z-height.
Fri Oct 28 16:15:23 2011	eileen	10/28	8:30 AM	4:15 PM	fishione	rttomo	tub4-S360A and wt Vogel/Winey	7		
Mon Oct 31 16:09:26 2011	Spilman	10/31	8:30 AM	4:30 PM	f30cryo	serviceman	Compustage			Adjusted V groove and aligned 
stage. Checked stage with single 
tilt and F20 cryo holders which 
worked fine.  Found F30 cryo holder 
to be dirty.  cleanend holder & 
checked for normal operation. cold 
& warm.
Tue Nov  1 19:01:44 2011	eileen	11/1	8:30 AM	7:00 PM	fishione	rttomo	tub4-360A Vogel/Winey	14		
Thu Nov  3 13:04:44 2011	Joey	11/03	8:30 AM	1:01 PM	f20cryo	cryotomo	Tilt series WT cytos	1		Messed up the vacuum-ooops! 
Fri Nov  4 13:19:21 2011	morphew	11-5	8 AM	1 PM	tiltrotate	rttomo	neurons	2-3x1		
Mon Nov  7 12:50:36 2011	Cindi	11/07	9:30 AM	12:45 PM	70	rttomo	2738ELG	4		
Tue Nov  8 15:42:44 2011	eileen	11/8	8:30 AM	3:40 PM	tiltrotate	rttomo	cardiac myocytes Tom (Kohl lab)	2		
Thu Nov 10 15:08:55 2011	Joey	11/10	8:00 AM	3:00 PM	f30cryo	cryotomo	TS WT giardia cytos	2		As mentioned previously.
Thu Nov 10 15:19:54 2011	Joey	11/10	8:00 AM	3:00 PM	f30cryo	cryotomo	cytos 	2		During TS, in view before second 
image beam was lost. Only about to 
find it again in LM. Fixed by 
pressing Reset Beam in TUI. Managed 
to finish TS.
Fri Nov 11 12:35:04 2011	morphew	11-11	8 AM	12:30 PM	f30cryo	cryoimaging	hela-correlativeLM/EM			stage moved to extreme position 
when touched the trackbal (or 
multifunction knobs).  Re-booted 
computer, fixed the problem.
Mon Nov 14 13:10:47 2011	Cindi	11/14	6:30 AM	1:15 PM	f30cryo	cryotomo	Giardia WT cytos	4	low mag obj should be ~7.9%	
Mon Nov 14 17:11:21 2011	eileen	11/14	1:30 PM	5:10 PM	tiltrotate	rttomo	H30 WT Tomoko	2 2x2		
Tue Nov 15 18:07:42 2011	eileen	11/15	3:15 PM	6:05 PM	tiltrotate	rttomo	Tomoko H30 wt	2 2x2		
Wed Nov 16 12:53:51 2011	Cindi	11/16	8:40 AM	12:50 PM	tiltrotate	rttomo	Septins 2738ELog	4		
Wed Nov 16 15:29:32 2011	morphew	11/16	1:30 PM	3:30 PM	f20cryo	cryoimaging	neurons			
Thu Nov 17 16:08:25 2011	Joey	11/17	11:00 AM	4:00 PM	f30cryo	cryoimaging	WT cytos	0		
Fri Nov 18 16:48:12 2011	eileen	11/18	8:30 AM	4:45 PM	tiltrotate	rttomo	tub4_S360A	9		
Mon Nov 21 14:28:40 2011	morphew	11/21	9 AM	2:30 PM	f30cryo	cryotomo	neuron	3		
Tue Nov 22 15:35:42 2011	eileen	11/22	8:30 AM	3:30 PM	tiltrotate	rttomo	Celegans anaphase sec16 b axis	3x1 supermontage of 3x2 montages		
Tue Nov 29 18:06:17 2011	eileen	11/29	8:30 AM	6:00 PM	tiltrotate	rttomo	tub4 S360A Vogel/Winey	10		
Wed Nov 30 13:28:16 2011	Kristin	11/30	8:30 AM	1:30 PM	tiltrotate	rttomo	lens capsule	2		
Thu Dec  1 12:46:22 2011	Cindi	12/01	6:30 AM	1:00 PM	tiltrotate	rttomo	septin correlative work	5		
Fri Dec  2 11:39:54 2011	Cindi	12/2	7:00 AM	11:45 AM	tiltrotate	rttomo	correlative septins	5	beam crazy bright (16 na)	
Mon Dec  5 12:10:13 2011	Joanna Duncan	12/05	8:30 AM	12:00 PM	f30cryo	cryoimaging	Viewed giardia cytos	0		
Tue Dec  6 17:55:00 2011	Matthew West	12-6	9:30 AM	6:00 PM	tiltrotate	rttomo	MIM mutants	8	Odorizzi lab	Stage controller is sticky in one 
spot
Thu Dec  8 16:12:49 2011	morphew	12/8	8 AM	4 PM	f30cryo	cryotomo	neurons	4		
Fri Dec  9 14:24:28 2011	Linda	12/9	8:30 AM	2:30 PM	f30cryo	training	streptomyces cryo		My account has a problem with 
intensity automatic control. And 
teh sample was too thick with 
bacteria	
Mon Dec 12 15:08:41 2011	morphew	12/12	8 AM	3 PM	f20cryo	cryoimaging	neuron			
Tue Dec 13 18:36:59 2011	eileen	12/13	9:00 AM	6:30 PM	tiltrotate	rttomo	tub4-S360A Vogel	6	Big dewar still had nitrogen after 
9.5 hours	
Wed Dec 14 15:33:21 2011	Spilman	11/14	8:00 AM	4:00 PM	standard	serviceman	test	0		PM due, Made electronic & vac 
checks as per PM checklist. Replace 
ODP oil, water filter, Vacuum hose, 
drain compressor & added oil.  
Added water to Haskris chiller, 
Checked beam & image for normal 
operation.
Thu Dec 15 16:37:59 2011	Spilman	11/15	1:00 PM	5:00 PM	standard	serviceman	test	0		Removed compustage cleaned & lubed 
orings.  Reinstalled, aligned and 
tested.  Ordered new set of orings. 
Left scope logged on and vacuum 
logger on over night.  Will return 
Friday to check status.
Fri Dec 16 08:00:31 2011	Spilman	12/16	7:15 AM	8:15 AM	standard	serviceman	test	0		Checked vacuum logger (OK) Tilted 
sample +&- 79 degrees with Y at +&- 
500.  Moved stage to multiple 
limits.  No vacuum leaks found.
Fri Dec 16 19:56:10 2011	Linda Sandblad	12/16	10 AM	8 PM	f30cryo	cryotomo	Streptomyces	5		
Mon Dec 19 15:20:49 2011	Cindi	12/19	8:00 AM	3:20 PM	f30cryo	cryotomo	Giardia WT cytoskeletons	7		column had higher vacuum when 
shutter was opened on holder. It 
took hours to get to 6 from around 
20. Compustage might need tweaking. 
Optical axis 5um from tilt-axis in 
extreme stage.
Tue Dec 20 21:55:39 2011	Linda Sandblad	12/20	10 AM	10 PM	f30cryo	cryotomo	cryo keratin	7		
Thu Dec 22 00:56:15 2011	Linda	12/21	12 AM	12 PM	f30cryo	cryotomo	FilP cryo	9		
Fri Dec 23 18:01:05 2011	Linda	12/23	10 AM	6 PM	f30cryo	cryoimaging	keratin			
Mon Dec 26 20:45:20 2011	Linda	12/26	1 PM	9:00 PM	f30cryo	cryotomo	Keratin	1		Was stucked in Cryo cycle since 
friday, restart, OK
Tue Dec 27 17:33:20 2011	eileen	12/27	9:00 AM	5:30 PM	tiltrotate	rttomo	Tetrahymena conjugation	8		
Wed Dec 28 15:49:11 2011	Cindi & Garry	12/18	9:00 AM	3:39 PM	f20cryo	cryotomo	Training and giardia cytos	1		
Thu Dec 29 18:26:21 2011	Linda	12/92	10 AM	18:30 PM	f30cryo	cryotomo	keratin	1		
Fri Dec 30 14:42:22 2011	morphew	12/29	8 AM	2:30 PM	f30cryo	cryotomo	neurons	4		
Sat Dec 31 01:41:47 2011	Linda	12/30	4 PM	2 AM	f30cryo	cryotomo	FilP	7		
Mon Jan  2 13:42:38 2012	morphew	1/2	9 AM	1:30 PM	f30cryo	cryoimaging	LM/EM tests			
Tue Jan  3 16:50:48 2012	eileen	1/3	9:00 AM	4:50 PM	tiltrotate	rttomo	scanning RPE1 mitotics	1		
Wed Jan  4 15:26:37 2012	morphew	1-4	8 AM	3:30 PM	f30cryo	cryotomo	neuron	4		
Thu Jan  5 14:33:15 2012	morphew	1/5	9 AM	2:30 PM	tiltrotate	rttomo	jon martin neurons	2 2x2		vacuum goes up w/o rod in.
Tue Jan 10 17:09:02 2012	Byung-Ho Kang	01/10	8:00 AM	5:00 PM	fishione	rttomo	Arabidopsis cotyledon	16	I did not sign yesterday (1/9)
time in 9:30, time out 6:00	vacuum failed on 9th Jan when the 
compustage move up to 60oC. Turned 
off according to Cindi's direction
Thu Jan 12 13:36:23 2012	Spilman	1/11	11:00 AM	5:30 PM	standard	serviceman	yest	0		Replaced compustage orings, bearing 
cup & ball as needed.  Cleaned & 
lubricated.  Aligned and checked 
for normal operation.
Thu Jan 12 13:37:11 2012	Spilman	1/12	8:00 AM	1:45 PM	standard	serviceman	yest	0		Replaced compustage orings, bearing 
cup & ball as needed.  Cleaned & 
lubricated.  Aligned and checked 
for normal operation.
Thu Jan 12 15:45:08 2012	morphew	1-12	2 PM	3:45 PM	standard	rtimage	rpe1			
Fri Jan 13 16:49:23 2012	Joey	01/13	9:00 AM	4:45 PM	f20cryo	cryotomo	tilt series of WT giardia cytos	4		
Tue Jan 17 18:09:31 2012	eileen	1/17	11:00 AM	6:05 PM	tiltrotate	rttomo	Vogel WT yeast	6		
Wed Jan 18 16:36:07 2012	Cindi	01/18	8:00 AM	10:00 AM	70	rtimage	Giardia supermontage			couldn't do a supermontage b/c the 
movement in X was bad. Aztlano 
worked on it.
Wed Jan 18 16:44:50 2012	Aztlano Archuleta	1/18	10 AM	4:45 PM	tiltrotate	serviceman	X drift			Found the X encoder sticking a 
little.  Cleaned off and verified 
it was smooth.  Also found the V-
Groove was misaligned.  Aligned the 
V-Groove but it still needs some 
slight adjustments.  Will be back 
in tomorrow for that adjustment.
Thu Jan 19 11:05:00 2012	Aztlano Archuleta	01/19	9:00 AM	11:00 AM	f30cryo	serviceman	V Groove Adjustment			Adjusted V Groove with the beta 
tilt/cryo holder.  Eucentric height 
adjusts great at -74um and has very 
little shift at +/- 60 degrees.  
When using SerialEM to move the 
sample it is with in spec less than or = 500nm 
in both directions.  Back to CU.
Thu Jan 19 15:01:59 2012	morphew	1-19	12:30 PM	3 PM	f20cryo	cryotomo	neuron	3		
Fri Jan 20 17:30:29 2012	eileen	1/20	8:30 AM	5:30 PM	tiltrotate	rttomo	Pacificus anaphase	4 2x2 montage		
Mon Jan 23 15:17:32 2012	Kristin	1/23	8:15 AM	3:15 PM	tiltrotate	rttomo	lens capsule	3	stage got completely stuck... 
luckily Ron was here to take it 
apart and get it moving again	
Tue Jan 24 12:24:20 2012	Joey	01/24	9:20 AM	12:20 PM	f20cryo	cryoimaging	Did not get any done	0		Compustage was not functioning 
correctly. It was not able to move 
the stage properly when I tried to 
take a lmm. Cedric helped me take 
out the rod and try to reset the 
stage position but it did not help. 
It seems to be having trouble 
finding the centre of the stage and 
then moving away from there. 
Wed Jan 25 16:39:04 2012	eileen	1/25	9:00 AM	4:30 PM	tiltrotate	rttomo	P.pacificus	4 2x2 montages	No stage issues!	
Wed Jan 25 16:39:47 2012	eileen	1/25	9:00 AM	4:30 PM	tiltrotate	rttomo	P.pacificus	4 2x2 montages	No stage issues!	
Thu Jan 26 14:39:49 2012	Joey	01/26	9:15 AM	2:30 PM	f30cryo	cryoimaging	None	0		Lots of problems again today. Stage 
is is still not working properly. 
Seems to be confused about its 
position and unable to move 
properly in the X direction. Also 
had problem with the GIF and 
ultracam. Could not get focuses or 
eucentric easily. Could not get the 
GIF to refine the zero loss peak. I 
am not sure if I should have 
completely retuned the GIF to fix 
this. Also, some of the settings on 
my account are now off from what 
they should be from the efforts of 
various people to fix it. 
Fri Jan 27 14:05:48 2012	morphew	1-27	8 AM	2 PM	f20cryo	cryotomo	neurons	4		
Mon Jan 30 16:54:58 2012	eileen	1/30	8:45 PM	5:00 PM	fishione	rttomo	chalmy bbs	6		
Tue Jan 31 15:00:47 2012	morphew	1-31	8 AM	3 PM	f20cryo	cryotomo	neurons	4		
Wed Feb  1 13:46:03 2012	eileen	2/1	8:30 AM	1:45 PM	fishione	rttomo	chlamy bbs	4		
Thu Feb  2 17:37:51 2012	Spilman	2/2	10:45 AM	5:45 PM	standard	serviceman	cal/spheres	0		Went through special alignments not 
satisfied with results more time 
needed.
Mon Feb  6 15:26:39 2012	morphew	2/6	8 AM	3:30 PM	f20cryo	cryotomo	neuron	2		
Tue Feb  7 19:20:50 2012	eileen	2/7	2:30 PM	7:20 PM	tiltrotate	rttomo	Vogel wt	4		
Wed Feb  8 07:46:36 2012	Cindi	02/07	7:30 AM	1:30 PM	standard	serviceman	Cryo cycle stuck			found cryo-cycle stuck. With advice 
from Ron, tried rebooting. Had to 
manually turn on IGP pumps after 
vacuum cycle started. HT couldn't 
come on. Ron came in and worked on 
it.
Wed Feb  8 08:44:41 2012	Cindi	02/08	7:30 AM	8:45 AM	standard	serviceman	none			Stuck in cryo-cycle. Called Ron: 
Close V6, turn off turbo. Reboot 
computer, after TEM server starts, 
start IGP2,1,4. Get out of manual 
mode! Start FEG, enable vacuum on 
scope. Start warm cycle. Turn on 
HT. Bring up HT and extraction 
voltage when warm cycle complete.
Thu Feb  9 13:20:16 2012	Spilman	2/7	3:00 PM	4:00 PM	standard	serviceman	none	0		Restarted pumping system after cryo 
cycle failed to complete
Thu Feb  9 13:22:34 2012	Spilman	2/9	8:00 AM	9:30 AM	standard	serviceman	none	0		Found P15 gage failed to read 
correct vacuum causing cryocycle 
failure. Tapped on gage to knock 
off whisker/dirt and scope worked 
fine.
Fri Feb 10 17:52:09 2012	Joey	02/10	9:15 AM	5:45 PM	f20cryo	cryotomo	wt  CYTOS	6		
Mon Feb 13 17:01:36 2012	eileen	2/8	8:30 AM	4:00 PM	tiltrotate	rttomo	Vogel wt yeast	8		
Mon Feb 13 17:01:53 2012	eileen	2/8	8:30 AM	5:00 PM	tiltrotate	rttomo	Vogel wt yeast	9		
Mon Feb 13 17:02:09 2012	eileen	2/13	8:30 AM	5:00 PM	tiltrotate	rttomo	Vogel wt yeast	9		
Tue Feb 14 08:23:24 2012	Cindi	02/14	8:00 AM	8:20 AM	standard	serviceman	stuck in cryo-cycle			tapped P15 guage and the cryo-cycle 
finished. guage needs replacing.
Tue Feb 14 14:17:20 2012	Spilman	2/14	8:00 AM	2:30 PM	standard	serviceman	gold calibration	0		Took head of new p15 gage and 
installed electonics of old gage to 
make p15 work.  Completed special 
alignments.
Wed Feb 15 16:35:49 2012	Joey	02/15	9:15 AM	4:31 PM	f20cryo	cryotomo	WT cytos	3		
Thu Feb 16 10:11:30 2012	morphew	2/16	8 AM	10 PM	tiltrotate	cryotomo	rpe1	1		
Thu Feb 16 15:31:25 2012	Joey	02/16	10:10 AM	3:30 PM	f20cryo	cryotomo	TS WT cytos	3		
Fri Feb 17 13:59:09 2012	morphew	2-17	8 AM	2 PM	tiltrotate	rttomo	rpe1	1		
Fri Feb 17 15:15:30 2012	cindi	02/17	2:45 PM	3:15 PM	standard	serviceman	PVP pump failure			Vacuum crashed during cryo-cycle. 
Restarted vacuum and had buffer 
tank error. Called Ron and got 
column under vacuum. Looks like 
buffer tank pump has failed and can 
no longer pull vacuum.
Mon Feb 20 14:43:55 2012	Aztlano Archuleta	2/20	10:30 AM	3:00 PM	standard	serviceman	Projection PVP Vacuum Error			Replaced the PVP and still had high 
vac ~46 log/ 1 torr.  Found the 
rubber hose to the buffer tank was 
cracked and when moved the pressure 
will spike up and you could hear 
the leak.  Replaced but still had 
high leak of 1 torr/min.  Blanked 
off V1 and still had high 
pressure.  Ordering V1.
Wed Feb 22 11:09:28 2012	Service	02/21	8:00 AM	5:00 PM	standard	serviceman	scope down bad V1			V1 needs replacement. Aztlano accidentally 
ordered the wrong part (used FEI V1, when he 
should have used Edwards V1). 
Thu Feb 23 15:15:37 2012	Aztlano Archuleta	2/23	8:30 AM	3:30 PM	standard	serviceman	Vacuum Problems			Replaced V1 and the ODP dipstick o-
rings.  Pumped the system down and 
checked the beam.  Looks good so 
back to CU.
Thu Feb 23 18:01:40 2012	DNM	2/24	5:30 PM	6:00 PM	standard	development	looking at rotation on GIF 220 x 
for handedness issue Jensen lab			
Fri Feb 24 12:06:39 2012	Joey	02/24	9:15 AM	12:01 PM	f30cryo	cryotomo	Cryo	0	I let a little air into the column 
and crashed the vacuum, faily minor 
and it was able to recover quickly. 
Had a problem getting the rod to go 
in the last bit when loading the 
cryo rod. 	
Mon Feb 27 16:53:38 2012	Cindi	02/27	6:45 AM	4:48 PM	f30cryo	cryotomo	WT cytos	9		Stage seemed to be far out of 
alignment when up at high tilt at -
60. Had to take it there manually 
and then realign frequently during 
TS. This was improved by starting 
the ts at +60 instead. 
Tue Feb 28 17:02:28 2012	eileen	1/28	9 AM	4:30 PM	tiltrotate	rttomo	Tomoko Augmin U2OS	3 2x3 montage		
Tue Feb 28 17:03:28 2012	David/Eileen	1/28	4:30 AM	5:00 PM	standard	rttomo	alignments	0		
Wed Feb 29 14:31:00 2012	morphew	2/29	7:45 AM	2:30 PM	f30cryo	cryotomo	neurons	0		
Wed Feb 29 17:57:35 2012	DNM	2/29	2:30 PM	6:00 PM	standard	serviceman	Trying to work with ill-fated 
aperture alignment, checking out 
GIF camera			
Thu Mar  1 12:36:08 2012	Aztlano Archuleta	3/01	08:30 AM	1:00 PM	standard	serviceman	Drift in (-) Alpha			Adjusted V-Groove.  Drift has been 
improved.  Please check and let me 
know if it needs additional 
adjustments.
Fri Mar  2 17:20:17 2012	eileen	3/2	8:45 AM	5:15 PM	tiltrotate	rttomo	Tomoko/Goshima augmin RNAi	3 2x3 montage		
Mon Mar  5 15:24:40 2012	morphew	3/5	7:30 AM	3:20 PM	f30cryo	cryoimaging	neuron			
Tue Mar  6 16:45:08 2012	eileen	3/6	8:45 AM	4:45 PM	tiltrotate	rttomo	C.elegans supermontage	1 2x3 and 1x3 supermontage of 2x3 montages		
Wed Mar  7 16:23:17 2012	cindi	03/07	7:30 AM	1:30 PM	f20cryo	cryotomo	giarida cytos	2		found F20 cryo holder pin slightly 
bent. Tested a -4um and -6um tilt-
series to see if ultracam ok (still 
unclear). GIF drifting consistenly 
to minus side (about -7eV over 45 
min).  
Thu Mar  8 15:19:13 2012	morphew	3/8	8 AM	3:15 PM	fishione	rttomo	ltfs giardia	4		
Fri Mar  9 13:13:53 2012	morphew	3/9	8 AM	1 PM	fishione	rttomo	ltfs giardia	4		
Fri Mar  9 18:00:39 2012	eileen	3/9	1:30 PM	6:00 PM	tiltrotate	rttomo	Tomoko normal control	1 3x3 montage and grid scans		
Tue Mar 13 17:08:42 2012	eileen	3/13	8:45 AM	5:00 PM	tiltrotate	rttomo	Tomoko wt- H51 serials	3 3x3 montage		
Wed Mar 14 15:28:09 2012	eileen	3/14	8:30 AM	3:30 PM	tiltrotate	rttomo	Celegans supermontage	3 2x3		beam seemed dim at spot 2 (~4ish)
Thu Mar 15 13:11:20 2012	morphew	3/15	8 AM	1 PM	tiltrotate	rttomo	ltfs chlamy	4		
Thu Mar 15 16:43:31 2012	DNM	3/15	1:00 PM	4:45 PM	70	serviceman	mostly futile GIF testing			
Fri Mar 16 17:58:41 2012	eileen	3/16	8:45 AM	6:00 PM	tiltrotate	rttomo	Celegans anaphase supermontage	6 2x3 montages		
Mon Mar 19 16:19:22 2012	eileen	3/19	8:30 AM	4:15 PM	tiltrotate	rttomo	Celegans anaphase supermontage	3 2x3 montage		
Mon Mar 19 19:53:42 2012	DNM	3/19	4:30 PM	7:50 PM	standard	development	testing autoalign changes and nav
correction for shrinkage in realign			
Tue Mar 20 16:41:55 2012	eileen	3/20	8:30 AM	4:40 PM	fishione	rttomo	DeCamilli WT_K_1min	8		Scope was giving stage end of range 
messages, but was in the middle 
when I logged on. Stage got stuck 
when doing eucentricity. Chuck 
reanabled the stage but it gave the 
message again when inserting rod. 
David said this happened to him 
last night. The scope worked fine 
and I only got the stage message 
one time after stage was enabled.
Wed Mar 21 15:57:35 2012	morphew	3/21	8 AM	4 PM	tiltrotate	rttomo	rpe1	5		
Thu Mar 22 18:05:26 2012	Janet Meehl Fox/Winey	3/22	12:40 PM	6:10 PM	tiltrotate	rttomo	poc1Delete tetrahymena basal bodies	6		
Fri Mar 23 11:11:23 2012	eileen	3/23	8:30 AM	9 AM	standard	rtimage	no work done- zem chiller out again	0		Zem chiller is out again. Service 
called.
Tue Mar 27 17:30:05 2012	Aztlano Archuleta	03/27	9:30 AM	5:30 PM	standard	serviceman	Zepher Zem 300 to Hasker Replacement			Replaced the ZEM with a new Haskers 
R050 chiller.  Ran into a few 
problems with flow and blocked 
lines.  Those were resolved but now 
the PVP will not pump down.  
Getting a spare and gauges to 
verify if the pump is ok or an 
internal leak.  System is still 
down.
Wed Mar 28 17:09:36 2012	Aztlano Archuleta	03/28	9:00 AM	5:00 PM	standard	serviceman	Haskers Install			Replaced the ODP O-Rings and 
restarted TEM Vac.  All is looking 
fine there.  Spent the rest of the 
day pump purging the Haskers to get 
rid of air in the water lines.  It 
is currently running with vacuum.  
Will check tomorrow and finish the 
Haskers ZEM swap.  System is still 
down.
Thu Mar 29 10:27:30 2012	Aztlano Archuleta	03/29	9:00 AM	10:30 AM	standard	serviceman	ZEM To Haskers Replacement			Objective lens water still has 
bubbles so repumping system and 
continuing to purge air out of the 
water.  Will check back later.  
System is still down.
Thu Mar 29 18:05:09 2012	Aztlano Archuleta	3/29	5:30 PM	6:00 PM	standard	serviceman	Haskers R050			Pumped the reservoir again with a 
little more air coming out.  Will 
leave overnight and check 
tomorrow.  System is down.
Fri Mar 30 14:53:54 2012	Aztlano Archuleta	3/30	9:00 AM	3:00 PM	standard	serviceman	Haskers R050			Spoke with Haskers and it sounds 
like we have multiple leaks in the 
water tank.  This was most likely 
caused in shipping.  Haskers is 
sending me a replacement tank.  
System is down.
Tue Apr  3 10:41:02 2012	Service	04/02	8:00 AM	5:00 PM	standard	serviceman	Scope down for bad chiller			Still down. Waiting for water tank on chiller.
Wed Apr  4 10:13:35 2012	Archuleta-Spilman	4/3	1:00 PM	4:00 PM	standard	serviceman	none			Resealled fittings of Haskriss 
chiller vacuum valve. and pumped 
over night.
Wed Apr  4 10:15:33 2012	Archuleta-Spilman	4/3	8:30 PM	11:00 AM	standard	serviceman	none			Switched on objective lens supply 
checked beam & lens currents. 
Adjusted wate flow.  Checked for 
normal operation.  Prepare ZEM for 
return.
Wed Apr  4 17:22:45 2012	eileen	4/4	11:00 AM	5:20 PM	tiltrotate	rttomo	Chlamy anaphase	4 1x2 montage		
Thu Apr  5 14:55:57 2012	morphew	4-5	8 AM	3 PM	f30cryo	cryotomo	neurons	3		
Fri Apr  6 16:30:27 2012	eileen	4/6	8:45 AM	4:30 PM	fishione	rttomo	DeCamilli WT K	8		
Mon Apr  9 10:56:06 2012	cindi	04/09	8:15 AM	10:45 AM	fishione	rttomo	Giardia in HM20	2		Looked at gain ref of Ultracam. Oil 
spots seem to be mostly gone. 
Tue Apr 10 14:18:43 2012	Spilman	4/10	8:00 AM	2:20 PM	standard	serviceman	none			Made electronic & vacuum checks, 
backed harddrive, saved align 
files, Tecnai configuration files, 
and fingerprint files.  Replaced 
vacuum hoses, and checked scope for 
normal operation.
Wed Apr 11 14:56:53 2012	morphew	4/11	8 AM	3 PM	f30cryo	cryotomo	neurons	5		
Fri Apr 13 10:42:03 2012	Cindi	04/13	8:00 AM	10:30 AM	f30cryo	cryoimaging	delta-giardin gfp cytos			crashed vacuum removing F30 holder 
second time. Also, holder couldn't 
get below 13.
Fri Apr 13 15:30:07 2012	cindi	04/13	10:15 AM	3:30 PM	tiltrotate	rttomo	giardia sections	5	sample looks like crap :(	
Tue Apr 17 13:30:00 2012	Cindi	04/17	7:30 AM	1:30 PM	f30cryo	serviceman	testing GIF			
Tue Apr 17 17:28:39 2012	cedric	04/17	4:00 PM	5:30 PM	f30cryo	cryoimaging	nothing	0		
Mon Apr 23 18:03:49 2012	DNM	4/23	10:15 AM	6:00 PM	70	development	testing realign to item, trying 
special alignment			
Tue Apr 24 16:24:04 2012	eileen	4/24	1:00 PM	4:20 PM	fishione	rttomo	Odde LLCPK1 scans and tilt series	1 2x2		
Tue Apr 24 21:35:18 2012	DNM	4/24	5:30 PM	9:30 PM	70	serviceman	Special alignments - success!			
Wed Apr 25 13:03:44 2012	Maria Pagratis	04/25	8:00 AM	1:00 PM	70	training	Lense capsule	1		
Wed Apr 25 16:33:47 2012	eileen	4/25	1:30 PM	4:30 PM	fishione	rttomo	Odde Mts	2 2x2		
Thu Apr 26 17:01:51 2012	Joey	04/26	9:15 AM	4:55 PM	f30cryo	cryotomo	WT cytos	5		GIF would not stay aligned, at one 
point I had to realign the ZLP 
about every two minutes. Also, a 
lot of problems taking tilt series. 
I had to stop two as it could no 
longer track and would move so far 
off imaging area that I could not 
realign it and continue. This 
occurred suddenly within a single 
loop. Not sure what was causing 
this. Also seemed to have a lot of 
trouble keeping the beam centered 
by the end of the session. 
Mon Apr 30 22:29:18 2012	DNM	4/30	10:00 AM	10:30 PM	70	development	Recalibrate for special alignment, 
debug realign to item			
Tue May  1 18:30:48 2012	eileen	5/1	8:45 AM	6:30 PM	tiltrotate	rttomo	Tomoko/Ghoshima augmin RNAi	3 3x3 montage		
Wed May  2 14:49:30 2012	Cindi	05/02	7:45 AM	2:45 PM	f30cryo	cryotomo	vitreous sxns giardia wt	2		view and preview aren't lined up.
Wed May  2 18:31:37 2012	DNM	5/2	4:15 PM	6:30 PM	70	development	realign to item maybe last bug			
Thu May  3 16:54:57 2012	Joey	05/03	9.00 AM	4.50 PM	f30cryo	cryotomo	WT cytos	6	No problems with GIF today	stage position shown on screen with 
crosshairs was a bit off from 
actual stage positions
Fri May  4 12:54:50 2012	eileen	5/4	8:30 AM	1:00 PM	tiltrotate	rttomo	Tomoko/Ghoshima augmin RNAi	1 3x3 montage and grid scans		
Mon May  7 15:01:32 2012	eileen	5/7	8:45 AM	3:00 PM	tiltrotate	rttomo	Tomoko/Ghoshima augmin RNAi	2 3x3 montage		
Tue May  8 15:29:53 2012	Maria Pagratis	05/08	9 AM	3:30 PM	tiltrotate	rttomo	Dr. Chris Goulbourne	4		
Mon May 14 18:02:32 2012	Kirmse	05/14	9:30 AM	6:00 PM	f20cryo	cryoimaging	IF	0		
Tue May 15 17:52:53 2012	eileen	5/15	9:00 AM	6:00 PM	fishione	rttomo	Candida wt and scans	6		
Wed May 16 16:08:22 2012	eileen	5/16	8:45 AM	4:00 PM	tiltrotate	rttomo	Tomoko/Ghoshima augmin RNAi	2 3x3 montage		
Thu May 17 15:22:41 2012	Kirmse	05/16	9:30 AM	3:30 PM	f30cryo	cryoimaging	IF	0		
Fri May 18 16:06:52 2012	cindi	05/18	9:00 AM	4:00 PM	f30cryo	cryotomo	delta giardin-gfp cytos	5		used special alignments with 30um 
aperture. It was awesome! also 
increased image shift limit to 2 
um. BUT stage needs some 
adjustment. It's very wonky during 
tilting
Mon May 21 15:20:37 2012	Maria Pagratis	5/21/12	9 AM	3 PM	tiltrotate	rttomo	liposomes	5		
Tue May 22 16:44:32 2012	eileen	5/22	8:45 AM	4:40 PM	tiltrotate	rttomo	RPE1 supermontage	2 3x3 montage	trouble with realign to item	
Wed May 23 13:17:40 2012	morphew	5/23	8 AM	1 PM	tiltrotate	rttomo	chromatin	4		
Thu May 24 13:09:23 2012	Spilman	5/24	8:00 AM	1:30 PM	f30cryo	serviceman	Test	1		Compustage alignments and 
adjustments for use with heavy 
holders. 
Tue May 29 18:28:56 2012	DNM	5/29	9:20 AM	6:30 PM	tiltrotate	development	Realign to item bugs with Eileen's 
sample			
Wed May 30 15:44:40 2012	pagratis	5/31	9 AM	3:30 PM	tiltrotate	rttomo	goulbourne			
Thu May 31 15:44:38 2012	Cindi	05/31	8:30 AM	4:00 PM	f30cryo	cryotomo	Giardia delta-gfp cytos	3		lots of issues tilting. probably 
need a floating rod. new alignments 
and SEM calibrations great! 
Fri Jun  1 13:23:04 2012	DNM	6/1	10:30 AM	1:30 PM	70	development	Non-centered trimming in realign to 
item			
Tue Jun  5 13:38:11 2012	eileen	5/4	8:45 AM	12:00 AM	tiltrotate	rttomo	supermontage with overnight	4 2x3		
Tue Jun  5 13:39:42 2012	eileen	6/5	12:00 AM	1:30 PM	tiltrotate	rttomo	supermontage overnight acquisition	2 3x2		
Tue Jun  5 16:16:45 2012	DNM	6/5	2:00 PM	4:15 PM	tiltrotate	development	Nailed weird realign failure on 
second tilt series			
Tue Jun  5 17:26:21 2012	eileen	6/5	4:15 PM	5:30 PM	tiltrotate	rttomo	RPE1 mitotic test series	1		
Wed Jun  6 08:51:50 2012	Cindi	06/06	8:30 AM	8:45 AM	standard	serviceman	move alignment files			moved alignment files and rebooted 
computer
Wed Jun  6 13:51:12 2012	Spilman	6/6/	11:00 AM	4:00 PM	standard	serviceman	none	9		Replace 5 volt supply
Wed Jun  6 19:22:42 2012	Cindi	06/06	2:00 AM	7:15 PM	tiltrotate	rttomo	Giardia sections	1	Tested plasmon imaging on 350 nm 
sections. 	
Thu Jun  7 15:18:35 2012	DNM	6/7	9:00 AM	3:20 PM	standard	training	going over special alignments with 
Cedric			
Fri Jun  8 15:27:28 2012	eileen	6/8	8:30 AM	4:30 PM	tiltrotate	rttomo	Tomoko/Ghoshima U2OS augmin wt	2 3x3		
Mon Jun 11 15:48:43 2012	eileen	6/10	8:45 AM	3:45 PM	tiltrotate	rttomo	Tomoko/Ghoshima WT U2OS	2 3x3 montage		
Tue Jun 12 16:13:41 2012	pagratis	6/12	9:00 AM	4:00 PM	tiltrotate	rttomo	goulbourne 	2		
Wed Jun 13 13:32:01 2012	Cindi	06/13	7:45 AM	1:30 PM	70	rttomo	Giardia gamma-matrigel	2		
Wed Jun 13 16:11:21 2012	pagratis	6/20	2:30 AM	4:15 PM	tiltrotate	rttomo	Goulbourne	1		
Thu Jun 14 14:27:20 2012	Joey	06/14	9:15 AM	2:22 PM	f30cryo	cryoimaging	None :(	0	Got the cryo rod in second time 
fine, think I have figured out what 
I am doing wrong	
Thu Jun 14 16:53:40 2012	DNM	6/14	3:00 PM	5:00 PM	70	serviceman	getting mag gradient files			
Fri Jun 15 16:59:19 2012	eileen	6/15	8:45 AM	5:00 PM	tiltrotate	rttomo	Tomoko/Ghoshima augmin RNai	2 3x3 montage		
Mon Jun 18 17:20:19 2012	eileen	6/18	8:30 AM	5:15 PM	fishione	rttomo	Ghoshima/Tomoko U2OS augmin RNAi	6 2x3 montages		
Wed Jun 20 17:21:18 2012	eileen	6/20	1:30 PM	5:30 PM	tiltrotate	rttomo	Basrai spc98	4 + 2 1x2 montage		
Thu Jun 21 15:30:32 2012	pagratis	6/21	9 AM	4 PM	tiltrotate	rttomo	goulbourne	4		
Tue Jun 26 17:30:48 2012	eileen	6/22	8:45 AM	6:30 PM	tiltrotate	rttomo	Basrai/Choy SPC98	6 (4 1x2 and 2 1x3)		
Tue Jun 26 17:31:57 2012	eileen	6/26	8:45 AM	5:30 PM	tiltrotate	rttomo	Basrai/Choy WT	9 (6 1x2 and 3 1x3 montage)		
Thu Jun 28 15:35:39 2012	Joey	06/28	9:15 AM	3:23 PM	f20cryo	cryotomo	image WT cytos	3		Some issue with beam not staying 
centred for the trial images. Did 
not cause a serious problem. 
Mon Jul  2 14:21:18 2012	cindi	07/02	8:00 AM	2:15 PM	f20cryo	cryoimaging	giardia cytoskeletons			crashed vacuum 2X. Once while 
putting holder in. Once while 
taking holder out. 
Tue Jul  3 18:35:16 2012	eileen	7/2	9:15 AM	6:30 PM	tiltrotate	rttomo	Celegans metaphase supermontage	2 3x3 montage		
Thu Jul  5 13:24:41 2012	Joey	07/05	9:00 AM	1:18 PM	f20cryo	cryotomo	WT cytos	1		Drift measured during autofocus 
does not correspond to drift 
observed during record images in 
ts, especially bad at high tilt. 
Some instability in the vacuum, 
especially when tilting. 
Fri Jul  6 19:09:18 2012	Cindi	07/06	8:30 AM	7:00 PM	tiltrotate	rttomo	gamma-matrigel	3	3X4 montages	
Mon Jul  9 13:52:29 2012	Spilman	7/9	10:00 AM	2:00 PM	standard	serviceman	service	0		leaky pnuematic solinoid for V3 
replaced
Mon Jul  9 16:50:16 2012	morphew	7/9	2 PM	4:45 PM	tiltrotate	rttomo	Hec1	2		
Tue Jul 10 18:13:07 2012	eileen	7/10	8:45 AM	6:10 PM	tiltrotate	rttomo	C.elegans anaphase supermontage	6 2x3 montages		
Wed Jul 11 15:14:02 2012	pagratis	7/11	9:00 AM	3:00 PM	tiltrotate	rttomo	boulbourne	4		
Fri Jul 13 09:36:03 2012	Spilman	7/12	9:00 AM	4:40 PM	standard	serviceman	test	0		Repair vacuum leak in compustage
Fri Jul 13 19:11:14 2012	Cindi	07/13	9:00 AM	7:00 PM	tiltrotate	rttomo	giardia gamma	3	3 4X3 montages	found a few bugs in serialEM with 
David.
Mon Jul 16 12:23:48 2012	pagratis	7/16	9:00 AM	12:00 PM	tiltrotate	rttomo	goulbourne	3		
Tue Jul 17 18:14:20 2012	eileen	7/17	8:45 AM	6:15 PM	tiltrotate	rttomo	Vogel/Kristy tub4_360A	8		
Wed Jul 18 15:43:44 2012	eileen	7/18	8:45 AM	3:45 PM	tiltrotate	rttomo	Vogel/Kristy WT	4		
Fri Jul 20 09:16:48 2012	cindi	07/19	12:00 PM	12:00 AM	tiltrotate	rttomo	gamma-matrigel	3	4X4 montages	had a com error that required a 
reset of the goniometer
Fri Jul 20 09:17:11 2012	cindi	07/20	12:00 AM	10:00 AM	tiltrotate	rttomo	gamma-matrigel	3	4X4 montages	worked fine!
Fri Jul 20 18:26:28 2012	eileen	7/20	11:30 AM	6:30 PM	tiltrotate	rttomo	Vogel/Krysty wt yeast	6		
Mon Jul 23 18:10:06 2012	eileen	7/23	8:30 AM	6:10 PM	tiltrotate	rttomo	RPE1 anaphase supermontage	dual axis 1x2 supermontage of 3x3 montages		
Wed Jul 25 14:56:08 2012	morphew	7-25	8 AM	3 PM	tiltrotate	rttomo	hec1	4		
Thu Jul 26 15:26:52 2012	cindi/karen davies	07/26	8:00 AM	3:30 PM	tiltrotate	training	training		training of Karen Davies of MPI for 
serialEM	
Fri Jul 27 15:32:17 2012	Cindi/Karen	07/27	8:00 AM	3:30 PM	tiltrotate	training	training		serialEM training	Stage very wonky. Called service. 
Need new float rod, but it's on 
backorder.
Mon Jul 30 14:58:50 2012	Kristin	07/30	8:00 AM	3:00 PM	fishione	rttomo	e. coli	2		
Tue Jul 31 16:54:55 2012	eileen	7/31	11:00 AM	5:00 PM	fishione	rtimage	section survey and montage	0		
Wed Aug  1 17:12:18 2012	eileen	8/1	8:45 AM	5:10 PM	tiltrotate	rttomo	RPE1 metaphase 3x3 montage	2 3x3 montage		The second data disk died and no 
data could be written to e drive. 
David disconnected the second drive 
and data can be written to the 
first drive now.
Thu Aug  2 14:12:08 2012	supervisor	08/02	2:00 PM	2:15 PM	standard	serviceman	Putting scope on standby			Scope will be on standby until the 
morning of Aug 13 when Ron brings 
it back up. Robert will watch the 
IGP pumps to be sure they are 
working.
Mon Aug 13 10:20:01 2012	Spilman	8/13	8:00 AM	10:30 AM	standard	serviceman	none	0		Brought scope up from standby, 
Restarted feg (cold start) Checked 
beam & image.
Tue Aug 14 16:58:15 2012	eileen	8/14	9:00 AM	5:00 PM	tiltrotate	rttomo	RPE1 metaphase 	4		
Fri Aug 17 17:02:55 2012	eileen	8/17	10:30 AM	5:00 PM	tiltrotate	rttomo	RPE1 metaphase scans and montage	1 3x3 montage and grid scans		
Mon Aug 20 15:28:55 2012	eileen	8/20	8:45 AM	3:25 PM	tiltrotate	rttomo	RPE1 metaphase	1 3x3	vac crash during second tilt series 
acquisition. Camera was also at 
room temp. Cindi increased water 
flow and temp back to normal now. 	
Thu Aug 30 13:42:21 2012	Spilman	8/21	9:00 AM	5:00 PM	standard	serviceman	service	0		Install, align compustage floating 
rod. monitor vacuum over multiple 
days 8/21/2012 thru 8/30/2012. 
untill call complete.
Thu Aug 30 13:42:41 2012	Spilman	8/22	9:00 AM	5:00 PM	standard	serviceman	service	0		Install, align compustage floating 
rod. monitor vacuum over multiple 
days 8/21/2012 thru 8/30/2012. 
untill call complete.
Thu Aug 30 13:42:51 2012	Spilman	8/23	9:00 AM	5:00 PM	standard	serviceman	service	0		Install, align compustage floating 
rod. monitor vacuum over multiple 
days 8/21/2012 thru 8/30/2012. 
untill call complete.
Thu Aug 30 13:43:11 2012	Spilman	8/24	9:00 AM	5:00 PM	standard	serviceman	service	0		Install, align compustage floating 
rod. monitor vacuum over multiple 
days 8/21/2012 thru 8/30/2012. 
untill call complete.
Thu Aug 30 13:43:46 2012	Spilman	8/27	1:00 PM	5:00 PM	standard	serviceman	service	0		Install, align compustage floating 
rod. monitor vacuum over multiple 
days 8/21/2012 thru 8/30/2012. 
untill call complete.
Thu Aug 30 13:44:11 2012	Spilman	8/28	8:00 PM	5:00 PM	standard	serviceman	service	0		Install, align compustage floating 
rod. monitor vacuum over multiple 
days 8/21/2012 thru 8/30/2012. 
untill call complete.
Thu Aug 30 13:44:22 2012	Spilman	8/29	8:00 PM	5:00 PM	standard	serviceman	service	0		Install, align compustage floating 
rod. monitor vacuum over multiple 
days 8/21/2012 thru 8/30/2012. 
untill call complete.
Thu Aug 30 13:44:35 2012	Spilman	8/30	8:00 PM	5:00 PM	standard	serviceman	service	0		Install, align compustage floating 
rod. monitor vacuum over multiple 
days 8/21/2012 thru 8/30/2012. 
untill call complete.
Thu Sep  6 13:50:00 2012	spilman	12/5	8:00 AM	6:30 PM	tiltrotate	serviceman	test			Adjusted compustage for normal 
operation.  Found heavy wear on 
cryo holder alowing excessive 
movement when tilting.  Dr 
Masternarde ran multiple maps in 
serial em.  Scope available for use.
Fri Sep  7 15:19:07 2012	eileen	9/7	9:00 AM	3:15 PM	tiltrotate	rttomo	RPE1 metaphase section 7 and baking	2 3x3 montage		
Mon Sep 10 17:28:06 2012	eileen	9/10	8:30 AM	5:30 PM	tiltrotate	rttomo	RPE1 metaphase 	4 3x3 montage		
Tue Sep 11 14:01:03 2012	cindi	09/11	9:00 AM	10:00 AM	70	rttomo	mouse heart	0		crashed vacuum with +/-70 holder. 
can't use in this scope because the 
compustage wasn't calibrated for 
it. Vacuum had to be reset, so 
switched to F20
Wed Sep 12 15:40:29 2012	Cindi/Chris Gilbourne	09/12	9:00 AM	4:00 PM	f30cryo	rttomo	heart wt	5		
Thu Sep 13 12:04:57 2012	Cindi/Chris Gilbourne	09/13	8:00 AM	10:30 AM	f30cryo	rttomo	heart wt sections	3		
Fri Sep 14 14:20:56 2012	cindi/minghua	09/14	9:00 AM	2:15 PM	f30cryo	training	training staved pombe			crashed vacuum removing cryo-holder
Mon Sep 17 15:49:19 2012	Cindi	09/17	1:40 PM	3:40 PM	standard	rtimage	giardia sections			
Tue Sep 18 14:43:09 2012	eileen	9/18	8:45 AM	2:30 PM	tiltrotate	rttomo	RPE1 metaphase	2 3x3 montage		
Thu Sep 20 14:18:56 2012	Cindi	09/19	9:00 AM	2:30 PM	tiltrotate	training	Making training videos	1		
Thu Sep 20 14:19:10 2012	Cindi	09/20	9:00 AM	2:15 PM	tiltrotate	training	Making training videos	4		
Fri Sep 21 15:43:50 2012	Cindi	09/21	8:30 AM	3:45 PM	f30cryo	training	making serialem videos	1	4th grid is the charm!	
Tue Sep 25 08:47:28 2012	Cindi	09/24	8:30 AM	4:00 PM	tiltrotate	training	SerialEM demo vidoes			lost internet/network connectivity
Tue Sep 25 08:47:50 2012	Cindi	09/25	8:30 AM	8:45 AM	tiltrotate	training	SerialEM demo vidoes			cannot connect to bio3d
Wed Sep 26 17:29:01 2012	Janet Meehl Fox/Winey	9/26	1:00 PM	5:10 PM	tiltrotate	rttomo	poc1Del tetrahymena basal bodies	4		When I pushed the eucentric focus 
button during the alignment, it was 
really off.
Fri Oct  5 15:48:23 2012	Cindi	10/05	8:45 AM	3:45 PM	tiltrotate	rttomo	giardia	2	3X3 montages	
Fri Oct  5 17:42:31 2012	DNM	10/5	4:15 PM	5:45 PM	standard	development	testing backlash adjust in Nav			
Mon Oct  8 13:02:17 2012	DNM	10/8	10:45 AM	1:00 PM	standard	development	backlash, realign changes			
Wed Oct 10 15:30:16 2012	Cindi	10/10	9:00 AM	3:30 PM	tiltrotate	rttomo	Giardia	2	2X3 montages	
Thu Oct 11 18:12:37 2012	Matt West	10/11	10 AM	6:15 PM	fishione	rttomo	Cos cell wt	8		
Wed Oct 17 13:45:24 2012	Cindi/Jared Polis	10/17	1:00 PM	2:00 PM	tiltrotate	rttomo	giardia		tour with jared polis	
Tue Oct 23 17:08:33 2012	spilman	10/23	8:00 AM	5:30 PM	standard	serviceman	none	0		PM. Backup harddrive, change PVP 
oil, Replace compress air hoses, 
clean regulators, install 
condensation vent valve. Replace 
water filter.
Wed Oct 31 17:11:50 2012	DNM	10/27	11:00 AM	11:00 PM	standard	development	nav testing		Was continuously logged in until 
10/31, mostly working remotely	
Wed Oct 31 17:13:09 2012	DNM	10/31	5:00 AM	5:00 PM	standard	development	nav testing		Just a wild estimate of how much 
time I was actually on scope since 
10/27	
Tue Nov  6 18:10:20 2012	eileen	11/5	12 PM	6:10 PM	fishione	rttomo	Murray lab/Natalie WT	6		
Thu Nov  8 17:41:15 2012	Matthew West	11/8	9:15 AM	6:00 PM	fishione	rttomo	Cos WT MVBs	12		
Fri Nov  9 10:33:07 2012	Gordana	11/9	9:30 AM	10:30 AM	standard	rtimage	EELS data analysis			
Tue Nov 13 18:20:54 2012	eileen	11/13/12	11:30 AM	6:20 PM	fishione	rttomo	Natalie 276_ndc10 yeast	4		
Wed Nov 14 18:25:29 2012	eileen	11/14	8:15 AM	6:30 PM	fishione	rttomo	Natalie ndc10 and wt	16		
Thu Nov 15 12:26:20 2012	Gordana Dukovic	11/15	9:45 AM	12:25 PM	standard	rtimage	EELS data analysis with 
DigitalMicrograph			
Thu Nov 15 17:04:13 2012	cynthia	11/15	1:00 PM	5:00 PM	tiltrotate	rttomo	human heart chem fixed	2 dual axis		
Fri Nov 16 15:55:12 2012	cynthia	11:16	9:15 AM	5:00 PM	tiltrotate	rttomo	human heart	3.5 dual axis		
Sat Nov 17 15:25:36 2012	DNM	11/17	1:30 PM	3:30 PM	standard	development	checking realign with view offset			
Mon Nov 19 16:15:36 2012	cynthia	11/19	1:30 PM	4:15 PM	tiltrotate	rtimage	human heart			
Tue Nov 20 13:10:24 2012	Cindi	11/20	8:30 AM	1:00 PM	f30cryo	cryotomo	bpv	4	tests for John Heumann	
Wed Nov 21 14:14:26 2012	cynthia	11/21	10:00 AM	2:15 PM	tiltrotate	rttomo	human heart	1		
Mon Nov 26 14:49:28 2012	DNM	11/26/12	10:30 PM	2:50 PM	standard	serviceman	redo ultracam quadrant correction			
Tue Nov 27 11:05:00 2012	Cindi	11/27	9:00 AM	11:00 AM	f30cryo	cryoimaging	BPV		sample not so good	vacuum spiked to 61 when inserting 
holder
Tue Nov 27 11:05:02 2012	Cindi	11/27	9:00 AM	11:00 AM	f30cryo	cryoimaging	BPV		sample not so good	vacuum spiked to 61 when inserting 
holder
Wed Nov 28 16:43:47 2012	eileen	11/28	8:30 AM	4:45 PM	fishione	rttomo	Murray/Natalie yeast 173_1.8	6		
Mon Dec  3 16:41:41 2012	eileen	12/3	9:00 AM	4:40 PM	fishione	rttomo	Murray/Natalie 293_3sk	6		
Wed Dec  5 17:51:17 2012	eileen	12/5	10 AM	5:50 PM	fishione	rttomo	Murray/Natalie ndc10	12		
Thu Dec  6 15:43:53 2012	Cindi	12/06	9:00 AM	3:45 PM	standard	development	trying to figure out pixel sizes			scope was at 10 when logged on
Fri Dec  7 15:39:28 2012	eileen	12/7	9:45 AM	3:40 PM	fishione	rttomo	Murray/Natalie ndc10	9		
Tue Dec 11 17:18:30 2012	eileen	12/11	9:30 AM	5:20 PM	fishione	rttomo	Murray/Natalie ndc10_3sk	11		
Tue Dec 18 15:25:35 2012	Kristin	12/18	8:30 AM	3:30 PM	standard	rttomo	M. barkeri and E. coli	14		
Wed Dec 19 15:35:12 2012	Cindi	12/19	8:30 AM	3:30 PM	tiltrotate	rtimage	crappy salivary gland samples			
Thu Dec 20 17:21:28 2012	Janet Meehl Fox/Winey	12/20	12:15 PM	5:20 PM	tiltrotate	rttomo	Tetrahymena basal bodies	4		Serial EM froze with 3 degrees left 
in a tilt series, David M. checked 
it out.  Data was not lost.
Thu Jan  3 15:36:18 2013	Cindi	01/03	10:00 AM	3:30 PM	tiltrotate	rtimage	giardia	0		
Mon Jan  7 16:19:52 2013	Cindi and Ackerson Lab	0/07	1:00 PM	4:20 PM	f30cryo	cryoimaging	Bacteria			
Wed Jan  9 18:15:26 2013	eileen	1/9	8:45 AM	6:15 PM	fishione	rttomo	Moore 282_8	8		
Tue Jan 15 15:58:49 2013	Cindi	01/15	11:15 AM	4:00 PM	standard	rtimage	bactera +/- selenium		tried some EFTEM and EELS. It is 
NOT clear that there is Se in the 
+Se cells. But there are densities 
not present in control.	
Wed Jan 16 15:43:20 2013	Cindi/Cynthia	01/16	2:45 PM	3:45 PM	standard	rtimage	salivary glands of fly			
Fri Jan 18 18:11:10 2013	eileen	1/18	8:45 AM	6:10 PM	tiltrotate	rttomo	stain tests and Moore 282_2	9		
Mon Jan 21 15:58:26 2013	DNM	1/21	3:40 PM	4:00 PM	standard	development	Looking at low-dose pictures before 
getting data elsewhere			
Tue Jan 22 17:59:36 2013	eileen	1/22	9:00 AM	6:00 PM	tiltrotate	rttomo	Murray/Natalie 293_3sk	6 2x1 montages		
Wed Jan 23 17:00:47 2013	eileen	1/23	9:00 AM	5:00 PM	tiltrotate	rttomo	testing unstained sections/ SF9	5 ts plus focus series	crashed vacuum- hair on o-ring! vac 
recovered fine.	
Fri Jan 25 15:18:26 2013	Joey	01/25	9:30 AM	3:15  PM	f30cryo	cryoimaging	Tried to find microvesicles	0	Crashed vacuum - ooopsee! It 
recovered very quickly though.	Problems with objective apature. I 
used the first 30 um apature and 
found that it was constantly 
drifting off centre. I had to 
recentre it about 20 times. 
Mon Jan 28 16:43:54 2013	eileen	1/28	9:00 AM	4:45 PM	tiltrotate	rttomo	Patrick Willey/ Candida	3		
Tue Jan 29 16:58:44 2013	eileen	1/29	9:00 AM	4:50 PM	tiltrotate	rttomo	Patrick Willey Candida	6		
Fri Feb  1 18:13:16 2013	eileen	2/1	9:00 AM	6:15 PM	tiltrotate	rttomo	Murray/Natalie 3sk, Celegans 
supermontage mapping	3		
Mon Feb  4 16:55:07 2013	eileen	2/4	9:00 AM	4:55 PM	tiltrotate	rttomo	Celegans supermontage	1x3 supermontage of 3x3 montage		
Tue Feb  5 19:42:25 2013	eileen	2/5	8:30 AM	7:40 PM	tiltrotate	rttomo	C.elegans anaphase supermontage	6 3x3 montage 		
Wed Feb  6 10:34:16 2013	DNM	2/6	9:20 AM	10;40 AM	standard	development	image pairs from ultracam			
Wed Feb  6 16:47:59 2013	eileen	2/6	10:30 AM	4:50 PM	tiltrotate	rttomo	C.elegans supermontage (1x3 of 3x3)	3 3x3 montages		
Fri Feb  8 17:03:11 2013	eileen	2/8	9:00 AM	5:00 PM	tiltrotate	rttomo	C.elegans anaphase supermontage 
(1x3 of 3x3)	3 3x3 montages		
Mon Feb 11 14:58:09 2013	eileen	2/11	8:45 AM	3:00 PM	tiltrotate	rttomo	C.elegans supermontage T0265_31 
(1x3 super of 3x3 montages)	3 3x3 montage		
Tue Feb 12 20:20:48 2013	Matt West	2/12	10:30 AM	8:15 PM	tiltrotate	rttomo	Cos mutants	8	Thanks to Eileen I was able to 
salvage a scope day after the f-20 
camera failed.	
Wed Feb 13 09:59:21 2013	Cindi	02/13	8:45 AM	10:00 AM	standard	rtimage	epon FS samples		looked like crap :(	
Wed Feb 13 16:38:54 2013	eileen	2/12	2:30 PM	4:40 PM	tiltrotate	rtimage	Scanning Drosophila embryo sections	0		
Tue Feb 19 15:57:16 2013	cindi	02/19	2:00 PM	4:00 PM	standard	rtimage	Giardia on intestines			beam very dim when I logged in, 
came up nice with alignments
Wed Feb 20 18:08:19 2013	Cindi	02/20	8:30 AM	6:15 PM	standard	rtimage	giardia intestines			
Fri Feb 22 15:57:58 2013	Minghua Liu	02/22	9:10 AM	4:00 PM	f30cryo	cryoimaging	Pombe			
Mon Feb 25 17:04:29 2013	eileen	2/25	8:25 AM	5:00 PM	tiltrotate	rttomo	mtth yeast	7		camera error
Tue Feb 26 16:53:52 2013	eileen	2/26	8:45 AM	5:00 PM	tiltrotate	rttomo	MTTH yeast Murray/Natalie 3sk yeast	8		
Wed Feb 27 17:08:31 2013	eileen	2/27	11:00 AM	5:10 PM	tiltrotate	rttomo	murray/natalie ndc10 yeast	4		vac went up to 20 when tilting (-
20). cleaned fibers off o-ring and 
put new grease. vac was fine the 
rest of the day.
Thu Feb 28 11:27:23 2013	Kristin	2/28	8:30 AM	11:30 AM	standard	rttomo	didn't get past alignments	0		Digital Micrograph not recognizing 
camera (detailed e-mail sent to 
Cindi and Eileen)
Fri Mar  1 15:20:00 2013	Cindi	03/01	9:00 AM	3:20 PM	f30cryo	cryoimaging	BPV, sample too thick			DM took 16 times to work. Crashed 
vacuum inserting cryo holder the 
first time.
Mon Mar  4 14:45:17 2013	DNM	3/4	11:00 AM	2:45 PM	standard	serviceman	Trying to get ultracam to work, 
pairs on upper camera			
Tue Mar  5 16:32:21 2013	DNM	3/5	10:0 AM	4:30 PM	standard	development	testing dose rates and counts			
Fri Mar  8 18:30:48 2013	eileen	3/8	9:00 AM	6:30 PM	tiltrotate	rttomo	Chlamy uni1	6		
Mon Mar 11 17:14:54 2013	eileen	3/11	1:00 PM	5:15 PM	tiltrotate	rttomo	Chlamy uni1	8		
Tue Mar 12 15:26:18 2013	eileen	3/12	9:00 AM	3:30 PM	tiltrotate	rttomo	Chlamy uni1	5		vac issues (23 at tilt 20)with both 
holders. Cleaned o-rings, and it 
resolved. no vac issues at all 
during tilt series acquisition and 
none at all with eucentricity.
Wed Mar 13 18:31:39 2013	Kristin	3/13	8:30 AM	6:30 PM	tiltrotate	rttomo	E. coli	2		
Thu Mar 14 15:37:15 2013	Spilman	3/14	8:00 AM	3:45 PM	standard	serviceman	stage	9		Lube stage orings to eliminate leak 
during tilt. align stage. adjust 
switch for sensing the holders.
Thu Mar 14 15:37:33 2013	Spilman	3/14	8:00 AM	3:45 PM	standard	serviceman	stage	9		Lube stage orings to eliminate leak 
during tilt. align stage. adjust 
switch for sensing the holders.
Tue Mar 19 18:10:39 2013	eileen	3/9	8:45 AM	6:10 PM	70	rttomo	pombe and Drosophila	6		the microscope recognized the 
holder on the first insertion, but 
then did not unless the turbo was 
off and Cindi used her magic touch!
Wed Mar 20 16:48:31 2013	eileen	3/20	9:30 AM	4:50 PM	tiltrotate	rtimage	pombe and Drosophilia supermontage 
mapping	2		
Thu Mar 21 15:25:23 2013	Cindi	03/21	8:30 AM	3:30 PM	f30cryo	cryotomo	BPV	2		
Fri Mar 22 17:12:49 2013	eileen	3/22	2:30 PM	5:10 PM	standard	rtimage	Drosophila scans	0		
Mon Mar 25 16:52:34 2013	eileen	3/25	8:45 AM	4:50 PM	tiltrotate	rttomo	Candida 102	6		
Mon Apr  1 13:51:51 2013	Cindi	04/01	8:30 AM	2:00 PM	tiltrotate	rttomo	Giardia	2		alpha wobbler stopped when pressing 
z-axis button. Reset goniometer, 
worked fine after.
Tue Apr  2 14:45:50 2013	Cindi	04/02	9:00 AM	2:45 PM	tiltrotate	rttomo	giardia for fib	4		
Wed Apr  3 12:46:02 2013	cynthia	04/02/13	9:00 AM	12:45 PM	f30cryo	cryotomo	bpv		got one grid in then we had two 
crashes trying to get the second 
grid in	got one grid in then we had two 
crashes trying to get the second 
grid in. We are baking out the F30 
rod as there was a alot od frost on 
the rod itself and the dewar was 
cold.
Fri Apr  5 17:34:28 2013	eileen	4/5	9:00 AM	5:30 PM	fishione	rttomo	pombe 2738	3 1x2		
Mon Apr  8 10:48:14 2013	Cindi	04/08	10:30 AM	10:45 AM	standard	rtimage	MTs		not polymerized	
Thu Apr 11 16:40:03 2013	Cindi	04:11	8:30 AM	4:30 PM	f20cryo	cryoimaging	BPV	0	The f30 rod crashed, Cindi will 
replace O-ring. The F20 rod worked 
OK even after the crash!	
Fri Apr 19 16:36:23 2013	Cindi	04/19	8:30 AM	4:30 PM	tiltrotate	rttomo	neurons for FIB	6		
Mon Apr 22 15:40:16 2013	Cindi	04/22	8:30 AM	2:30 PM	f30cryo	cryoimaging	BPV		sample concentration way too low	
Mon Apr 22 15:40:41 2013	Cindi	04/22	2:30 PM	3:45 PM	tiltrotate	rttomo	giardia, testing auto acquire TS	1		
Fri Apr 26 15:47:58 2013	Joey	04/26	9:30 AM	3:45 PM	f30cryo	cryotomo	Mth tagged cytos	1		
Tue Apr 30 11:51:53 2013	Cindi	04/30	9:30 AM	12:00 PM	f30cryo	cryoimaging	giardia with 17053-mth		crappy sample	
Thu May  2 14:49:15 2013	Joanna	05/02	8:15 AM	2:45 PM	f30cryo	cryoimaging	none	0		Had problems with the vacuum. 
Crashed with F30 rod, then had 
trouble with F20 rod. Airlock not 
counting down, instead constant 
pumping and could not get rod in. 
Later, (3 hours) Mary was about to 
insert warm cryorod twice and 
everything seemed normal. 
Fri May  3 11:30:49 2013	Cindi	05/03	9:30 AM	11:30 AM	standard	serviceman	testing vacuum when inserting 
holders			Tried 3X with F30 cryo holder, 2X 
with T/R holder, 3X with F20 cryo 
holder. Took some time to come down 
after inserting holders, but it 
always worked. Also, changed o-ring 
on F20 cryo holder.
Mon May  6 15:39:10 2013	DNM	5/6/13	11:00 AM	3:40 PM	70	development	Realign to item with huge view 
defocus offset			
Fri May 10 16:51:54 2013	cynthia	05/10/13	2:00 PM	5:00 PM	standard	rtimage	k4m sections of DES_MTH		When I fisrt started the scope 
colum vacuum was 30. Removed rod 
and cleaned of O ring there were 
about 8 threads/hairs! And once the 
rod was in the vacuum was back down 
again.	
Fri May 17 13:13:20 2013	DNM	5/17/13	11 AM	1:15 PM	tiltrotate	development	testing stability of rotating rod 
with cable tethered to stage	1		Rotation stopped working - seems to 
be the cable
Tue May 21 12:30:52 2013	eileen	5/22	9:15 AM	12:30 PM	fishione	rttomo	candida 102	1		
Wed May 22 16:12:33 2013	eileen	5/22	8:45 AM	4:15 PM	fishione	rttomo	Candida 102 anaphase	4		
Tue May 28 17:31:08 2013	eileen	5/28	1:00 PM	5:30 PM	fishione	rttomo	Odde Candida	4		
Wed May 29 16:56:46 2013	eileen	5/29	8:45 AM	5:00 PM	fishione	rttomo	Odde Candida kip1del	5		
Fri May 31 15:23:47 2013	eileen	5/31	9:00 AM	3:30 PM	fishione	rttomo	Odde Candida 12711	5		
Mon Jun  3 16:33:26 2013	morphew	6/3	9 AM	4:30 PM	f30cryo	cryotomo	BPV_focus series	4		
Tue Jun  4 17:59:41 2013	eileen	6/4	8:45 AM	6:00 PM	fishione	rttomo	Candida strain 102	7		The stage was not always moving to 
the right place when dragging in 
SerialEM.
Wed Jun  5 16:55:00 2013	eileen	6/5	8:45 AM	5:00 PM	fishione	rttomo	Candida strains 102 and 12711	10		
Thu Jun  6 15:36:27 2013	morphew	6/6	1 PM	3:30 PM	fishione	rtimage	tet-MTH			
Fri Jun  7 13:42:27 2013	eileen	5/7	10 AM	1:30 PM	fishione	rttomo	Candida 12711	3		
Thu Jun 13 17:17:13 2013	Joey	06/13	9:30 AM	5:15 PM	tiltrotate	rttomo	Montage of cyto	17		
Mon Jun 17 17:27:40 2013	eileen	6/17	12:00 PM	5:30 PM	fishione	rttomo	Candida 111	3		
Tue Jun 18 17:44:18 2013	eileen	6/18	9 AM	5:45 PM	fishione	rttomo	Candida 111 and 12377	5		
Wed Jun 19 17:05:44 2013	eileen	6/19	9:00 AM	5:10 PM	fishione	rttomo	Odde Candida 27377	7		
Fri Jun 21 13:10:41 2013	morphew	6-21	8:45 AM	1 PM	fishione	rttomo	tetrehymena MTH	2		
Mon Jun 24 17:06:41 2013	eileen	6/24	9:45 AM	5:00 PM	fishione	rttomo	Tetrahymena K4M 3MTH	4		
Tue Jun 25 18:25:04 2013	eileen	6/25	9:00 AM	6:30 PM	fishione	rttomo	SPC42_MTH freeze sub resins	6		
Wed Jun 26 15:45:58 2013	morphew	6-26	10 AM	3:45 PM	fishione	rtimage	mth			
Thu Jun 27 16:16:04 2013	c page	6/27	10 AM	4:15 PM	tiltrotate	rtimage	prasad rotovirus	0		
Fri Jun 28 17:48:16 2013	cynthia	06/28	10:00 AM	5:30 PM	fishione	rttomo	Prasad rotavirus epon	6		
Tue Jul  2 15:42:13 2013	cynthia	07/01	10 AM	12:00 PM	fishione	cryotomo	prasad	0		Did not recognize the fishione the 
second time I tried to put it in.
Tue Jul  2 15:44:51 2013	cynthia	07/01	12:00 PM	4:00 PM	tiltrotate	cryotomo	prasad plastic 	1	didn't work because I was havinfg 
trouble with my account need 
alighment tab and files to load 
lost of problems with aperature no 
way to load new alignments	need and updated account
Tue Jul  2 15:46:25 2013	cynthia	07/02	10:00 AM	3:00 PM	tiltrotate	rttomo	prasad plastic	4 , 2X1	I used Eileens account	
Wed Jul  3 17:02:04 2013	eileen	7/3	11:30 AM	5:00 PM	tiltrotate	rttomo	Natalie/Murray 273_2	6		
Mon Jul  8 14:03:21 2013	Cindi	07/09	1:30 PM	2:00 PM	standard	rtimage	Brain Sections			
Mon Jul  8 16:20:40 2013	morphew	7/8	2:30 PM	4:30 PM	tiltrotate	rttomo	desmin_MTH	1		
Tue Jul  9 12:47:30 2013	eileen	7/9	11:30 AM	12:45 PM	tiltrotate	rtimage	yeast examples	0		
Tue Jul  9 18:06:49 2013	morphew	7/9	1:00 PM	4:15 PM	tiltrotate	rttomo	desmin MTH	2		
Tue Jul  9 18:07:55 2013	eileen	7/9	4:15 PM	6:00 PM	tiltrotate	rttomo	SPC42_MTH_FS_HM20_green	2		
Wed Jul 10 18:18:30 2013	Joey	07/10	11:05 AM	6:15 PM	tiltrotate	rttomo	neg stain cyto montages	16		
Thu Jul 11 10:06:59 2013	Spilman	7/11	8:15 AM	10:15 PM	fishione	serviceman	test	0		Adjust specimen airlock switches.
Thu Jul 11 18:13:31 2013	Cindi	07/11	2:30 PM	6:15 PM	70	rttomo	MTH samples	4		
Fri Jul 12 16:45:34 2013	Joey	07/11	9:30 AM	4:45 PM	tiltrotate	rttomo	Neg stained cytos	16		
Tue Jul 16 17:34:45 2013	eileen	7/16	8:45 AM	5:30 PM	tiltrotate	rttomo	SPC42_MTH samples	6		
Wed Jul 17 16:54:04 2013	eileen	7/17	2:00 PM	5:10 PM	70	rtimage	spc42MTH scans	0		
Thu Jul 18 09:12:32 2013	Cindi	07/18	8:30 AM	9:15 AM	standard	rtimage	samples for FIB			
Thu Jul 18 14:36:18 2013	cynthia	07/19	1:30 PM	2:30 PM	tiltrotate	rttomo	prasad plastic		grid broke!	
Fri Jul 19 17:47:21 2013	eileen	7/19	8:45 AM	4:00 PM	70	rttomo	SPC42-MTH	2	The microscope no longer recognizes 
the +/- 70 holder.	
Mon Jul 22 13:29:21 2013	cynthia	07/22	10:30 AM	1:30 PM	tiltrotate	rttomo	orasad plastic			
Mon Jul 22 17:57:17 2013	eileen	7/22/13	2:30 PM	6:00 PM	tiltrotate	rttomo	spc42_mth	1		
Tue Jul 23 11:39:04 2013	Cindi	07/23	9:00 AM	11:40 AM	standard	rtimage	PTA stained samples			rebooted computer to boost intensity
Tue Jul 23 17:55:27 2013	Cynthia	07/23	12:30 PM	6:00 PM	tiltrotate	rttomo	prasad plastic 	2		
Fri Jul 26 10:47:22 2013	Spilman	7/23	8 AM	5:00 PM	standard	serviceman	Test	0		PM due,  Made electronic & vaccum checks,  
replaced pump oil,  water filter, Xray checks, 
inspected hoist. re[aced 50 volt supply for HT.
Fri Jul 26 10:48:10 2013	Spilman	7/24	8:00 AM	5:00 PM	standard	serviceman	Test	0		PM due,  Made electronic & vaccum checks,  
replaced pump oil,  water filter, Xray checks, 
inspected hoist. re[aced 50 volt supply for HT.
Fri Jul 26 10:48:51 2013	Spilman	7/26	8:00 AM	11:00 AM	standard	serviceman	Test	0		PM due,  Made electronic & vaccum checks,  
replaced pump oil,  water filter, Xray checks, 
inspected hoist. re[aced 50 volt supply for HT.
Fri Jul 26 14:16:50 2013	morphew	7/26	11:00 AM	2:30 PM	f30cryo	cryotomo	BPV			
Fri Jul 26 16:00:58 2013	Cindi	07/26	2:30 PM	4:00 PM	standard	serviceman	FEG Crash			Mary had 2 vacuum crashes. Probably 
coincidental. HT died at 268. Ron 
coming Monday to look at it again.
Mon Jul 29 17:27:21 2013	Byung-Ho	07	11:30 AM	5:30 PM	tiltrotate	rttomo	C elegans	10		
Tue Jul 30 16:44:42 2013	Byung-Ho	07/30	8:30 AM	4:40 PM	fishione	rttomo	Arabidopsis leaf	16		
Wed Jul 31 14:25:11 2013	Byung-Ho	07/31	8:40 AM	2:20 PM	fishione	rttomo	Arabidopsis	8		
Wed Jul 31 17:09:44 2013	cindi	07/31	2:45 PM	5:15 PM	fishione	rttomo	PTA tests	3		
Thu Aug  1 16:21:47 2013	morphew	8/1	10 AM	4:30 PM	f20cryo	cryotomo	bpv_focal pairs			
Mon Aug  5 14:45:15 2013	Cindi	08/05	9:00 AM	2:45 PM	fishione	rttomo	FIB samples	5		Fischione rod makes the program ask 
you multiple times for what holder 
you need.
Tue Aug  6 18:58:37 2013	Cynthia	08/06	10:30 AM	7:00 PM	tiltrotate	rttomo	prasad plastic sucrose	4 duals	took a while 20 minutes for vacuum 
to go below 20 with t/r, standard 
went below right away when I cryo 
cycled	
Wed Aug  7 17:35:38 2013	Cynthia	08/07	10:15 AM	5:30 PM	tiltrotate	rttomo	prasad sucrose plastic	6		
Thu Aug  8 17:30:17 2013	Cynthia	08/08	11:00 AM	5:30 PM	tiltrotate	rttomo	prasad sucrose plastic	5	vacuum went down quickly when I put 
in T/R rod	
Fri Aug  9 20:35:02 2013	Joey	08/09	10:50 AM	8:30 PM	tiltrotate	rttomo	Whole cytos	32		
Mon Aug 12 12:36:49 2013	Cindi	08/12	9:00 AM	12:40 PM	tiltrotate	rttomo	FIB tests	1		
Mon Aug 12 18:02:19 2013	eileen and Jayne (Moore lab)	8/12	1:00 PM	6:00 PM	tiltrotate	rttomo	282 yeast spindles	6		
Tue Aug 13 18:41:28 2013	eileen	8/13	9:00 AM	6:40 PM	tiltrotate	rttomo	Candida 102	8		
Wed Aug 14 10:38:08 2013	eileen	8/14	8:45 AM	10:30 AM	tiltrotate	rttomo	Candida 102	2		
Fri Aug 16 17:04:35 2013	Cindi	08/16	9:00 AM	5:00 PM	tiltrotate	rttomo	FIB Samples	4		noticed vacuum spikes were less and less each 
time I inserted rod.
Mon Aug 19 16:55:07 2013	Joey	08/19	10:15 AM	5:00 PM	tiltrotate	rttomo	whole discs	16		Vacuum took a long time to come 
down with tilt rotate rod. Removed 
rod, added vacuum grease and that 
helped a little. Also started 
imaging with iptg 1 = 17. Then 
seemed to come down after that. 
Tue Aug 20 18:13:41 2013	eileen	8/20	8:45 AM	6:10 PM	tiltrotate	rttomo	Desmin-MTH, WTyeast_Au(III)CL, 
Candida 102	7		
Wed Aug 21 16:24:52 2013	eileen	8/21	8:45 AM	4:30 PM	tiltrotate	rttomo	WT-desmin secondary only, Candida 
102	5		
Fri Aug 23 16:05:30 2013	Joey	08/22	10:15 AM	4:00 PM	tiltrotate	rttomo	Montage whole discs	16		The last montaged image of the tilt series 
contained two images that were completely 
black. All other images were fine. I do not this 
this was due to a grid bar because they were 
right in the center of the montage and the rest 
of the other images were fine. 
Fri Aug 23 17:00:17 2013	cindi	08/23	9:30 AM	4:30 PM	tiltrotate	rttomo	fib samples	8		
Mon Aug 26 17:32:56 2013	eileen	8/26	11 AM	5:30 PM	tiltrotate	rttomo	Moore 282	8		
Tue Aug 27 18:20:59 2013	eileen	8/27	9:00 AM	6:20 PM	tiltrotate	rttomo	Moore lab 282  alpha SPC42MTH	5		
Fri Aug 30 17:02:55 2013	Joey	08/30	10:00 AM	5:00 PM	tiltrotate	rtimage	checking out grids	0		
Wed Sep  4 20:43:24 2013	Joey	09/04	9:30 AM	8:45 PM	tiltrotate	rttomo	montage whole discs	32		
Fri Sep  6 15:00:36 2013	morphew	9-6	9 AM	3 PM	f20cryo	cryotomo	bpv	0		3 vac crashes--bailed
Mon Sep  9 16:17:59 2013	morphew	9/9	12:30 PM	4:30 PM	f30cryo	cryotomo	bpv focal pairs	6		
Wed Sep 11 16:12:49 2013	Cynthia	09/11	10:00 AM	4:00 PM	tiltrotate	rtimage	dedecker		take the cryo box out dingbat :)	
Mon Sep 16 15:08:14 2013	morphew	9/16	1 PM	3 PM	f30cryo	cryotomo	bpv	0		vac crash, slow recovery
Fri Sep 20 08:48:48 2013	Ben DE-12 Install	09/17	10:00 AM	7:00 PM	standard	serviceman	try and install DE-12 on scope			
Fri Sep 20 08:49:09 2013	Ben DE-12 Install	09/18	10:00 AM	7:00 PM	standard	serviceman	remove DE-12 and put on F20			
Fri Sep 20 08:49:59 2013	Ron Spilman	09/18	8:00 AM	5:00 PM	standard	serviceman	Reinstall gatan camera			vacuum leak. reinstall gatan 
camera. looking better friday 
morning
Fri Sep 20 17:17:59 2013	cindi	9/20	8:30 AM	5:15 PM	tiltrotate	rttomo	Selenium bacteria	4		
Mon Sep 23 16:50:25 2013	eileen	9/23	9:30 AM	5:00 PM	tiltrotate	rttomo	SPC42_MTH and wt preps	4		
Tue Sep 24 14:38:37 2013	Minghua Liu	9/24	10:00 AM	2:30 PM	f30cryo	cryoimaging	Pombe			
Wed Sep 25 18:14:04 2013	eileen	9/25	8:30 AM	6:10 PM	tiltrotate	rttomo	Odde Candida	5		
Wed Sep 25 19:04:00 2013	Minghua	09/25	6:00 PM	7:00 PM	tiltrotate	rtimage	pombe			
Fri Sep 27 15:50:27 2013	eileen	9/27	8:45 AM	3:34 PM	tiltrotate	rttomo	nsp1_MTH	6		
Mon Sep 30 18:01:02 2013	Cynthia	09/30	10:30 AM	6:00 PM	f20cryo	cryotomo	desmin wt	7	Nice little F30, Thank you!	
Tue Oct  1 17:00:11 2013	eileen	10/1	8:30 AM	5:00 PM	tiltrotate	rttomo	eg5_mts_A3 and wt	6		
Wed Oct  2 15:05:59 2013	Cynthia	10/02	9:30 AM	11:30 AM	f20cryo	cryotomo	des mth		lousy grids, tooooo thick	
Fri Oct  4 13:32:36 2013	eileen	10/4	9:00 AM	1:30 PM	tiltrotate	rttomo	SPC42MTH nano only	2		
Fri Oct  4 15:12:23 2013	morphew	10/4	1:30 PM	3 PM	f30cryo	cryoimaging	vitreous sections			vac crash
Sat Oct  5 22:10:29 2013	Byungho Kang	10/05	2:00 PM	10:00 PM	tiltrotate	rttomo	C elegans Arabidopsis	20		
Sun Oct  6 18:41:00 2013	Byung-Ho Kang	10/06	10:00 AM	6:30 PM	tiltrotate	rttomo	C elegans, Arabidopsis	22	Thanks. F30 worked well.	
Tue Oct  8 18:39:54 2013	eileen	10/8	9 AM	6:30 PM	tiltrotate	rttomo	nic96-2mth, wt, Milan_cmk yeast	7		
Wed Oct  9 17:38:21 2013	eileen	10/9	8:45 AM	5:45 PM	tiltrotate	rttomo	milan yeast, nic96MTH, chlamy	12		
Thu Oct 10 16:38:18 2013	Cindi	10/10	4:15 PM	4:30 PM	standard	rtimage	neg stain MTs			
Fri Oct 11 19:02:34 2013	Cynthia	10/11	9:30 AM	7:00 PM	tiltrotate	rttomo	desmin wt and mth tail 9/24 HPF2	7		
Mon Oct 14 17:44:54 2013	eileen	10/14	8:45 AM	5:45 PM	tiltrotate	rttomo	Milan yeast and nup1008MTH	6		
Mon Oct 14 22:26:45 2013	DNM	10/13	11:00 PM	12:15 AM	standard	development	Testing remote access from 64-bit 
version			
Tue Oct 15 16:49:25 2013	minghua	10/15	11:00 AM	5:00 PM	f30cryo	cryoimaging	pombe			
Wed Oct 16 19:40:55 2013	eileen	10/15	9:00 AM	7:45 PM	tiltrotate	rttomo	nup100 8MTH chlamy	10		
Thu Oct 17 12:06:47 2013	DNM	10/18	11:00 AM	12:10 PM	standard	development	testing remote server nd 64-bit 
version	1		
Thu Oct 17 16:04:17 2013	Cynthia	10/17	1:30 PM	4:00 PM	tiltrotate	rtimage	Prasad Time course t2			screen meter not registering beam
Mon Oct 21 18:54:07 2013	eileen	10/21	8:30 AM	7:00 PM	tiltrotate	rttomo	nup100_8MTH and Milan yeast dmso	10		
Tue Oct 22 14:55:08 2013	Cindi	10/22	8:30 AM	2:45 PM	tiltrotate	rttomo	71+selenium bacteria	8		
Tue Oct 22 18:11:38 2013	eileen	10/23	3:00 PM	6:10 PM	tiltrotate	rttomo	wt control nup100	4		
Mon Oct 28 16:07:19 2013	Spilman	10/28	9:15 AM	4:30 PM	standard	serviceman	test	0		Communication issue cause by faulty 
Right handpanel.
Tue Oct 29 17:54:51 2013	DNM	10/29	11:00 AM	5:00 PM	standard	serviceman	reinstalling Gatan camera and 
recovering vacuum			
Tue Oct 29 17:55:35 2013	DNM	10/29	5:00 PM	6:00 PM	standard	development	checking if focus change between 
GIF and uooer camera changes 
intensity			
Wed Oct 30 17:46:58 2013	Cindi	10/30	8:30 AM	5:45 PM	70	rttomo	spc42mth and se bacteria	4		
Thu Oct 31 17:11:42 2013	Cynthia	10/31	12:00 PM	5:00 PM	tiltrotate	cryotomo	prasad	3		cryo box was out. I could not get 
comeplete tilt series would stop 
after about 112 images at about 48 
degrees and claim to be more than 
30% out of alignment, yet the 
record in 2 of 3 cases was in the 
correct area.kept taking pictures 
at the same spot doing "reset shift 
and realign" and I got poletouch 
messages from Microscope Control 
program
Fri Nov  1 17:37:39 2013	Cynthia	11/01	10:30 AM	5:30 PM	tiltrotate	rttomo	prasad time course time 2	2x1, 2X2	I could only go -54 to 60 because I 
was getting pole touches at tilts 
above -54	
Mon Nov  4 18:11:17 2013	c page	11/04	3:00 PM	6:00 PM	tiltrotate	rttomo	a3desmin hm20	3	worked great, +_60!	
Wed Nov  6 17:25:20 2013	Cynthia	11/06	10:30 AM	5:30 PM	tiltrotate	rttomo	prasad timepoint3	1		
Thu Nov  7 17:19:57 2013	Cynthia	11/07	10:15 AM	5:15 PM	tiltrotate	rttomo	prasad timepoint 2	7		
Tue Nov 12 17:32:17 2013	eileen	11/12	9 AM	5:30 PM	tiltrotate	rttomo	SPC42MTH post stain tests	3		
Wed Nov 13 16:39:10 2013	Cindi	11/13	9:00 AM	4:30 PM	tiltrotate	rttomo	drp1-2xmth-mcherry	5		
Thu Nov 14 15:51:53 2013	eileen	11/14	9:00 AM	2:00 PM	tiltrotate	rttomo	DRB1 MTH gold/HM20	3		
Fri Nov 15 16:44:17 2013	Cindi	11/15	8:30 AM	9:30 AM	standard	rtimage	cos7			
Mon Nov 18 17:43:11 2013	Cynthia	11/18	11:00 AM	5:45 PM	tiltrotate	rttomo	prasad time course2 pt 2	5		
Wed Nov 20 18:12:56 2013	eileen	11/20	9 AM	6:10 PM	tiltrotate	rttomo	Yumei/DeCamilli hpf neurons	12		
Thu Nov 21 15:35:06 2013	Cynthia	11/21	11:00 AM	3:30 PM	tiltrotate	rttomo	prasad tc2 tp3	5		
Fri Nov 22 10:36:07 2013	DNM	11/22	10:00 AM	10:40 AM	standard	development	testing dependence of stage 
calibration on rotation angle			
Fri Nov 22 17:26:42 2013	eileen	11/21	1:15 PM	5:30 PM	tiltrotate	rttomo	Yumei/DeCamilli Dynamin1KO	4		
Mon Nov 25 10:57:50 2013	DNM	11/25	9:55 AM	11:00 AM	standard	development	Testing stage cal with whole image 
correlation			
Tue Nov 26 15:55:39 2013	Cynthia	11/26	11:00 AM	4:00 PM	tiltrotate	rttomo	prasad tc2,4	4	The beam centering is really off 
when you go from low to high mag. 
Is this due to the fact that the C2 
aperture is off center a bit?	
Sun Dec  1 14:43:15 2013	DNM	12/1	8:00 AM	2:45 PM	standard	development	tsting image shift/stage calibration			
Mon Dec  2 17:02:41 2013	eileen	12/2	10:30 AM	5:00 PM	tiltrotate	rttomo	Yumei/De Camilli neurons	3		
Tue Dec  3 15:41:35 2013	eileen	2/3	9:00 AM	3:40 PM	tiltrotate	rttomo	Yumei/DeCamilli neurons Vogel yeast	5		
Wed Dec  4 21:01:57 2013	Johanna	12/4	09:00 AM	9:00 PM	f30cryo	cryotomo	Spermatozoa	11		
Thu Dec  5 20:30:47 2013	Johanna	12/5	9:00 AM	8:30 PM	f30cryo	cryotomo	Human spermatozoa	12		
Sun Dec  8 20:40:48 2013	Johanna	12/7	11:00 AM	4:00 PM	tiltrotate	rtimage	Human spermatozoa	0		
Sun Dec  8 20:43:22 2013	Johanna	12/8	10:30 AM	8:40 PM	f30cryo	cryotomo	Human spermatozoa	9		There was a contamination on the 
uppercamera for a while, but 
disappeared by it self. 
Mon Dec  9 15:51:35 2013	eileen	12/9	8:30 AM	3:50 PM	tiltrotate	rttomo	SPC42-MTH ATM only + imaging	2		
Wed Dec 11 18:08:51 2013	eileen	12/11	4:00 PM	6:00 PM	tiltrotate	rttomo	SPC42-MTH ATM acetate 60min	2		
Fri Dec 13 15:38:01 2013	Cynthia	12/13/13	11:00 AM	3:30 PM	f30cryo	cryotomo	desmin	0	Grid no good, but scope seems fine	
Sat Dec 14 19:31:52 2013	Johanna	14/12	11:00 AM	07:30 PM	f30cryo	cryotomo	Spermatozoa	3		
Sun Dec 15 17:29:52 2013	Johanna	15/12	12:00 AM	5:30 PM	f30cryo	cryotomo	Spermatozoa	3		
Mon Dec 16 17:44:38 2013	eileen	12/16	10:30 PM	5:40 PM	tiltrotate	rttomo	SPC42_MTH pH expts	5		
Tue Dec 17 21:20:53 2013	Johanna	17/12	10:00 AM	09:15 PM	f30cryo	cryotomo	Spermatozoa	5		
Wed Dec 18 21:14:56 2013	Johanna	18/12	09:00 AM	09:15 PM	f30cryo	cryotomo	Spermatozoa	10		
Fri Dec 20 14:50:42 2013	morphew	12/20	11 AM	3 PM	tiltrotate	rttomo	spc42MTH	0		2 vac crashes
Mon Dec 23 16:39:01 2013	Cindi	12/23	9:00 AM	5:00 PM	standard	serviceman	scope down			restart vacuum, no water flow. 
Atzlano came and cleaned out lines 
with compressed air. Got flow, but 
now one of the connectors is 
leaking. Scope is down.
Mon Dec 30 11:06:29 2013	Ron Spilman	12/29	8:00 AM	11:00 AM	standard	serviceman	replace water connectors			replaced connector, now works
Thu Jan  2 16:47:55 2014	Cindi	01/02	9:00 AM	4:45 PM	tiltrotate	rttomo	bacteria +se	10		
Fri Jan  3 17:54:04 2014	eileen	1/3	9:00 AM	5:50 PM	tiltrotate	rttomo	C.elegans anaphase supermontage	3 3x3 montages		
Sat Jan  4 15:03:45 2014	eileen	1/4	8:40 AM	3:00 PM	tiltrotate	rttomo	C. elegans anaphase supermontage	3 3x3 montages		
Mon Jan  6 17:57:22 2014	eileen	1/6	8:20 AM	5:55 PM	tiltrotate	rttomo	C.elegans anaphase supermontage	4 3x3		
Tue Jan  7 18:57:26 2014	eileen	1/7	8:00 AM	7:00 PM	tiltrotate	rttomo	C. elegans anaphase supermontage	6 3x3 montages		
Wed Jan  8 15:02:02 2014	Cindi	01/08	8:30 AM	3:00 PM	tiltrotate	rttomo	giardia snmts	6		vacuum at 37, cleaned Oring, OK now
Thu Jan  9 11:16:15 2014	Cindi	01/09	7:15 AM	11:15 PM	tiltrotate	rttomo	giardia snmts	8		
Thu Jan  9 16:18:36 2014	Minghua	01/09	11:00 AM	4:30 PM	f20cryo	cryoimaging	Pombe	0	Crashed twice. Used F30 Cryo holder 
first and then F20 Cryo holder.	
Fri Jan 10 15:24:42 2014	Cindi	01/10	9:15 AM	3:30 PM	tiltrotate	rttomo	giardia snmts	11		
Sun Jan 12 16:25:37 2014	eileen	1/12	10:15 AM	4:20 PM	tiltrotate	rttomo	C. elegans anaphase	3 3x3 montages		
Mon Jan 13 17:19:15 2014	eileen	1/13	8:45 AM	5:15 PM	tiltrotate	rttomo	C.elegans anaphase supermontage	4 3x3 montages		
Mon Jan 20 18:04:51 2014	spilman	1/13	8:00 AM	5:00 PM	standard	serviceman	test	0		Replace objective lens coils & 
flush water lines
Mon Jan 20 18:05:04 2014	spilman	1/14	8:00 AM	5:00 PM	standard	serviceman	test	0		Replace objective lens coils & 
flush water lines
Mon Jan 20 18:05:15 2014	spilman	1/15	8:00 AM	5:00 PM	standard	serviceman	test	0		Replace objective lens coils & 
flush water lines
Mon Jan 20 18:05:26 2014	spilman	1/16	8:00 AM	5:00 PM	standard	serviceman	test	0		Replace objective lens coils & 
flush water lines
Mon Jan 20 18:05:39 2014	spilman	1/17	8:00 AM	5:00 PM	standard	serviceman	test	0		Replace objective lens coils & 
flush water lines
Mon Jan 20 18:05:54 2014	spilman	1/20	8:00 AM	5:00 PM	standard	serviceman	test	0		Replace objective lens coils & 
flush water lines
Wed Jan 22 15:52:39 2014	eileen	1/22	9:00 AM	3:50 PM	tiltrotate	rttomo	Drosophila and yeast	2		
Thu Jan 23 16:01:12 2014	MINGHUA	01/23	11:30 AM	4:30 PM	f30cryo	cryoimaging	Pombe		crashed when putting the single 
tilt rod back in after viewing.	
Fri Jan 24 17:36:46 2014	eileen	1/24	8:30 AM	5:30 PM	tiltrotate	rttomo	ellutriated yeast	1		HT and FEG were off when I logged 
on. Focus issues.
Tue Jan 28 13:12:46 2014	Cindi	01/27	1:00 PM	2:15 PM	standard	serviceman	alignments			
Tue Jan 28 13:33:33 2014	Cindi	01/28	1:00 PM	1:30 PM	standard	serviceman	alignments			
Wed Jan 29 10:22:09 2014	Chuck Moe	1/29/14	08:15 AM	10:30 AM	standard	serviceman	alignments EFTEM			
Wed Jan 29 13:52:31 2014	Cindi	01/29	1:14 PM	1:45 PM	standard	serviceman	scope down			HT is off, but reads 60kV and 50uA
Wed Jan 29 13:52:33 2014	Cindi	01/29	1:14 PM	1:45 PM	standard	serviceman	scope down			HT is off, but reads 60kV and 50uA
Thu Jan 30 15:53:32 2014	Spilman	1/30	8:00 AM	4:00 PM	standard	serviceman	none	0		Troubleshoot HT

System down
Tue Feb  4 15:17:51 2014	Spilman	2/3	9:00 AM	3:30 PM	standard	serviceman	Service	0		High dark current when HT Off.  
Replace HT feedback board and 
aligned scope as needed.  tourble 
shoot beam instability. Issue went 
away and cause not found.
Tue Feb  4 15:18:13 2014	Spilman	2/4	9:00 AM	3:30 PM	standard	serviceman	Service	0		High dark current when HT Off.  
Replace HT feedback board and 
aligned scope as needed.  tourble 
shoot beam instability. Issue went 
away and cause not found.
Wed Feb  5 17:14:50 2014	eileen	2/5/14	9:00 AM	5:10 PM	tiltrotate	rttomo	Moore 596 yeast	12	worked great!	
Fri Feb  7 16:55:20 2014	Cynthia	02/07	11:00 AM	5:00 PM	tiltrotate	rtimage	prasad exp4			
Mon Feb 10 17:28:46 2014	Cynthia	02/10	12:00 PM	5:30 PM	tiltrotate	rttomo	prasad exp 4 sample 3	5		
Tue Feb 11 18:32:41 2014	eileen	2/11	10:00 AM	6:30 PM	tiltrotate	rttomo	Kristy wt ellute	3		
Wed Feb 12 17:02:50 2014	Cynthia	02/12	10:30 AM	5:00 PM	tiltrotate	rttomo	Prasad exp4 sample 6 grid 17c	4		
Mon Feb 24 14:10:16 2014	FEI	2/14	8:00 AM	5{00 PM	standard	serviceman	Test	0		Replaced defective lens coil. 
reassembled, and aligned.
Mon Feb 24 14:10:32 2014	FEI	2/15	8:00 AM	5{00 PM	standard	serviceman	Test	0		Replaced defective lens coil. 
reassembled, and aligned.
Mon Feb 24 14:10:45 2014	FEI	2/16	8:00 AM	5{00 PM	standard	serviceman	Test	0		Replaced defective lens coil. 
reassembled, and aligned.
Mon Feb 24 14:10:56 2014	FEI	2/17	8:00 AM	5{00 PM	standard	serviceman	Test	0		Replaced defective lens coil. 
reassembled, and aligned.
Mon Feb 24 14:11:15 2014	FEI	2/18	8:00 AM	5{00 PM	standard	serviceman	Test	0		Replaced defective lens coil. 
reassembled, and aligned.
Mon Feb 24 14:11:49 2014	FEI	2/21	8:00 AM	5{00 PM	standard	serviceman	Test	0		Replaced defective lens coil. 
reassembled, and aligned.
Wed Feb 26 17:43:59 2014	Cynthia	2/25	12:30 PM	6:00 PM	tiltrotate	rttomo	Prasad exp 6 sample 6 grid a2	4		
Wed Feb 26 17:45:13 2014	Cynthia	2/26	12:00 PM	5:45 PM	tiltrotate	rttomo	Prasad exp 6 sample 6 grid a2	5		
Thu Feb 27 17:27:54 2014	Cynthia	02/27	9:30 AM	5:30 PM	tiltrotate	rttomo	Prasad exp 6 sample 6 grid a4	6	Eucentricity is way off between the 
standard and the t/r - don't 
remember that being the case. Must 
set tilt axis offset by around 6 
for t/r rod to be eucentric. Some 
issues with auto focus only 
measuring a % of actual focus at 60 
degrees mag 31k and tilt series 
stopped. OK went down in mag to 23K	
Fri Feb 28 18:03:59 2014	Cynthia	02/28	11:00 AM	6:15 PM	tiltrotate	rttomo	Prasad exp 6 sample 6 grida3	9		
Mon Mar  3 18:00:56 2014	Byungho Kang	03/04	8:20 AM	6:00 PM	tiltrotate	rttomo	Celegans	20	works good	none
Tue Mar  4 17:46:41 2014	Byungho Kang	03/05	9:10 AM	5:40 PM	tiltrotate	rttomo	Celegans, Arabidopsis	18	Worked ok other than the vacuum.	Vacuum degraded slowly to reach 20 
after collecting 2 tilt series.
Wed Mar  5 18:13:16 2014	Byungho Kang	03/05	10:00 AM	06:10 PM	tiltrotate	rttomo	Celegans, Arabidopsis	18	The vacuum degradation problem. 
Other than that, it worked ok.	
Thu Mar  6 11:08:29 2014	Ron Spilman	03/06	8:45 AM	11:00 AM	standard	serviceman	diagnosing leak in compustage area			line needs replacing maybe V8
Thu Mar  6 18:38:46 2014	Cindi, Eva, Alexa	03/06	11:15 AM	6:30 PM	tiltrotate	rttomo	muscle	4		
Fri Mar  7 13:20:40 2014	Spilman	3/7	9:00 AM	1:30 PM	standard	serviceman	none	0		Replace & adjust V8 valve, Replace 
vacuum line between V8 &V 41
Fri Mar  7 18:00:42 2014	eileen and Ava/Alexa	3/7/14	2 PM	6:00 PM	tiltrotate	rttomo	heart muscle montages	2 1x2		
Mon Mar 10 14:52:31 2014	Cindi	03/07	9:00 AM	2:45 PM	tiltrotate	rttomo	neurons	2	sample had to be destained	tilt-axis offset high (6 um)
Tue Mar 11 15:45:48 2014	Minghua	03/11	10:00 AM	16:00 PM	tiltrotate	rttomo	Pombe	2		
Wed Mar 12 16:54:19 2014	eileen and Eva/Alexa	4/11	8:25 AM	5:00 PM	tiltrotate	rttomo	heart muscle	3 2x2 montages		
Fri Mar 14 14:24:37 2014	Cindi,Cynthia,Eva	03/13	9:00 AM	8:45 PM	tiltrotate	rttomo	muscle	10		tilt-axis offset high
Fri Mar 14 18:09:10 2014	eileen	3/14	8:45 AM	6:05 PM	tiltrotate	rttomo	Vogel wt elutriated	11		
Tue Mar 18 18:07:30 2014	eileen	3/18	9:00 AM	6:05 PM	tiltrotate	rttomo	wt Jayne/Moore lab and DeCamilli 
neurons	3 1x3 montage and 4 regular tilt series		
Thu Mar 20 17:32:50 2014	Eileen/Cynthia/Eva	031914	9:00 AM	6:45 PM	tiltrotate	rttomo	Peter Kohl samples	3 1X2 duals, 1 dual		
Thu Mar 20 17:34:39 2014	Mary/Cynthia/Eva	03/14	9:45 AM	5:30 PM	tiltrotate	rttomo	Peter Kohl samples cardiomyocytes	3 duals		
Fri Mar 21 17:48:13 2014	eileen and Eva	3/21	1:00 PM	5:45 PM	tiltrotate	rttomo	heart muscle Kohl lab	4 1x2 montage		
Mon Mar 24 15:35:56 2014	eileen	3/24	8:30 AM	3:30 PM	tiltrotate	rtimage	Yumei/De Camilli neurons	10		
Tue Mar 25 15:31:31 2014	eileen	3/25	9:30 AM	3:30 PM	tiltrotate	rttomo	Yumei/De Camilli neurons	8		
Wed Mar 26 18:01:24 2014	Mary/Eva/Cynthia	03/26	8:45 AM	6:00 PM	tiltrotate	rttomo	cardiac cells	3, 1X2 duals		
Thu Mar 27 18:01:46 2014	Mary/Eva/Cynthia	03/27	9:00 AM	6:00 PM	tiltrotate	rttomo	cardiac cells	2, 1X2 duals 2 duals, 2single		
Tue Apr  1 14:15:48 2014	Cindi	04/01	9:00 AM	1:00 PM	tiltrotate	rttomo	neurons 14-day	2		no issues with inserting holder
Wed Apr  2 18:04:46 2014	Mary/Eva/Cynthia	04/02	9:00 AM	6:00 PM	tiltrotate	rttomo	Cardiac cells	4 duals		
Wed Apr  2 19:01:06 2014	Minghua	04/02	6:00 PM	7:00 PM	tiltrotate	rtimage	pombe			
Thu Apr  3 18:09:23 2014	Cindi/Eva/cynthia	04/03	9:00 AM	6:00 PM	tiltrotate	rttomo	Cardiac Cells	3, 1X3 duals, 1, 1X1 dual		
Fri Apr  4 16:14:25 2014	Minghua	04/04	9:50 AM	4:15 PM	tiltrotate	rttomo	Pombe	3		
Mon Apr  7 17:14:11 2014	eileen	4/7	11:30 AM	5:10 PM	tiltrotate	rttomo	Jaspersen EM1350 yeast	6		
Wed Apr  9 18:07:36 2014	Cindi?Eva/Cynthia	04/09	8:30 AM	6:00 PM	tiltrotate	rttomo	Cardiac Cells	3, duals, 2, 1X2 duals, 1 single axis		
Thu Apr 10 18:07:48 2014	Mary/Eva/cynthia	0410	9:00 AM	6:00 PM	tiltrotate	rttomo	Cardiac	4dual, 1, 1X2 dual	heard the turbo(?) wind down sound 
when I removed the T/R rod	
Mon Apr 14 16:30:29 2014	eileen	4/14	8:30 AM	4:30 PM	tiltrotate	rttomo	Jaspersen EM1350	9		
Wed Apr 16 17:28:10 2014	eileen	4/16	9:15 AM	5:30 PM	tiltrotate	rttomo	Jaspersen EM1350	10		
Thu Apr 17 19:06:07 2014	Matthew West	4/17	9:00 AM	7:00 PM	tiltrotate	rttomo	yeast mutants for Gia Voeltz	12		
Fri Apr 18 18:35:16 2014	Cynthia	04/18	10:00 AM	6:30 PM	tiltrotate	rttomo	prasad exp 7 red 	3 duals, 4 single axis	Eucentricity was way off (LD about 
6)with T/R rod as usual of late.	
Mon Apr 21 20:46:37 2014	Cynthia	04/21	9:45 AM	8:45 PM	tiltrotate	rttomo	prasad exp 7 blue	6 duals		
Thu Apr 24 17:46:38 2014	Cynthia	04/24	10:30 AM	5:30 PM	tiltrotate	rttomo	prasaad exp 7 blue	4 duals 2 singles		
Fri Apr 25 13:05:02 2014	morphew	4-25	9 AM	1 PM	f20cryo	cryoimaging	vit sections	0		3 vac crashes
Mon Apr 28 18:08:47 2014	Cynthia	04/30	10:00 AM	6:00 PM	tiltrotate	rttomo	prasad exp 7 blue grid a3	4 duals	Yah, no crashes!	
Tue Apr 29 12:45:07 2014	Cindi	04/28	8:30 AM	12:45 PM	tiltrotate	rttomo	neurons	4		
Wed Apr 30 17:17:47 2014	Cindi	04/29	11:00 AM	5:15 PM	tiltrotate	rttomo	neurons	3		
Thu May  1 15:35:12 2014	Cindi	5/1	8:30 AM	5:30 PM	tiltrotate	rttomo	7			compustage wonky. service coming 
next week.
Fri May  2 18:04:21 2014	eileen	5/2	9 AM	6 PM	tiltrotate	rttomo	CDC20 yeast	6 2x2 montages		
Wed May  7 13:24:30 2014	Cindi	05/07	8:30 AM	1:30 PM	tiltrotate	rttomo	Bacteria	6		
Thu May  8 13:59:58 2014	Cindi	05/08	9:30 AM	2:00 PM	tiltrotate	rttomo	bacteria	5		
Fri May  9 16:39:02 2014	eileen	5/9	10:30 AM	4:30 PM	tiltrotate	rttomo	Moore wt anaphase	4		
Tue May 13 12:48:20 2014	David	05/12	10:00 AM	11:00 AM	standard	serviceman	fix rotation issue			
Tue May 13 12:48:51 2014	Cindi	05/12	11:00 AM	1:00 PM	tiltrotate	rttomo	neurons +10 nm gold	2		david had to fix SerialEM
Tue May 13 17:13:16 2014	eileen	5/13	1 PM	5:10 PM	tiltrotate	rttomo	Moore 596_wt	4		
Wed May 14 13:47:25 2014	eileen	5/14	9 AM	1:30 PM	tiltrotate	rttomo	CDC20	6		
Thu May 15 15:35:10 2014	morphew	5/15	11 AM	3:30 PM	f30cryo	cryoimaging	vitreous sections U2OS			
Fri May 16 22:38:03 2014	Matt West	5/16	9:30 AM	10:30 PM	tiltrotate	rttomo	Yeast mutants	12		
Mon May 19 17:10:52 2014	eileen	5/19	8:45 AM	5:05 PM	tiltrotate	rttomo	CDC20	6		
Wed May 21 17:22:35 2014	Cynthia	05/21	9:30 AM	5:15 PM	tiltrotate	rttomo	prasad exp7 blue	7	Turbo kicked on each time I removed 
the T/R rod. This is unusual.	Turbo kicked on each time I removed 
the T/R rod. This is unusual.
Thu May 22 15:21:52 2014	morphew	5/22	8:45 AM	3:15 PM	tiltrotate	cryoimaging	pombe			
Fri May 23 17:48:05 2014	eileen	5/23	8:45 AM	5:50 PM	tiltrotate	rttomo	CDC20	10		
Tue May 27 16:10:26 2014	Cynthia	05/27	10:00 AM	3:30 PM	tiltrotate	rttomo	prasad exp7 blue 	7	airlock pumped a second time after 
inserting standard rod	airlock pumped out a second time 
after inserting standard rod at end 
of session
Thu May 29 15:38:38 2014	Cynthia	05/29	9:30 AM	3:30 PM	tiltrotate	rttomo	Prasad exp7 blue	9		
Fri May 30 17:31:27 2014	Cynthia	05/30	9:30 AM	5:15 PM	tiltrotate	rttomo	prasad blue exp 7	10		putting in standard rod was asked 
2X to select rod, and the airlock 
was prepumped 2X.

Additionally when I first put in 
the T/R rod the system did not 
allow me to select a rod, ie it 
appeared not to know the rod had 
been inserted. I slippped the rod 
out slightly and returned it. Then 
the rod was detedted.
Mon Jun  9 15:52:33 2014	morphew	6/9	2:30 PM	3:45 PM	standard	rtimage	shuhb boranephosphanate			
Tue Jun 10 17:22:41 2014	Cynthia	06/10	11:00 AM	5:15 PM	tiltrotate	rtimage	cytoskeleton		Scope worked great!  Did have a 
blip due to power outage.	
Wed Jun 11 17:29:04 2014	Cynthia	06/11	10:00 AM	5:15 PM	tiltrotate	rttomo	Giardia cytoskeletons	1 almost		I did Eucentricity with the 
standard rod for alignments, then 
the lateral displacement was 4.25 
when I put in the T/R rod.
Thu Jun 12 17:13:40 2014	Cynthia	06/12	11:00 AM	5:00 PM	tiltrotate	rttomo	Cytoskeletons	1	The F30 was very well behaved 
today, I think that is because 
Cindor was here!	
Wed Jun 18 14:16:24 2014	Cynthia	06/17/14	0:0 AM	0:0 PM	standard	development	NONE please see note		On June 11th, the high tension was
off in the morning after an cryo
cycle done the night before	
Thu Jun 19 10:59:02 2014	Cynthia	06/19	10:20 AM	10:50 AM	standard	development	trouble shoot		I had trouble with 4X4 montage, 
looked like every other panel in Y 
was out of focus. Was going to do a 
1X5 montage to begin to determin 
problem, however the beam was, 
shaky and pulsing larger then 
smaller.	During alignments beam was shaking 
and pulsing larger, smaller. David 
said to call for service.
Tue Jun 24 15:07:04 2014	Chuck Moe	6/23	09:00 AM	5:00 PM	standard	serviceman	service		replaced upper obj lens	completed upper obj lens 
replacement and verified proper 
operation
Tue Jun 24 15:08:36 2014	Chuck Moe	6/24	9:00 AM	4:00 PM	standard	serviceman	service		conditioned column and HT. Checked 
for proper operation after lens 
coil replacement	conditioned column and HT. Checked 
for proper operation after lens 
coil replacement
Tue Jun 24 17:28:23 2014	DMN	6/24	1:00 PM	1:01 PM	standard	development	testing log mailing			Here is a test message to mary and
eileen
Tue Jun 24 20:11:10 2014	DNM	6/24	1:02 PM	1:03 PM	standard	development	email test 2			Here is a test problem sent to mary,
eileen, and mast
Tue Jun 24 20:13:18 2014	DNM	6/24	1:02 PM	1:03 PM	standard	development	email test 3			Here is a test problem sent to mary,
eileen, and mast@bio3d
Tue Jun 24 20:14:06 2014	DNM	6/24	1:02 PM	1:03 PM	standard	development	email test 4			Here is a test problem sent to 
mast@bio3d
Tue Jun 24 20:28:55 2014	DNM	6/24	1:02 PM	1:03 PM	standard	development	email test 5			Here is a test problem sent to  mary
and eileen
Tue Jun 24 20:42:01 2014	DNM	6/24	1:05 PM	1:06 PM	standard	development	Test email with subject?			Having no subject is a problem too
Thu Jun 26 15:50:10 2014	Chuck	06/26	8:30 AM	4:00 PM	tiltrotate	rttomo	E. coli	2		
Mon Jun 30 20:57:37 2014	chuckr	06/30	8:30 AM	9:00 PM	tiltrotate	rttomo	e coli seg mutant minicells	13		
Tue Jul  1 20:55:31 2014	chuckr	07/01	10 AM	9 PM	tiltrotate	rttomo	e coli seq mutant minicells	14		tui became non-responsive.  
computer reboot required... done by 
Chuck of FEI.
Wed Jul  2 07:53:53 2014	Cynthia	06/27	12:00 PM	7:00 PM	tiltrotate	rttomo	Giardia neg stain	1, 4X4		
Wed Jul  2 14:18:45 2014	Cynthia	07/02	8:00 AM	2:00 PM	tiltrotate	rttomo	Giardia	1, 4X4		
Thu Jul 10 20:33:36 2014	chuckr	07/10	8:00 AM	8:30 PM	tiltrotate	rttomo	e coli seg mutants	10	Good.	
Mon Jul 14 13:43:54 2014	David/Eileen	7/10	12 PM	6:00 PM	standard	serviceman	Gif service	0		The GIF was at room temp. Camera 
cooling panel in controller burned 
out. Switched out the entire camera 
controller with spare from Jensen 
lab. Cooled to -40
Mon Jul 14 13:46:34 2014	David/Eileen	7/14	9:30 AM	1:30 PM	standard	serviceman	Gif quadrant calibration	0	GIF quadrant correction went well. 
New gain reference acquired. Things 
seem ok with the GIF	
Tue Jul 15 17:45:05 2014	eileen	7.15	3:15 PM	5:40 PM	tiltrotate	rtimage	mapping strategies	0		
Wed Jul 16 17:09:38 2014	eileen	7/16	9:00 AM	5:00 PM	tiltrotate	rtimage	GIF tuning, etc. also mapping 
strategies	0		
Thu Jul 17 14:36:18 2014	morphew	7/17	10 AM	2:30 PM	f20cryo	cryoimaging	synaptophysin		Two vac crashes when inserting cryo 
rod, one success. Gif needed -10eV 
energy shift to align ZLP. 	
Fri Jul 18 18:30:11 2014	Matthew West	7-18-14	12:00 PM	6:15 PM	tiltrotate	rttomo	Cos mutant tomography	6		The washer for the tilt rotate is 
very sticky, so sticky that it bent 
and ruined one of my grids just 
trying to remove it.  Probably just 
needs an etoh/acetone 
bath...Otherwise the scope ran 
flawlessly.
Mon Jul 21 16:50:13 2014	Cynthia	07/21	9:00 AM	4:00 PM	tiltrotate	rtimage	Giardia cyto skel morpholino			
Thu Jul 24 08:16:25 2014	chuckr	07/23	8:10 AM	8:30 PM	tiltrotate	rttomo	wt ecoli	10	good!	
Thu Jul 24 16:32:20 2014	chuckr	07/24	8:15 AM	4:30 PM	tiltrotate	rttomo	wt e coli	10	great!	
Mon Jul 28 16:07:45 2014	eileen	7/26	8:30 AM	4:00 PM	tiltrotate	rttomo	C. elegans supermontage	3 3x3	stage tracking error, but otherwise 
fine.	
Mon Jul 28 16:09:18 2014	eileen	7/28	8:30 AM	3:00 PM	tiltrotate	rttomo	C. elegans supermontage	2 3x3	stage making clicking sounds. 
Autofocus errors, not keeping focus 
with montage panels. tilt series 
not usable. Service called.	
Mon Aug  4 14:36:36 2014	FEI Service	7/29	9:15 AM	5:00 PM	standard	serviceman	checking objective lens	0	objective lens failure. Chuck Moe 
replaced upper objective lens and 
did alignments 7/29-8/1	
Mon Aug  4 14:39:33 2014	eileen	8/4	9:00 AM	12:00 PM	standard	serviceman	Candida	0	Power to objective lens went out 
during direct alignments. Called 
Chuck/service. Power supply panel 
reading 0 volts/0 amps. Chuck 
ordered new power supply will 
replace tomorrow (8/5).	
Mon Aug 25 17:10:40 2014	eileen	8/25	9:00 AM	5:05 PM	tiltrotate	rttomo	Jing pilot tomos	2		
Tue Aug 26 15:04:56 2014	Minghua	08/26	1:30 PM	3:00 PM	standard	rtimage	Pombe			
Wed Aug 27 17:11:41 2014	eileen	8/27	9:00 AM	5 PM	tiltrotate	rttomo	Wu pombe wt	4 2x1 montage		
Fri Aug 29 17:48:13 2014	Matt West	9/29	10:30 AM	5:45 PM	tiltrotate	rttomo	Yeast mutants	8		A combined C2 and beam alignment 
problem held me back for an hour, 
but thanks to Eileen's help I was 
able run my tilt series.  Thanks 
Eileen for your help!!!
Tue Sep  2 09:27:31 2014	eileen	8/28	10 AM	11 AM	standard	training	Kim Erickson intro	0		
Tue Sep  2 17:13:22 2014	eileen and Kim Erikson	9/2	9:00 AM	4:00 PM	tiltrotate	training	Kim's KDE virus-infected cells	2 1x3 montage		one panel of montage completely 
black two times, then not again. 
sample very unstable at high minus 
tilt. Needed large drift settling.
Thu Sep  4 15:41:06 2014	eileen and Kim Erikson	9/4	10 AM	3 PM	tiltrotate	training	virus infected cells	2 1x3 montage		
Tue Sep  9 17:09:57 2014	kim erickson	9/9	10 AM	4:00 PM	tiltrotate	rttomo	virus infected cells	4 1x3 montages		
Wed Sep 10 15:21:55 2014	eileen	9/10	8:45 AM	3:00 PM	tiltrotate	rttomo	Wu pombe mutant 'pilot'	2 1x2 montage		
Thu Sep 11 17:45:41 2014	eileen	9/11	9 AM	5:30 PM	tiltrotate	rttomo	Jing antibody treated infected cells	3		
Tue Sep 16 14:28:51 2014	Minghua	9/16/2014	11:00 AM	2:30 PM	standard	rtimage	Pombe		Crashed once and check the rod's 
ring after that. Everything went 
fine when putting the rod back in 
finally.	
Wed Sep 17 15:24:39 2014	morphew	9/18	10 AM	3 PM	tiltrotate	rttomo	ackerson protein crystals	3		one vac crash
Wed Sep 17 17:35:08 2014	Minghua	09/17/2014	4:00 PM	5:40 PM	standard	rtimage	Pombe			
Thu Sep 18 17:31:26 2014	eileen	9/18	9 AM	5:30 PM	tiltrotate	rttomo	Candida wt	5		
Wed Sep 24 16:23:26 2014	eileen	9/24	9:00 AM	4:15 PM	tiltrotate	rttomo	Candida wt	3		
Mon Sep 29 13:13:23 2014	mary/Feldheim	9/29	11 AM	1 PM	standard	rtimage	gold lattice	0		
Tue Sep 30 17:03:57 2014	kim erickson	9/30	10 AM	4:30 PM	tiltrotate	rttomo	virus infected cells	4 2x2 montages		
Wed Oct  1 16:48:38 2014	kim erickson	10/1	10 AM	4:30 PM	tiltrotate	rttomo	virus infected cells	3 3x3 montages		
Thu Oct  2 15:01:03 2014	kim erickson	10/2	10:00 AM	12:15 PM	tiltrotate	rttomo	virus infected cells	1 3x3 montage		
Fri Oct  3 15:16:04 2014	eileen	10/3	9:00 AM	3:00 PM	tiltrotate	serviceman	making guest account and testing	1		
Tue Oct  7 17:54:29 2014	Eva	10/7	9 AM	5:30 PM	tiltrotate	rttomo	heart muscle	10 single frame, 2 1x2 montage		
Mon Oct 13 17:47:36 2014	morphew/eva	10/13	8:30 AM	5:40 PM	tiltrotate	rttomo	muscle	9		
Tue Oct 14 17:50:35 2014	eva (kohl)	10/14	8:30 AM	5:45 PM	tiltrotate	rttomo	heart muscle	6 single frame, 4 1x2 montage 		
Thu Oct 16 17:36:16 2014	eva (kohl)	10/16	8:30 AM	5:30 PM	tiltrotate	rttomo	hear muscle	6 single frame, 2 1x3 montage		
Fri Oct 17 16:39:18 2014	eileen	10/17	8:25 AM	4:30 PM	tiltrotate	rtimage	Candida	0		
Mon Oct 20 16:34:45 2014	morphew/eva	10/20	8:30 AM	4:30 PM	tiltrotate	rttomo	muscle	6 montages		
Wed Oct 22 17:45:19 2014	eva (kohl)	10/22	9 AM	5:30 PM	70	rttomo	heart	3,1X3, 1 - 3X3		
Wed Oct 22 21:40:40 2014	DNM	10/19	2:00 PM	11:30 PM	standard	development	continuous acquisition			
Wed Oct 22 21:41:03 2014	DNM	10/20	4:30 PM	11:30 PM	standard	development	continuous acquisition			
Fri Oct 24 18:07:28 2014	eva (kohl)	8/24	8:30 AM	6:06 PM	tiltrotate	rttomo	heart muscle	10		
Tue Oct 28 18:04:17 2014	eva (kohl)	10/28	8:30 AM	6 PM	tiltrotate	rttomo	heart muscle	9 (5 single, 2 1x2 montage, 2 3x2 montage		The scope was asking for rod 
selection again when inserting the 
rod. Did not recognize the t/r 
holder in the morning, but then no 
issues the rest of the day with t/r 
or standard rod.
Wed Oct 29 17:24:19 2014	eva (kohl)	10/29	8:30 AM	5:30 PM	tiltrotate	rttomo	Heart muscle		All went well with the scope today. 
Actual hours of beam time 8	
Thu Oct 30 16:35:58 2014	eva	10/30	9 AM	3 PM	tiltrotate	rttomo	heart muscle	5		The grid was not screwed in tightly 
in the T/R holder. Grid, hex ring, 
washer fell out of holder when 
holder was removed. Called service.
Mon Nov  3 17:04:01 2014	eva (kohl)	11/3	9:00 AM	5:00 PM	tiltrotate	rttomo	muscle			
Tue Nov  4 17:13:15 2014	eva (kohl)	11/4	8 AM	5:15 PM	tiltrotate	rttomo	heart muscle	6 (including 3x1 adn 2x1 montages)		
Wed Nov  5 17:21:09 2014	eva	11/5	9:30 AM	5:00 PM	tiltrotate	rttomo	heart			
Thu Nov  6 17:32:31 2014	eileen	11/6	9 AM	5:30 PM	tiltrotate	rttomo	vogel elutriated yeast	10		
Mon Nov 10 11:01:02 2014	Chuck Moe	11/10	0830 AM	1100 AM	standard	serviceman	adjustment of obj aperture		adjusted obj aperture	
Mon Nov 10 17:10:07 2014	eva (Kohl)	11-10	12:00 PM	5:15 PM	tiltrotate	rttomo	muscle			
Tue Nov 11 15:53:30 2014	eva (Kohl)	11/11	8:30 AM	3 PM	tiltrotate	rttomo	heart muscle	4 (2x1 and 4x1 montages)		
Wed Nov 12 17:26:28 2014	eva/kohl	11/12	9:30 AM	4:30 PM	tiltrotate	rttomo	heart cells	4 duals	When I closed Filter control app 
got an error message saying start 
up Fliter Control before Digital 
Micrograph! I did this this morning.	
Thu Nov 13 17:15:41 2014	ron	11/13	11 AM	1 PM	tiltrotate	serviceman	problem with objective lens/power 
suppy	0		objective lens power off- changed 
water filter and scope is back up
Thu Nov 13 17:15:43 2014	ron	11/13	11 AM	1 PM	tiltrotate	serviceman	problem with objective lens/power 
suppy	0		objective lens power off- changed 
water filter and scope is back up
Thu Nov 13 17:16:47 2014	eva (Kohl)	11/13	1:30 PM	5:30 PM	tiltrotate	rttomo	hear muscle	6 (including montages)		
Fri Nov 14 15:36:22 2014	eva (Kohl)	11/14	8:30 AM	3:30 PM	tiltrotate	rttomo	heart muscle	4 (2 1x2 and 2 2x3 montage)		
Wed Nov 19 13:14:34 2014	eileen	11/18	8:45 AM	1 PM	tiltrotate	rttomo	Candida- kip1 delete	0		objective lens power went out 
again. Service has been called.
Wed Dec 10 16:56:18 2014	eileen	12/10	8:45 AM	5 PM	tiltrotate	rttomo	cooper pombe mutant pink	8		
Wed Dec 31 15:49:17 2014	eileen	12/31	8:45 AM	3:50 PM	tiltrotate	rttomo	WT elute Kristy	6		the objective lens chiller went out 
and shut off lenses for no apparent 
reason. called ron, restarted 
chiller and all worked well the 
rest of the day.
Tue Jan  6 16:08:40 2015	eileen	1/6	1 PM	3 PM	tiltrotate	rttomo	screening grids for Kamesh's visit	2		
Wed Jan  7 17:08:49 2015	eileen/kamesh	1/7	2:00 PM	5:05 PM	tiltrotate	rttomo	rice leaf	2		
Thu Jan  8 17:18:43 2015	eileen/kamesh	1/7	8:45 AM	5:00 PM	tiltrotate	rttomo	rice leaf	6		
Fri Jan  9 15:58:18 2015	eileen/kamesh	1/9	10 AM	4 PM	tiltrotate	rttomo	rice leaves	2		
Tue Jan 13 16:51:38 2015	eileen/cooper	1/13	11:30 AM	4:50 PM	tiltrotate	rttomo	cooper pombe 'yellow'	8		
Thu Jan 15 15:11:42 2015	Janet Meehl Fox/Winey	1/15	9:15 AM	2:45 PM	tiltrotate	rttomo	Tetrahymena sfr13Delete basal bodies	6		
Fri Jan 16 13:29:58 2015	morphew	1_16	9:40 AM	1:30 PM	tiltrotate	rttomo	Cooper Pombe mut	2		
Tue Jan 20 15:19:13 2015	mary(cooper)	1/20	9 AM	3:00 PM	tiltrotate	rttomo	Cooper pombe	4		
Wed Jan 21 15:33:46 2015	morphew	1/21	11:00 AM	3:30 PM	tiltrotate	rttomo	cooper wt	2		
Fri Jan 23 18:06:17 2015	eileen/cooper	1/23	9:00 AM	6:00 PM	tiltrotate	rttomo	cooper pombe wt	6 2x1 montage		
Tue Jan 27 16:01:33 2015	eileen/cooper	1/27	2:00 PM	4:00 PM	tiltrotate	rttomo	cooper pombe	1 1x2 montage		
Thu Jan 29 17:28:06 2015	eileen	1/29	9:00 AM	5:30 PM	tiltrotate	rttomo	Cooper wt pombe	8		
Fri Jan 30 18:16:17 2015	eileen	1/30	9:00 AM	6:00 PM	tiltrotate	rttomo	Cooper wt pombe and test batch 
series	10		
Sat Jan 31 14:19:01 2015	DNM	1/31/15	1:30  PM	2:20 PM	standard	development	Fixing diffraction mode setting 
through plugin			
Fri Feb 13 18:28:32 2015	eileen	2/13	9 AM	6:15 PM	tiltrotate	rttomo	cooper wt	2		vac crash when taking out standard 
holder. vac not recover due to 
compustage leak. Ron came and 
fixed. no other problems
Thu Mar  5 13:51:21 2015	eileen	3/5	10 AM	1 PM	tiltrotate	rttomo	Cooper mutant	6		
Mon Mar  9 16:47:47 2015	eva/kohl	3/9/15	8:45 AM	4:45 PM	tiltrotate	rttomo	heart	4(2-1x1, 2 4x1)		
Tue Mar 10 18:52:08 2015	eva	3/10	8:45 AM	6:50 PM	tiltrotate	rttomo	heart muscle	1x3 supermontage of 2x3 montages		
Wed Mar 11 17:00:02 2015	eva/kohl	3/11	9 AM	4:30 PM	tiltrotate	rttomo	heart muscle	1x2 supermontage of 2x4 montage		
Thu Mar 12 17:08:26 2015	eva (Kohl)	3/12	9 AM	4:30 PM	tiltrotate	rttomo	heart muscle	1x2 supermontage of 2x3 montage		
Fri Mar 13 16:38:55 2015	eva (Kohl)	3/13	9 AM	3:30 PM	tiltrotate	rttomo	heart muscle	2 2x2 montage		
Wed Mar 18 16:41:23 2015	eileen/Cooper	3/18	9 AM	4:30 PM	tiltrotate	rttomo	pombe mutants 'yellow'	8		
Thu Mar 19 14:09:22 2015	eileen/cooper	3/19	9:00 AM	1:30 PM	tiltrotate	rttomo	cooper mutant (yellow)	8		
Tue Mar 24 16:29:49 2015	eileen/Kohl	3/24	9:00 AM	4:30 PM	tiltrotate	rttomo	higher res cardiac contacts	10		
Fri Apr 17 11:51:59 2015	Service (Ron)	4/16	9 AM	5:30 PM	standard	serviceman	service on lens power	0	Power was out to all lenses. Ron 
did troubleshooting and found a 
loose connection in back panel.	
Fri Apr 17 11:52:44 2015	Service (Ron)	4/17	9 AM	11:50 AM	standard	serviceman	service on lens power	0	Power was out to all lenses. Ron 
did troubleshooting and found a 
loose connection in back panel.	
Fri Apr 17 17:33:57 2015	eileen	4/17	1:00 PM	5:30 PM	tiltrotate	rttomo	yeast tomos testing	6		
Tue Apr 21 16:41:41 2015	Ron	4/21	9 AM	12:15 PM	standard	serviceman	service	0	Ron reconnected lens power again	
Wed Apr 29 14:08:11 2015	DNM	4/29	1:25 PM	2:10 PM	standard	development	test precook montage		Installed new SerialEM and scope 
plugin	
Tue May  5 16:28:15 2015	eileen	5/1/15	9 AM	6 PM	tiltrotate	rttomo	dutcher (ramp down) cnk11-2	9		
Tue May  5 16:28:42 2015	eileen	5/5/15	9 AM	4:30 PM	tiltrotate	rttomo	dutcher (ramp down) cnk11-2	6		
Fri May  8 16:51:21 2015	eileen/McIntosh	5/8	11 AM	5 PM	tiltrotate	rttomo	mapping PTK2 spindle for 
supermontage	1 3x3		
Tue May 12 16:49:24 2015	eileen	5/12	10 AM	5 PM	tiltrotate	rttomo	evaluating Laura's COS cells 
(Voeltz)	0		
Fri May 15 15:20:08 2015	eileen	5/15	10 AM	3:20 PM	tiltrotate	rttomo	PTK2 single axis and GIF testing	1		
Thu May 21 15:07:16 2015	kim/eileen	5/21	8:45 AM	3 PM	tiltrotate	rttomo	virus-infected cells	6		
Sun May 24 12:11:14 2015	DNM	5/24	9:40 AM	12:15 PM	standard	development	testing pixel size routine and 
montage macro	0		scope-signup network wire hooked to 
wrong switch
Wed May 27 17:19:21 2015	Kim/eileen	5/27	8:45 AM	5:20 PM	tiltrotate	rttomo	virus-infected cells	6	Training Kim- one hour eot	
Thu May 28 15:54:16 2015	kim/eileen	5/28	8:45 AM	3:45 PM	tiltrotate	rtimage	stuff on cell- no tilt series	0		
Fri Jun  5 09:51:11 2015	eileen	6/2-6/3	9 AM	5 PM	standard	serviceman	scopes on standby for overnight 
electrical work	0		
Fri Jun  5 09:51:56 2015	eileen	6/4	9 AM	10 AM	standard	serviceman	scopes on standby briefly- tolin 
cancelled	0		
Tue Jun  9 17:01:19 2015	Cynthia	06/08	3:00 PM	5:00 PM	standard	training	training myself to use Gif!		I was not sucessful tuning the GIF. 
Message Found beam on slit but not 
on CCD Please check that the TV 
retracts and that the shutter works.
I did not log this use into the 
equipment log as beam time.	
Thu Jun 11 17:40:32 2015	Cynthia	06/11	11:00 AM	5:15 PM	tiltrotate	rttomo	Cos cells Plastic	1 dual	only 1.5 hours of beam time	
Mon Jun 15 16:33:24 2015	eileen	6/15	11 AM	4:30 PM	tiltrotate	rtimage	mapping PTK2 for supermontage	0		
Tue Jun 16 20:52:42 2015	eileen	6/16	9:00 AM	8:50 PM	tiltrotate	rttomo	PTK2 supermontage	6 3x3 montages		
Fri Jun 19 17:34:50 2015	Cynthia/Laura	06/15	10:30 AM	5;30 PM	tiltrotate	rttomo	cos cells	almost 1	Lots of problems. Insert rod didn't 
activate airlock pump, crashed 
vacuum. Was a fiber on oring, 
removed ok. Image shift(?)issue 
above 15K mag	Appeared to be image shift movement 
that did not correspond to 
mouse/stage movement at mags above 
15000. Coould not reliably move 
stage with mouse. Did tile at 15K, 
OK
Mon Jun 22 15:53:55 2015	eileen	6/22	11:00 AM	3:50 PM	standard	serviceman	redid image shift calibration and 
redid DM gain reference	0	There was a bad image shift 
calibration in the short term file. 
redid this from scratch.	
Tue Jun 23 17:24:45 2015	eileen	6/23	12:00 PM	5:30 PM	tiltrotate	rttomo	dutcher DB947	6		
Wed Jun 24 18:39:41 2015	Cynthia/Laura	06/24	9:30 AM	6:30 PM	tiltrotate	rttomo	cos cells	3.5 duals	Everything went fine!	
Tue Jun 30 17:16:46 2015	morphew	6/30	2:30 PM	5:15 PM	tiltrotate	rttomo	Ptk	2		
Wed Jul  1 13:09:43 2015	morphew	7/1	10:30 AM	12:15 PM	tiltrotate	rtimage	ptk			
Wed Jul  1 18:05:44 2015	eileen	7/1	3 PM	6 PM	tiltrotate	rttomo	dutcher DB20	9		
Thu Jul  2 15:47:04 2015	eileen	7/2	10 AM	3:00 PM	tiltrotate	rttomo	dutcher DB20	9		
Wed Jul  8 12:30:21 2015	Chuck Moe	7/8	0830 AM	1230 PM	f30cryo	serviceman	stage lockup recovery		Removed holder in safe state. reset 
EPD sensors on stage for proper id 
of holders. Tested cryo holder at 
60 degrees for issues. none noted.	
Thu Jul  9 12:06:34 2015	eileen	7/7	10 AM	2 PM	tiltrotate	rttomo	dutcher DB19	7	Holder jammed at high tilt	
Thu Jul  9 12:07:07 2015	eileen	7/8	3 AM	5 PM	tiltrotate	rttomo	dutcher DB19	3		
Thu Jul  9 12:07:33 2015	eileen	7/8	3 AM	5 PM	tiltrotate	rttomo	dutcher DB19	3		
Fri Jul 10 15:33:36 2015	kim	7/9	10 AM	4 PM	tiltrotate	rttomo	virus	4		
Fri Jul 10 15:34:01 2015	kim	7/10	10 AM	4 PM	tiltrotate	rttomo	virus	6		
Mon Jul 13 09:15:11 2015	Chuck Moe	7/13	0730 AM	0930 AM	standard	serviceman	service on P15 gauge		Replaced P15 gauge and restarted 
vacuum system	
Tue Jul 14 16:08:43 2015	morphew	7/14	2:30 PM	4:15 PM	tiltrotate	rttomo	ptk fix/freeze	2		
Wed Jul 15 16:48:17 2015	Cynthia/Laura	0715	10:00 AM	4:30 PM	tiltrotate	rttomo	cos cells transfected	1		
Fri Jul 17 15:01:44 2015	morphew	7/17	2:30 PM	3 PM	standard	rtimage	quick check			
Mon Jul 20 16:30:47 2015	Cynthia	0720	12:00 PM	3:30 PM	tiltrotate	rtimage	cos cells disc 3			
Fri Jul 24 13:03:52 2015	morphew	7/24	11 AM	1 PM	tiltrotate	rttomo	ptk	1		
Tue Jul 28 15:28:25 2015	morphew	7/29	1 PM	3:15 PM	tiltrotate	rttomo	mcintosh	2		
Thu Jul 30 10:43:36 2015	eileen	7/30	10 AM	11 AM	tiltrotate	rttomo	checking settings and images 
(service/maintenance)	1		
Thu Jul 30 15:53:51 2015	janet	7/29	12 PM	4 PM	tiltrotate	rtimage	mth spc29	0		
Thu Jul 30 15:54:31 2015	eileen	7/30	10 AM	3 PM	tiltrotate	rttomo	ptk2 chem/hpf 3x3 montage	1		
Fri Jul 31 17:24:10 2015	DNM	07/31	2:30 PM	5:30 PM	standard	development	realign to item bugs	0		
Sat Aug  1 13:25:54 2015	DNM	8/1	9:00 AM	1:30 PM	standard	development	realign to item			
Tue Aug 11 16:28:29 2015	eileen	8/11	3 PM	4:30 PM	tiltrotate	rtimage	setting up PTK2	0		
Wed Aug 12 20:53:15 2015	eileen	8/12	8:45 AM	9 PM	tiltrotate	rttomo	PTK2 supermontage sec2 McIntosh	2x3 super of 3x3 montages		
Fri Aug 14 14:46:22 2015	eileen	8/14	9 AM	2:30 PM	tiltrotate	rttomo	organoids 	5		
Mon Aug 24 16:18:05 2015	morphew	8/23	2:30 PM	4 PM	tiltrotate	rttomo	PTK DeLuca	1		
Tue Sep  1 10:30:43 2015	DNM	9/1	10:30 AM	10:31 AM	standard	development	test log			
Tue Sep  1 10:39:03 2015	DNM	9/1	10:30 AM	10:31 AM	standard	development	test log			
Wed Sep  2 11:57:30 2015	eileen/McIntosh	9/2	9:15 AM	11:45 AM	tiltrotate	rttomo	Deluca/McIntosh wt kinetochores	4		
Thu Sep  3 13:55:03 2015	eileen/McIntosh	9/3	9:00 AM	1:30 PM	tiltrotate	rttomo	McIntosh/Deluca 9A3	7		
Tue Sep 15 16:58:45 2015	eileen	9/15	2 PM	5:00 PM	tiltrotate	rttomo	Deluca/McIntosh 1.5 hours	2	CCD camera had warmed. Let cool for 
several hours before starting 
session.	
Wed Sep 16 16:58:05 2015	eileen	9/16	1:30 PM	5:00 PM	tiltrotate	rttomo	mcintosh/deluca 9A1 and wt1	6		
Tue Sep 22 16:20:16 2015	eileen	9/22	11:30 AM	3:00 PM	tiltrotate	rttomo	Eric Cole WT TET 	4 3x1 2 single		
Wed Sep 23 12:29:01 2015	eileen	9/23	10:30 AM	12 PM	tiltrotate	rttomo	Eric Cole WT Tet	2 1x2 montage		
Tue Sep 29 17:23:36 2015	eileen	9/29	4:00 PM	4:45 PM	tiltrotate	rttomo	Shanshan block12	2	vac took a while to come down	
Wed Sep 30 14:48:38 2015	eileen	9/30	12:45 PM	2:30 PM	tiltrotate	rttomo	Shanshan block12 cell1	3		
Fri Oct  2 17:16:43 2015	DNM	10/1	12:01 PM	3:30 PM	standard	development	Defocus from FFT			
Fri Oct  2 17:17:04 2015	DNM	10/2	12:01 PM	4:00 PM	standard	development	Defocus from FFT			
Wed Oct  7 16:38:52 2015	eileen	10/7	2:00 PM	4:30 PM	tiltrotate	rttomo	Shanshan block11 cell1	5		
Fri Oct  9 16:15:44 2015	eileen	10/9	12:30 PM	3:30 PM	tiltrotate	rttomo	Cole Tet 45min and autoacquire 
testing	3		
Sun Oct 11 12:15:37 2015	eileen	10/11	9:30 AM	12:00 PM	tiltrotate	rttomo	Eric Cole class demo	1 1x2 montage		
Wed Oct 21 16:11:16 2015	eileen	10/21	10:30 AM	3:00 PM	tiltrotate	rttomo	Shanshan B1_1 adn B_2 plus auto 
testing	6		
Fri Oct 23 14:07:09 2015	eileen	10/23	11 AM	1:30 PM	tiltrotate	rttomo	Shanshan B3_1 and automation testing	4		
Fri Oct 30 15:35:11 2015	Chuck Moe	10/30	07 AM	16:00 PM	standard	serviceman	PM		PM	PM
Mon Nov  9 14:26:41 2015	eileen	11/9	12:00 PM	2:15 PM	tiltrotate	rttomo	Shanshan B3 cell1 grid I3	4		
Mon Nov  9 14:28:24 2015	Chuck (service)	11/5	7 AM	2:30 PM	standard	serviceman	replaced relay on objective lens 
power supply	0		
Mon Nov  9 14:30:08 2015	Chuck (service)	11/9	8 AM	11:00 AM	standard	serviceman	reset network connector on computer 
and replaced air lines from 
compressor	0		
Tue Nov 10 17:19:37 2015	eileen	11/10	8:30 AM	4:30 PM	tiltrotate	rttomo	PTK2 supermontage section3	2x2 superof 3x3 montage		
Wed Nov 11 17:05:34 2015	eileen	11/11	8:45 AM	4:30 PM	tiltrotate	rttomo	PTK2 supermontage section4	2x2 super of 3x3 montages		
Thu Nov 12 15:25:14 2015	eileen	11/12	9:00 AM	12:30 PM	tiltrotate	rttomo	PTK2 montage reshoot	2 3x3 montage		
Wed Nov 18 16:47:23 2015	DNM	11/17	11:30 AM	4:50 PM	standard	development	high defocus stage calibrations			
Mon Nov 23 06:35:19 2015	DNM	11/22	10 AM	8 PM	standard	development	Focus-dependent mag testing			Stage positioning seems to have 1-2 
micron errors even with backlash 
correction
Thu Dec  3 14:50:36 2015	eileen	12/3	2 PM	2:30 PM	standard	rtimage	Tour with Linda Shoemaker Regent	0		
Mon Dec  7 16:54:02 2015	Cynthia/Laura	12/07	10:30 AM	4:45 PM	tiltrotate	rttomo	cos cells scram, drp1 kd	3 duals	beam/crud, loaded new alignments 
(5/11/15)and was OK! 5 hours of 
actual scope time.	
Tue Dec  8 18:58:01 2015	Cynthia/Laura	12/08	12:30 PM	6:30 PM	tiltrotate	rttomo	cos cells exp 8	6 duals	w9orked great!	
Thu Jan 28 11:35:14 2016	Chuck Moe	1/28	07:30 AM	11:30 AM	standard	serviceman	FEG replacement		replaced FEG and saved new 
alignments	no issues
Tue Mar  1 16:59:02 2016	eileen	3/1	11:00 AM	5:00 PM	tiltrotate	rttomo	plunge pombe	7 2x1 montage		
Wed Mar  2 13:56:57 2016	eileen	3/2	9:30 AM	2:00 PM	tiltrotate	rttomo	plunge pombe	6 (4 single frame, 2 1x2 montage)		
Mon Mar  7 16:42:34 2016	eileen	3/7	10:30 AM	4:30 PM	tiltrotate	rttomo	plunge pombe	5 2x1 montage		
Tue Mar  8 17:20:20 2016	eileen	3/8	9:30 AM	5:15 PM	tiltrotate	rttomo	pombe plunge	6		
Wed Mar  9 11:15:28 2016	eileen	3/9	9:00 AM	11:00 AM	tiltrotate	rttomo	plunge pombe	2 1x2 montage		
Thu Mar 10 17:25:32 2016	DNM	3/10	11:30 AM	5:30 PM	standard	development	testing for low mag montage crash	0		
Tue May 24 16:04:51 2016	Eileen (Natalie)	5/24	11:30 AM	3:45 PM	tiltrotate	rttomo	maize meoitic spindle	3 2x2 montage		
Thu May 26 15:50:27 2016	DNM	5/26	3:00 PM	5:00 PM	standard	development	looking at diffraction and lateral 
offset	0	Tilt axis offset is not ~4.5 
microns and should be fixed	
Thu May 26 15:51:23 2016	DNM	5/26	1:00 PM	4:00 PM	standard	development	looking at astigmatism with Garry 
offset	0		
Tue Dec 13 09:18:11 2016	DNM	12/12	9:30 PM	9:20 AM	standard	development	Testing focus limiting in tilt 
series			
Wed Jan 11 17:38:07 2017	Matthew West	1/11/17	12:45 PM	5:45 PM	fishione	rttomo	mutant yeast	6		Crashed the f-20 column, but was 
able to work from the f-30.  A few 
Serial EM glitches, but everything 
corrected itself if given a moment.